A Complete Bibliography of Publications in Journal of Fish Diseases (1980–1989)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

18 January 2025 Version 1.07

Title word cross-reference

14 [SLMH86]. ^{3} [SLMH86]. ^{35} [HD82]. _{2} [Mey80]. _{P 2} [Mey83]. 0-group [TJGA86]. 1254 [SSKP88]. 13 [Mey83]. 13p [Mey80]. 13P2 [BP88]. 2-year-old [Tra86]. 4-quinolone [TAR89]. 6 [AM83]. 8-tetrachlorodibenzo-p-dioxin [SSKP88]. = [AM83, HR80, TBS82, TBS83].

A-layer [EBS88]. Åland [Tho82]. abalone [Els83, EW83]. Abnormal [ML84, PMW80]. Abnormalities [RHI80, Poy87]. Abramis [SK82]. Abudefduf [RBML87]. acanthias [LL85a, Woo82]. Acanthocolpidae [Som81]. Acanthopagrus [Rou86]. accumulation [BG85, KNDLR86]. acheilognathi [ACCD81]. achromogenic [BR80]. Acid [KNDLR86, KNBLR86, SCM^{+}82, Buc80, Jac89, NAS88, WB85]. acid-alcohol-fast [Buc80]. acids [LMS85]. across [BA82]. actinosporean [Mar89]. Action [NS81]. activation [Sak84a]. activity [nCJS88, KYW83, LS80, Sak84a, TW89a]. acuminata [Ran86]. acute [ENH86, SRS80]. Addendum [Ano83a]. additional [AURE80]. adenoma [Fit88]. adenomas [FBV88]. Adenovirus [BMN86]. Adenovirus-like [BMN86]. adhesion [HB83a]. adjuvant [HRTW84]. administered [AJ84]. administration [GND^{+}87, LS80]. adult [CCK83, FTH87, LRM^{+}88, PR84, SI80, TBRH89, WBB85]. adults [LPPM86]. aeglefinus [BM88, MS81b, MC86, MP89a]. Aerococcus [Gje84]. Aeromonas [AJ84, BIII88a, BIII88b, BDDS88, BH86, BR80, CMZS84, DP86, DH84, EHM81, EBS88, EdGLV88, FFCM86, FCM88, nCJS88, HG82, HGB83, HE88, JA84, LD84, LdHvM85, MK89a, MK89b, ML89, MAJB83, Mcc83, MP80, MCSZ84, OEL85, PSDS89, REM89, Sak84a, Sak85, Sak87, TJGA86, TBRH89, TAR89, TRT89, VG88, WGC^{+}87, WC88]. aeruginosa [PRSC85]. aethiopicus [MIT84]. aetiological [CH88]. Aetiology [ES80b, ES80a, STEM80, CB82, CKGB^{+}84, SNZH80]. affecting [AJCM83, SZ87, Wen87]. affinis [CB81]. African [HTV88, MIT84]. after [BIII88b, CD86, EPPK86, GNV87, HTV88, Hof89, Jac89, Neu84, THJS87, Wil89a]. against [AHW83, AKF^{+}84, ASF87, BIII88a, BS81b, CD86, EHHJ88, EdGLV88, Gri89, HJ87, HRTW84, JA84, Lil89, ML89, MS82a, MCA84, Sak85, THJS87, Wil89a]. agassizi [FO85]. Age [BG85, DT81]. Age-dependent [BG85]. agent [AR85, BT86, CH88, DPS88, EHHJ88, FB87, Gri89, MB81, PDW81, STEM80]. agglutinating [MF82, Sak85]. aggregates [BG85]. Aggregation [Gat87]. Alabama [TH80]. Alaska [BG84, MHBM86]. Alaskan [FTH87]. alcian [HF80]. alcian-blue [HF80]. alcohol [Buc80]. alevin [BS84b]. alga [Lan86]. algal [LL85a]. alginolyticus [NB87a, NB87b]. alimentary [WTR88]. Allocreadiidae [AC80]. alpine [HKM86]. alpinus [HKM86, SK85]. altitude [BC85b]. altivelis [AKK81, AKF^{+}84, HFEJ86, KMKK83, Miy87]. alum [HRTW84]. aluminium [KNDLR86, KNBLR86]. ambigua [Lan86]. ambloplitis [BR84a]. Ameca [CP88]. America [DL89, SMRF84]. American [DH84, FHM85, MJ82, Mey80, MS83a]. americanus [EPB85a, EPB85b, Bog86, CW84]. Ameson [VA88]. Amia [AL87, MBH87]. aminoglycoside [SKSA^{+}86]. Ammodytes [DP86]. ammonia [DF84, Sod85]. Amoco [HNS^{+}82, SHNS83]. amoebiasis [NNS88]. among [AJ84, DF89, IIK^{+}84, LPPM86, SA84, SWDM83]. amongst [Mur81]. amphibians [SM80]. Amphiprion [BW83]. Amyloodinium [Pap80]. Anabantidae [PVdM87]. anaemia [TW88, TW89a]. anaesthesia [HD82]. anal [JA83b]. analyses [WSEL88]. Anguilla [Cop81, Cop82, CW82, Cop83, DH84, DBFW80, EDS83, HBF81, HM82, MB86b, MS83d, SON81, SLK89, SWDM83, WM82]. anguillae [MB86b]. anguillarum [JFA82a, JFA82b, AJ84, AGA80, AKK81, BKF83, HB83a, JA83b, KAF^{+}83, Miy87, RLRF84]. anguittiseptica [EDS83]. anisakid [Ela82]. Anisotremus [CNR81]. Ankylosing [LK84]. annual [SNZH80]. anomalies [TBS82]. anti [HC88, GNV87]. anti-Myxosoma [HC88]. anti-SRBC [GNV87]. antibacterial [AM88, Gri89, KYW83]. antibiotic [SKSA^{+}86]. antibiotics [CGBF87, GNV87, NS81]. antibodies [ACWM86, ASF87, BMR^{+}89, EHHJ88, MS82a, MF82, WAC89]. antibody [ACWM86, ABG^{+}86, FA84, HTV88, Lai80, Mar89, Olu86, PK88, SP87, Sak84b, Sak85, SG81, THJS87]. antibody-based [ABG^{+}86]. Antigen [Ell80, AHW83, FA84, HTV88, HSJE88, MAJB83, SM80]. Antigen-trapping [Ell80]. Antigenic [LD84, OSHF83, AJ84, NB87a, PK88]. antimicrobial [Aus85]. antimicrobials [TAR89]. antiserum [HC88]. Aphanomyces [AP86, APF87, PCS87]. Apicomplexa [KMF89, Lev84, MP89b]. appearance [RAL81]. Application [Olu86]. applied [JA83b]. Aqaba [PSC82, NP85, PHK80]. aquaculture [Ber80]. aquarium [SKS86]. aquatic [Bai83, OJBM88]. Arapaima [KNY^{+}87]. Arctic [HKM86]. Ardea [PN86]. areas [BB89, BA82, CMRM85, DF89]. Argopecten [Lei89, MS82b, MS83b]. Argulus [Ahn85]. Aristichthys [Sha81, SZA86, SR89]. arizonae [KNY^{+}87]. Aroclor [SSKP88]. Arthropoda [LR85]. artificial [SS88]. Ascetospora [BMR^{+}89]. aspects [Els80, FB87, RBRR87]. aspergillomycosis [OR86]. Aspergillus [Olu86]. assay [nCJS88, Lew86, RPT^{+}88, WAC89]. assays [ABG^{+}86]. assessment [BL86, BM86a, HG84, PK88]. associated [Ben80, BMN86, BR81b, CRC88, DK82, FRR^{+}86, GBC^{+}89, GCM83, Hau84, KWYE87, LPPM86, MAJB83, MHBM86, MGB83, NAS88, ND86, NS81, RL87, TT84, VG88, Wat82, YKN85]. Association [TBRH89]. astaci [AP86, APF87, PCS87]. Astacus [STF^{+}88]. asymptomatic [AMJS82, ARW83, BC85a, MK89a, MK89b, WBB85]. Atherinomorus [KMF89]. Atkinsiella [BE80]. Atlantic [ALWH87, BS84b, BM85a, BB89, Bru86a, BM86a, Bru86b, BE86, BS87, Bru89, BVN80, CH88, CW84, Cor86, CRS^{+}89, CMZS84, DG82, EJL81, EWM83, EMM85, EPB85a, EPB85b, EHHJ88, EH81, FRR^{+}86, FH85, FFCM86, FCM88, HS85, HAEE87, Hof89, Hog89, HJ87, JJ88, KWYE87, LPHA85, MB87, MS81a, MCSZ81, MCSZ84, MC86, MW86, Mur81, NND89, PDW81, RN81b, RVF86, RAL81, REM89, RLF89, SNZH80, TNH88, VK81, WSEL88]. atlanticum [VA88]. attached [RLF89]. attachment [LS83]. Attempts [FH80]. attributed [MS83d]. Atypical [CMZS84, DP86, FSC85, Hay80, MCSZ84, WGC^{+}87, WC88]. aurata [NP85, PHK80, Pap82, PSC82, Pap83]. auratus [AE80, BSG83, BH86, ES80b, ES80a, FODS89, STEM80, WC88]. aurea [TBS82, TBS83]. aureus [TBS82, TBS83]. aurofrenatum [GW83]. Australia [CK89a, CFV89, LH87, Lan87, LHW88, Lle80, MR88, PWL^{+}89, RPL89, WGC^{+}87, WC88]. Australian [HLGC87, LHC^{+}86, LHW^{+}86]. australis [Rou86]. Austrobdella [Rou86]. Austropotamobius [AP85, AFP86]. autoradiographic [RRB81]. auvatus [ML84]. avitaminosis [AM80]. Ayres [MS86, SW86, Ste86]. ayu [AKK81, AKF^{+}84, HFEJ86, KMKK83, Miy87].

B [AM83, Bul85, PW85, RLP85, TJGA86]. B-6 [AM83]. B-haemolysin [TJGA86]. B. [ACCD81]. baccatus [Som81]. bacteraemia [EBF^{+}87]. Bacteria [LPPM86, BA82, Buc80, BR81b, Don86, HLGC87, NS81, SWM^{+}87]. bacterial [Aus85, AR85, BSB89, Bru86b, CB82, CC85, CC86, CRC88, FODS89, HLGC87, KYW83, Lai80, LPHA85, MCA84, PDW81, SAAK87a, SCM^{+}82]. bactericidal [nCJS88]. bacterin [CD86, JA83a, LdHvM85, MAJB83, Miy87, NM82]. bacterins [AHW83, AJCM83, AJ84, JFA82a, JFA82b, JA83b]. Bacteriology [CK89a]. bacterium [AK85, DH84, EPPK86, Lle80]. baculoviral [MS88, MS89]. baculovirus [CK89b, Lew86, BNV^{+}86]. Baird [CB81, RLP85]. bairdi [MS80]. Bangladesh [SA80]. barbatulus [WTR88]. Barents [EJL81]. barnacle [Wen87]. based [ABG^{+}86]. Basin [HSR^{+}88, Hor87, APGL86]. bass [ABLY88, BR84a, CC84, MSWL84, Nog86, PAPBJ89, SMKH84, SRH85, SCR87, TH80, WWK^{+}87]. Bate [BE81, LRBB84, MS88, MS89, TB87]. bath [ALWH87]. batrachus [KRP87]. Battle [Hor87]. Bauer [Fij81]. Bay [KYBN88, Lei89, Mix83, MS82b, MS83b, BS81a, SNZH80]. bearded [Dav85]. beauty [GW85]. belonging [OJBM88, RR89]. Beneden [MSWL84]. Benoit [HNS^{+}82, SHNS83]. Benthiocarb [KOA81]. beryllina [SO82]. between [EBS88, MSWL84, NB87a, RB87]. Beyma [AP85]. big [DKHJ83, Sha81]. big-eyes [DKHJ83]. bighead [SZA86]. Bilateral [BF85]. bilobata [Rou86]. Binucleated [BS84b]. bioaccumulation [LS80]. biochemical [DF89, HG82, WSEL88]. Biography [Ahn86, BHS84, OM86, San82, SM89]. Birds [Wil89b]. birnaviruses [OJBM88]. bivalve [Els86, NB86, NB87b]. bivalve-pathogenic [NB86]. bivalves [NB86]. Bizzare [Ell84]. bjoerkna [AMS82]. BKD [Lai80]. Black [Mcv87, BE81, CB82, GSS81]. blacktail [WMC^{+}81]. Bleeker [Sha82]. Blennius [SE80b]. blenny [SE80b, SE80a]. Blicca [AMS82]. Bloch [AM83, CNR83, GW85, HHC^{+}81, LK84, MM88, SO82, Urd82]. Blood [MM88, BS84b, Ela82, HFH85, LPD85, WAC88]. blot [HCLW89, SAAK87b, SAAK87a]. blue [GC85a, HF80, Mac81]. bluegill [JC82, Mey80, MGB83, TR81]. BMN [MS88, MS89]. bodies [KOA81]. bodonid [Bul85]. body [CRC88, TBS82]. Bolivia [BC85b]. Bonamia [BMR^{+}89]. bonnethead [MRLB89]. bony [PWL^{+}89]. Book [Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano81c, Ano81d, Ano81a, Ano81e, Ano81b, Ano82a, Ano82b, Ano82d, Ano82c, Ano82e, Ano83b, Ano83c, Ano84a, Ano84b, Ano85a, Ano86a, Ano86b, Ano86c, Ano86d, Ano87a, Ano87b, Ano88a, Ano89a, Ano89b, Ano89d, Ano89c]. borne [Neu84]. Boschma [Wen87]. Bothriocephalus [ACCD81]. Botswana [MR82]. botulinum [CT84]. botulism [CT82, CT84]. bowfin [AL87, MBH87]. boxes [MB83]. Brachydanio [De 80]. brackishwater [NAS88]. Brain [LFDK89, MR82]. brama [SK82]. Branchial [NP85, Els86, TRT89]. branchialis [Lom80]. branchiphila [DLG83]. bream [AMS82, NP85, PWL^{+}89, Rou86, SK82]. brevifilum [MM80]. brevirostris [CNRC82, GGM85]. Brevoortia [FVO87, NND89, ND86, SNZH80]. Britain [APFH84, CT82, CS86]. British [Tra86, ACCD81, DBFW80]. Brittany [HNS^{+}82, SHNS83]. broodstock [BC85a, BMJ^{+}85]. brook [HFH85, ML89, MS83c, SCG82]. brown [BSB89, DPS88, DARD87, EMM85, FR80, KNDLR86, KNBLR86, Poy87, SFH89, SKSA^{+}86, WWB86, CNR81, Pap80]. Brunswick [Hog89]. Bucephalidae [PM84]. Buchanan [De 80]. bullatarudis [Sco85]. bullhead [DPS88, LFDK89, SFH89]. bullheads [GSS81]. bullminnows [RLP85]. Bunodera [AC80]. burbot [DSK87]. Burchell [HTV88]. Burkenroad [IKA87]. butterflyfish [CNR83]. byo [EDS83, SWDM83].

C [AM80, SLMH86, YKN85]. C-uracil [SLMH86]. Cadiz [HNS^{+}82, SHNS83]. cage [BS84a, MB87]. cage-reared [BS84a, MB87]. calcarifer [LK84]. calcified [MRLR84]. calcium [BW83]. California [Hor87]. calva [AL87, MBH87]. Canada [Con82, DSK87, Hog89, SZ87, SFH89]. Cancer [VA88]. canis [LL85b]. capillary [FOB85]. capricornensis [KMF89]. captive [CNRC82, NND89]. captivity [Gje84]. capture [BBN87]. Carassius [AE80, BSG83, BH86, ES80b, ES80a, FODS89, ML84, STEM80, WC88]. carbon [PR84]. Carcharhinus [SKSA^{+}86]. carchariae [GBC^{+}89]. carcinoma [Cor86, CRS^{+}89, FVO87, MJG84]. cardiomyopathy [FRR^{+}86]. Carolina [Nog86]. carolini [JC82]. carp [AM80, Ahn85, BIII88a, BIII88b, BC85a, CKGB^{+}84, EdGLV88, FA84, GNV87, GND^{+}87, KN89, LdHvM85, MG83, MRLR84, Mol80, MB81, Mol88, ME81, PD84, SOAK89, SM80, Sha81, SZA86]. carpio [BIII88a, BIII88b, BC85a, GNV87, GND^{+}87, KN89, KOA81, LdHvM85, MG83, MRLR84, Mol80, Mol88, ME81, PD84, SOAK89, SM80]. carps [SA80]. carrageenin [TTR85]. carrier [AMJS82, ARW83, MA88, MK89b, YMM82]. carriers [HKF80, MK89a]. cartilage [HC88]. caryophyllid [Hay80]. case [HFSPP88, Hor87, KMKK83, MBH87, MK83]. caseinase [DDMS89]. cases [ABG^{+}86, RR89]. cataractae [HFH85]. catfish [BMJ^{+}85, BMM85, BP88, BL85, CPH88, DLG85, HTV88, KRP87, LMS85, Lew85, MP80, MP85a, MM88, NNS88, NWGP89, SP87, SR80, SI80, TJGA86, VG88, WAC88, WAC89, WB85, WBB85, ZEK84]. Catla [MK83]. Catostomus [Hay80, SZ87, SFH89]. caught [BM88]. cauliflower [DBFW80, HM82]. causal [AR85, MB87]. causative [BT86, DPS88, EHHJ88, FB87, PDW81, STEM80]. caused [BE81, BG84, CNRC82, DT81, EDS83, EWM83, HFEJ86, Hau84, Hog89, JWS86, LRM^{+}88, Mcc83, MB83, RLRF84, Rob83, WTR88]. causes [SS88]. causing [Lle80]. cavefish [FO85]. Cell [BIII88a, MM80, SLK89, AHW83, DK82, FA84, FGL82, FVO87, FHW89, KK82, KPBS80, MJG84, RBML87, SLMH86, SLH88, TT84, Wat82]. cell-free [KK82]. cell-line [FHW89]. Cell-mediated [BIII88a, SLK89]. cells [BS84b, CK89b, HCLW89, LM85, Mod81, PSKS83, Sak84b, Sec85]. cellular [HA82, PM84, TH89]. central [MGB83]. centrifuge [MKS83]. Centropomus [Urd82]. Cépède [Cop81, Cop83]. cepedianum [TBRH89]. cephalus [LR81, BR81b, CK89a, CFV89, LR81, RB81]. Ceratomyxa [BSRF89, CCK83, HSR^{+}88]. cerebralis [HC87, HC88]. cernua [KP85]. Ceroid [BW83]. ceroidosis [RAL81]. Cestoda [ACCD81, SD86, TH89]. cestode [BS84a, HA82, WB80]. Chaetodon [CNR83]. challenge [ALWH87, BT86, Miy87, Sco85, WWB86]. Changes [AKK81, Bru86a, DARD87, CHBR87, DK82, Eck87, MR82, NAS88, SO82, TNH88, WS81]. Channa [AM83, NRC^{+}88]. channel [BMJ^{+}85, BMM85, BL85, CPH88, DLG85, LMS85, Lew85, MP80, MP85a, NWGP89, SP87, SR80, SI80, TJGA86, VG88, WAC88, WAC89, WB85, WBB85, ZEK84]. Chanos [LPPM86]. Characterization [BSRF89, BR84b, KNMY83, BNBKS81, CN80, Gje84, HSJE88, STEM80]. charr [HKM86]. Chelon [RM86]. chelonei [AFS86]. chemiluminescence [ABLY88]. Chemiluminescent [SRH85, SCR87]. chemistry [CMRM85]. chemosterilants [Mur81]. chemotherapeutants [Ald82]. chemotherapy [FH80]. Cherax [Her88]. Chesapeake [BS81a, SNZH80]. Chile [MR84]. chilodonelliasis [CVM80]. China [DKHJ83]. Chinook [BKF83, FTH87, Hau84, MK89a, MM85]. Chionocetes [MS80]. Chlamydia [MS82b]. Chlamydia-like [MS82b]. chlamydiae [DPS88]. Chlamydiosis [Lei89]. chlorinated [GSS81]. Chloromyxum [Pap83]. Chloromyxum-like [Pap83]. Chologaster [FO85]. chondrosarcoma [BKUW87]. Chowan [Nog86]. Chromatoblastomas [BB89]. Chronic [MM85, DF84, EH83, SRS80]. chum [BP88, FTH87]. churn [FTH87]. Ciliata [Dav85, DS88]. Ciliatea [CNR80]. ciliates [BAL87]. Ciliophora [FH80]. cinerea [PN86]. Cirrhina [AM80]. clams [EEC82]. Clarias [HTV88, KRP87, WB80]. classification [DL81, MS80]. clay [GFDM86, GFMD88]. clearance [HD82]. Cleidodiscus [TR81]. Clinical [WGC^{+}87, ABG^{+}86, BR81a, CHBR87, MB87, OR86]. clinically [SR89]. closed [GGL^{+}83]. Clostridium [CT84]. clownfish [BW83]. Clupea [BM86b, MHBM86]. clupeaformis [DR82]. co [ARW83]. co-cultivation [ARW83]. coastal [DLN^{+}89, SE80b]. coating [EEC82]. coccidia [DL81, LD82, MP89b]. Coccidian [KMF89, Lev84, MP89a]. coccidians [MR88]. coccidiosis [Mol82]. coccidium [HSO81, HSO83a, HSO83b, MB86b]. coccolithophorid [LL85a]. cod [CMZS84, EJL81, MS81a, MCSZ84, MP89b, RN81b, SE80b]. cognatus [Fit88]. Coho [EPPK86, LS81, NM82]. Coitus [Fit88]. Cold [EACR81, CVM80, TNH88]. Cold-water [EACR81]. coldwater [EHHJ88, HJ87, WSEL88]. colemanensis [CRC88]. colisarum [FHM85]. collare [CNR83]. collected [TB88]. colloidal [PR84]. Colombian [CVM80]. colonies [TRT89]. Columbia [HSR^{+}88, Tra86]. column [Poy87]. columnaris [KSDM86]. combination [KYW83, LS80]. combinations [AJ84]. combined [FCM88]. comments [Els86]. commercial [Wen87]. commersoni [Hay80, SZ87, SFH89]. common [Dia87, JC82, MRLR84, Mol80, Mol88, NS81]. Comparative [FGL82, SAAK87b, SAAK87a, SOAK89, SM80, TM82, Som81]. compared [FCM88]. Comparison [JA83a, RLRF84, AHW83, BP88, DPS88, FA84, FHW89, HGB83, HL85, SR89]. competition [AJ84]. Complement [TW89a, Sak84a, Sak84b]. components [TW89b]. compounds [Aus85, AM88]. concentrations [Bog86]. Conchiolin [EEC82]. condition [Mac81, SK87, Wat82]. conditions [APF87, AC80, KP85]. confidence [SS84]. connective [Wat82]. consideration [CHRB87]. containing [HG84, LMS85]. contaminated [OR86]. contamination [CT84, SLMH86]. Conte [MLG^{+}88]. contejeani [STF^{+}88]. Continued [Wil89a]. Control [BR84a, LPHA85, Aus85, Ber80, BP83, CVM80, FH80, GGL^{+}83, MR82, SCM^{+}82]. conversion [PW85]. Cope [SO82]. copepod [Don86, RR89]. Copepoda [Ben80, Hog89, RN81a, TH80]. copper [BKF83, Kni81]. Coregonus [DR82, RVF86]. correlation [Sak84a]. cortisol [WLL87]. Coryphaena [RL87]. Cottus [LFDK89]. counts [MS80]. course [BM85b]. Cowichan [Tra86]. crab [BE80, MS80, Wen87]. Cranial [BKUW87, KWYE87]. Crassostrea [Els80, EEC82, EBF^{+}87, Mey80]. crassus [BS84a, DR82, SD86]. Crayfish [APFH84, BS89, AP85, AFP86, PCS87, STF^{+}88]. Creek [BG84, Hor87]. Crenimugil [PM82]. critical [DL81]. croaker [EH81]. cross [Mar89, MJ82]. cross-protection [MJ82]. cross-react [Mar89]. cruciata [Nog86]. crude [BIII88b, EH81, HNS^{+}82, KK82, SO82, SHNS83]. Crustacea [BNS86, LR86, Ran86]. Cryptobia [BM85b, BS81a, JWS86, JW87, Lom80, TW88, TW89a, TW89b, WLL87]. cryptobiosis [TW89a, WLL87]. cryptosporidiosis [HHC^{+}81]. Crystallospora [Lev84]. crystals [BW83]. cultivated [BE80]. cultivation [ARW83]. culture [AP86, BGB82, BNBKS81, FA84, GGL^{+}83, KK82, LFYL81, NAS88, NB86, RPT^{+}88]. cultured [AKK81, AKF^{+}84, BAL87, BW83, BNS86, BVN80, CK89b, Cop81, Cop83, EH83, EEC82, EW83, FH85, Hog89, KK82, LHW88, LRBB84, NP85, NCL87, NRC^{+}88, NNS88, NSRS89, Pap80, PHK80, Pap82, Pap83, RLF89, SMKH84, Som82a, Som82b, SWDM83, WM82]. curvatures [Lan87]. Cutaneous [MG83, LR81, TM88]. Cuvier [DP86, DKHJ83, KNY^{+}87]. cyanobacterium [PRSC85]. cycle [Bog86, HC87, LPD85, PSZ81, PD84, VA88, WMH86]. cycles [DL81]. Cymothoidae [Ran86]. Cypnnus [Mol80]. cyprinacea [SKS86, TH80]. cyprinid [APGL86, DL89, MLG^{+}88]. Cyprinus [BIII88a, BIII88b, BC85a, GNV87, GND^{+}87, KN89, KOA81, LdHvM85, MG83, MRLR84, Mol88, ME81, PD84, SOAK89, SM80]. cysts [BE86]. cytopathology [RLF89].

dab [BMN86, BWF84, Dia87, OJBM88, Wat82]. dace [HFH85]. Dactylogyridae [TR81]. Dakota [HS82]. damage [HG84]. Dana [PCS87]. danio [De 80]. dead [FTH87, MP85b]. Decapoda [BNS86, Her88, LR85, LR86]. deepwater [BB89]. defence [Wil89a]. deficiency [AM80, AM83]. deficient [Sak87]. degeneration [BW83, FH85]. Degenerative [LMS85, NND89, FRR^{+}86]. degradation [Lil89]. delayed [BBN87]. delivery [JA83a]. demonstration [MPJ83a, Olu86]. densities [SK87]. density [MB83]. dependent [BG85, Jør82b]. deposition [BW83]. depression [EEC82]. deprivation [LS80]. derived [HC88]. Dermal [AE80, GW83, JSF82]. dermatitis [LL85a, LL85b, PWL^{+}89]. dermatobia [Cop82]. Dermocystidium [WM82]. Dermophthirius [CNRC82]. description [CNR80]. descriptions [LA89]. detect [nCJS88, SAAK87b, SAAK87a]. detecting [LM85, Lew86]. Detection [AK85, Lai80, MPJ83a, WB85, WBB85, FA84, HL85, HCLW89, RPT^{+}88, SP87, SLMH86, SS84, SG81, WAC89, YMM82]. determination [MAJB83]. developed [NAS88]. Development [CCK83, AC80, Fer81, LdHvM85, Mol88, SZ87, TM85a, TM85b, TNH88, TAR89]. Developmental [MS89, BSB89, PSZ81, PD84]. diagnose [MKS83]. diagnosis [ACWM86, ABG^{+}86]. Dicentrarchus [ABLY88, PAPBJ89]. diet [OR86]. dietary [HHF84, LPHA85, LS80]. Differences [RB87, DF89, LS81]. different [Bog86, Eck87, GND^{+}87, REM89, SS84, SK87, Wen87]. differential [MS80, SLMH86]. differentiated [AL87]. differentiation [LD84]. Digenea [Gat87, HFH85, PM84, Som81]. digenean [MS82a]. digestive [Lil89, MS82b]. Dinoflagellata [CNR81]. Dinoflagellida [Pap80]. Diodon [RN81a]. dioxin [SSKP88]. Diplostomatidae [Gat87]. diplostomiasis [BS81b]. Diplostomulum [MM88]. Diplostomum [Gat87, SRS80]. Direct [AURE80, NM82, AGA80, JFA82a, JFA82b, LdHvM85, Mar89]. discontinuous [WAC88]. disease [AM80, Aus85, AR85, BT86, BL87, BE81, BP83, Bru86b, CK89a, CFV89, CH88, CHBR84, CHA87, CHRB87, CHBR87, CD86, DMdK84, DF85, DF86, DBFW80, DLN^{+}89, EACR81, EH83, ES80b, ES80a, EWM83, EMM85, Els80, EPB85a, EPB85b, EHHJ88, EPPK86, FB87, Fer81, FRR^{+}86, FODS89, FH85, FHW89, FBS83, HL84, HJ87, Hor87, HLGC87, HM82, KN89, KSDM86, Lai80, LPHA85, LHC^{+}86, LH87, Lei89, LLGV83, LRBB84, LR85, Mar89, MCA84, MS80, MCSZ81, ML84, NSRS89, OBB^{+}85, PDW81, RB81, RLF89, SWM^{+}87, SBHM81, STEM80, SMRF84, TB87, WSEL88, WGC^{+}87, WMH86, ZEK84, HAEE87]. diseased [CMRM85, Con82, DP86, ES80b, KAF^{+}83]. diseases [Ber80, CC86, HLGC87, LHC^{+}86, NS81, SAAK87a]. disorder [BMM85]. dissolved [CC84, KH89]. distribution [Cop81, GND^{+}87, HSR^{+}88, MPJ83b, SBHM81]. diurnally [Sod85]. diverticula [MS82b]. dogfish [LL85a, LL85b, Woo82]. dollfusi [MRS88]. dolphin [RL87]. dominant [TBS83]. Dorosoma [TBRH89]. dot [HCLW89, SAAK87b, SAAK87a]. Dover [CB82]. Dr [Ahn86, San87]. drainage [Hor87]. Droncit [BS81b]. drug [AKK81, HG82]. Dual [HR80]. due [AM80, LFDK89, RM86, WGC^{+}87]. Duero [APGL86]. Dujardin [CNR80]. Dunbar [MS83c]. duorara [IKA87]. Duration [JFA82a, THJS87]. during [AM80, Bog86, CHBR87, EPPK86, MR82, NB87a, SS89, TNH88]. duroarum [IKA87]. Duszynski [HSO81]. Duwamish [PMW80]. dying [Gje84]. Dyková [LD82, Mol88]. Dynamics [Sco85]. dysfunction [LS81].

early [AM80, SWM^{+}87]. east [Bal83]. eastern [BR81b, CK89a, CFV89, GC82, WGC^{+}87]. Ecteinascidia [DH84]. ectocommensal [Lom80]. ectoparasite [Bul85]. Editorial [RW80, RW84]. edulis [BMR^{+}89, EEC82, GC85a, MB86a, Mix83, NB87a]. Edwardsiella [ACWM86, KRP87, MP85a, NWGP89, PS83, PK88, SP87, WAC89]. Edwardsiellaictaluri [SB87]. eel [DH84, EDS83, HM82, MS82a, MB86b, OSHF83, SLK89]. eels [Cop81, Cop82, Cop83, DP86, DBFW80, HBF81, MS83d, SWDM83, WM82]. Effect [HD82, LS80, LFYL81, MHSF84, MB83, SD80, Bul85, FFCM86, FCM88, KK82, LPSBL82, Mac81]. Effects [ABLY88, BM85b, HG82, HTV88, HKM86, SSKP88, BL87, CHA87, CC84, EH81, Fer81, GFMD88, HNS^{+}82, HSO81, HRTW84, LPBLS85, SR80, SHNS83, TBS82]. Efficacy [BAL87, JA83b, ML89, Sak85, JA83a, MCA84, MP80]. eggs [BSB89, EKPP84, EPPK86, LPSBL82, MB83, MP85b, SWM^{+}87, SS88]. Egtved [Jør82a, Jør82b]. Egusa [San87]. EHNV [LHW88, Lan89]. EHNV-like [LHW88]. Eilat [Pap80]. Eimeria [APGL86, ARV86, Bal83, DLG83, HSO81, HSO83a, HSO83b, Mac81]. Elbe [AM85, KP85]. electrolytes [ENH86]. Electron [MS81a, RPL89, AURE80, BSRF89, CPH88, DK82, LFDK89, MB81, MB86b, MS81b, RAL81, RDB84, SE80a, SS88, YKN85]. electrotherapy [FH80]. eliminate [MK89b]. ELISA [Lew86, RPT^{+}88, THJS87, WAC89]. elvers [CW82]. Emaciation [NRC^{+}88]. embedded [HC88]. Emeritus [San87]. emphasis [DS88, LS83]. encephalotropic [Lan87]. encystment [Som81]. endosulfan [MR82]. endothelial [PR84, TM85a, TM85b]. England [Bal83, Cop81, Cop82, SBHM81]. English [CMRM85, KPBS80, Rob83]. enhanced [BG84]. enhancement [DH84]. enlargement [Cou84]. Enteric [HHC^{+}81, ABG^{+}86, BP83, CD86, FBS83, LLGV83, NWGP89]. entry [Neu84]. environments [Wen87]. enzootic [Lan87, MW86]. Enzyme [RPT^{+}88, ABG^{+}86, DDMS89, Lew86, WAC89]. Enzyme-linked [RPT^{+}88, ABG^{+}86, Lew86, WAC89]. enzymes [DARD87]. Eosinophilic [RBML87]. eperlanus [AM85]. epidemic [SFH89]. epidemiological [BDDS88, EMM85, WGC^{+}87]. Epidermal [YKN85, BR80, BVN80, KNMY83, SZ87, SFH89]. epidermis [PR85, RRB81]. Epieimeria [LD82, MB86b]. Epiemeria [LD82]. Epistylididae [FHM85]. epithelial [BMN86]. Epitheliocystis [MS83a, ZEK84, DPS88, PSZ81, PD84, RDB84]. epithelium [Eck87]. epitopes [BMR^{+}89]. Epizootic [LH87, BG84, KMKK83, Lan89, MHBM86, MW86, Rob83, SON81, SMKH84, Tho82, Tra86]. epizootics [ML84, SNZH80]. Epizootiology [AR85, Pap80]. erebi [PWL^{+}89]. Ergasilus [Hog89]. Erratum [Ano86e, Ano88b, Ano85b]. erythrocytes [Ell84, SE80b, SE80a]. erythrocytic [MHBM86, PAPBJ89, RN81b, SE80b]. erythrodermatitis [EdGLV88]. erythromycin [HG84]. erythrophoroma [ME81]. erythrophthalmus [ARV86]. Esox [SD86, Tho82, TM88]. especially [BB89]. essential [WMH86]. establish [MK89a]. establishment [JJ88, Wil89a]. estuarine [RB81, SO82, TM82]. estuary [AM85, KP85, PMW80]. ethoxyquin [MHSF84]. Eubothrium [HKM86]. Eucoccidorida [KMF89]. Eudactylinidae [Ben80]. European [CPH88, Cop81, DT81, EDS83, FBV88, MB86b, NNS88, OSHF83, SWDM83]. Evaluation [Aus85, BDDS88, MK89a, CT84, MAJB83, MCA84, MS80, SLMH86]. Evidence [AJ84, EPB85a, EPB85b, TRT89, EKPP84, Lom80, PSZ81, WWB86]. examination [ES80b, TJGA86]. excretion [ENH86]. exlguus [PM82]. Exocrine [FBV88]. exophthalmos [ECM84]. exotoxin [BR84b, KK82]. experiences [Ber80]. Experimental [DMdK84, DR82, DLG85, GGM85, HAEE87, Lan89, Mey80, BL86, BH86, DG82, ES80a, EdGLV88, Hof89, KK82, MAJB83, Mcc83, Mey83, MF82, MS88, Neu84, NWGP89, NB87a, SR89, SS89, TTR85, VG88, WLL87, WWB86]. experimentally [CPH88, KNBLR86, MP80]. experiments [ENH86]. exposed [Cou84, MR82, Sod85, SO82]. exposure [BKF83, EH81, KNDLR86, KNBLR86, Kni81, NWGP89, REM89]. expression [BL87]. extracellular [EHM81, EBS88, FFCM86, FCM88, HE88, Sak84a]. extract [AHW83, DH84]. extracts [Dix87]. exudate [Sak84b]. eye [Gat87, SRS80, Sha81]. eyes [DKHJ83].

F [Wpl80]. Fabricius [CK89b, MS83a, NAS88, NNAS88]. factor [MB87]. factors [AJCM83, EBS88, nCJS88, GBC^{+}89]. faeces [Hof89]. failure [EKPP84]. fall [MM85]. farm [CT84, Rob83]. farmed [BE86, BC85b, CT82, KNDLR86, Kir80, RVF86]. farming [EACR81]. farms [ABG^{+}86]. fast [Buc80, SCM^{+}82]. FAT [Olu86]. fatal [EBF^{+}87]. fate [PR84]. fathead [Min81, MKS83]. fatty [CRS^{+}89, LMS85]. fed [HG84]. feed [PW85]. feeding [HTV88, OR86]. female [Con82]. fenestrae [KWYE87]. feral [MPJ83b, ML84]. fibroma [GW83, MRLR84]. fibromas [LR81]. fibrosarcoma [AE80]. fibrosis [MGB83]. field [APF87, AC80, MJ82]. filaments [Ben80, DLG83]. filtration [MB86a]. Fin [SK87, DBT89, Min81]. findings [WGC^{+}87]. Fine [Cor86]. fingerlings [Mey80]. Finland [BVN80, RVF86, Tho82, VK81]. First [LHW^{+}86, MHBM86, DBFW80, SMRF84]. Fish [DL81, AF81, APGL86, AK85, AFS86, AR85, ABG^{+}86, BL86, BLFR86, BDDS88, Ber80, Bru86b, BM83, CT84, CNR81, CHBR84, CVM80, CC86, DK82, DT81, DL89, DL80, Ela82, ES80b, FGL82, FHW89, Gat87, nCJS88, GGL^{+}83, Gri89, Hof89, HCLW89, KMF89, KYW83, KAF^{+}83, LD84, LdHvM85, LHW^{+}86, LM85, Lom80, LD82, LS83, LPD85, MGJ81, MR82, MLG^{+}88, Mcc83, MHSF84, NB87a, NS81, OMS83, Pap80, PN86, PS83, RL87, RBML87, RB81, Sak84b, SLMH86, SR89, SS84, SE80b, SE80a, SLH88, WC88, Wil85, Wil89a, Wil89b]. fish-pathogenic [LD84]. fisheries [Ald82, DLN^{+}89]. fishes [AJ84, BS81a, CNR80, FHM85, MH83, Mod81, MR88, MS81c, Sak84b, SM80, SO82]. Fitzsimons [CP88]. five [Dav85, FGL82, IIK^{+}84, PS83]. five-bearded [Dav85]. fixation [HF80]. fixed [Olu86]. flatfish [PSKS83, Som81]. flavescens [BG85]. flesus [FBV88]. flora [BSB89]. Florida [RL87]. flounder [CW84, EPB85a, EPB85b, FBV88, PMW80]. fluctuating [Sod85]. fluid [Ahn83, ENH86]. fluids [BM83, MPJ83b, NB86, RPT^{+}88]. fluke [Gat87]. flukes [SRS80]. fluorescent [ACWM86, FA84, Lai80, Mar89, Olu86, RAL81, SG81]. fluviatilis [DTC^{+}84, LHW^{+}86, LH87, Lan87, Lan89, PR84]. fly [MR82]. Foissner [FHM85]. foliaceus [Ahn85]. following [BKF83, CT84, GND^{+}87, Kni81, LdHvM85, NWGP89, SCG82, WWB86]. fontinalis [SCG82, HFH85, ML89, MS83c]. food [Eck87, LS80]. Forel [DR82, SD86]. form [DF85, MC86]. formalin [BAL87, Olu86]. formalin-fixed [Olu86]. formation [Sec85]. forms [Ell84]. Forsskål [BE80, LPPM86]. fossilis [MM88]. Fouquet [DLG85, FH80, GGL^{+}83, SS89]. four [MK89a, MH83, NS81, Som81]. Fowler [NRC^{+}88]. fractions [EH81, KK82, SO82]. fragility [EEC82]. France [HNS^{+}82, LLGV83, SHNS83]. Francisco [MSWL84]. free [Jør82a, KK82, LS81]. free-living [Jør82a]. frequency [Mur81]. fresh [GC82, TM82]. freshwater [BNS86, FHM85, Jac89, KAF^{+}83, LPD85, MR88, NCL87, PCS87, PWL^{+}89]. fry [AMS82, AT86, BS87, BR81a, CKGB^{+}84, DG82, HFEJ86, HK85, JJ88, JFA82b, KH89, MP85b, SD80, SS89, TH89, WS81]. function [HD82, Lom80, ML84]. funduli [HSO83a, HSO81, HSO83b]. Fundulus [RLP85]. fungal [EWM83, MS80, Mur81]. fungi [ND86]. fungicides [Bai83, LPSBL82, LPBLS85]. fungus [LRM^{+}88, PCS87, Wil89a, WWB86]. Furanace [MP80]. furdfer [Ran86]. Further [ARW83, CHRB87, DS88, EKPP84, MM86]. furunculosis [ALWH87, ABG^{+}86, BDDS88, DDMS89, EHM81, Mcc83, MF82]. Fusarium [AP85, BE81, MRLB89].

Gadidae [Dav85, DS88]. Gadus [CMZS84, EJL81, MS81a, MCSZ84, MP89b, RN81b, SE80b]. Gairdner [BT86]. gairdneri [ALWH87, AMJS82, ARW83, AHW83, Ahn83, AURE80, Aus85, AM88, BKF83, BL86, BSB89, Ber84, BA82, Bru86a, BM86a, Bru86b, BR80, BC85b, BS81b, CT82, CHRB87, CHBR87, Con82, CD86, CRC88, DMdK84, DF84, DF85, DF86, Dix87, DT81, EH83, ENH86, FB87, Fer81, FSC85, GD80, GC82, GFDM86, GFMD88, Gri89, HC88, HE88, Her85, HG84, HHF84, HD82, HL84, HTM^{+}82, HRTW84, HKF80, Jac89, JA83a, JWS86, JW87, Jør82a, Kni81, LHW88, LS80, LLGV83, Lil89, MJG84, MA88, MR84, MCA84, Mey83, MF82, MHSF84, MS83c, Neu84, NM82, OBB^{+}85, PR85, PRSC85, PW85, Rob83, RPT^{+}88, RBRR87, RDB84, Sak84a, SOAK89, SDC89, SD80, SRS80, Sod85, SSKP88, TM85a, TM85b, TW88, TW89a, THJS87, WGA^{+}84, WS81, WLL87]. Gallstone [Mat84]. Gambusia [CB81]. gammarus [Gje84]. Gams [AP85]. ganglioneuroma [KMKK83]. gariepinus [HTV88]. gas [CC84, KH89]. gasterostei [Don86]. Gasterostomum [PM84]. Gastrointestinal [MRS88]. gen [MM80]. gen.nov [RM86]. gene [TBS83]. genera [LD82]. genes [Sak87]. Genus [CN80, Lev84, LPD85, MS81a, RR89]. Geographic [HSR^{+}88]. geographical [DF89]. geometra [Ahn85]. Georgia [TH80]. Germany [FBS83]. Giant [TT84, LK84, Sec85]. Giant-cell [TT84]. giardi [Cop81, Cop83]. gigas [Els80, EBF^{+}87, KNY^{+}87]. Gilchrist [HS85, MS86, SW86]. Gill [SA80, TR81, AP85, Ben80, BE81, DF85, DF86, DLG83, KNDLR86, KNBLR86, MS81b, RLF89]. gill-attached [RLF89]. gills [BS81a, DK82, GFDM86, GFMD88, Lom80, MS81a, MS81c, RB87, RBRR87, TRT89]. gilthead [NP85]. Girard [CMRM85, CB81, KPBS80, RLP85, SK85]. gizzard [PWL^{+}89, TBRH89]. gland [MS88, MS89]. glanis [CPH88, NNS88, OMS83]. Glenn [OM86]. Glomerular [EH83]. glomeruli [Pap82]. glomerulonephritis [MM85]. glycogen [HD82]. Gmelin [Mey80, RVF86]. Gnathiidae [MS83d]. Gobio [Bal83, LFDK89]. gobionis [LPD85]. goby [Sha82]. golden [BP88, Lan86]. goldfish [AE80, BH86, CH88, ES80b, ES80a, FODS89, ML84, STEM80, WGC^{+}87, WC88]. gonad [ML84]. gonadal [DL89]. Goode [FVO87]. gourami [PVdM87]. gouramies [SCM^{+}82]. Goussia [LD82, MP89b]. gowkongensis [ACCD81]. gradient [WAC88]. grandis [RLP85]. granifera [Wen87]. granular [RBML87]. granuloma [PHK80, TTR85]. Granulomatous [BL86]. granuloplastic [AL87]. grayling [Eck87, HFSPP88, MW88]. Great [DL89, LS81]. greater [FBV88]. Greece [ARV86]. green [Ald85, Bog86, CHA87, Lan86]. grey [PM82, RM86]. Grimes [Lle80]. Group [BNS86, RLP85, TJGA86]. growth [AM80, Eck87, EEC82, LPSBL82, PW85]. growths [BM86b]. gudgeon [Bal83]. guideline [LA89]. gulf [FVO87, NP85, PHK80, PSC82]. Günther [HR80, PWL^{+}89, Rou86]. guppies [Sco85]. guppy [HFOW88]. Gurley [CN80]. gut [Mol82, MS88, MS89]. Gymnamoebae [RLF89]. Gymnocephalus [KP85]. Gyrodactylus [CC85, CRC88, JJ88, Sco85].

H [SLMH86]. H-uridine [SLMH86]. haddock [BM88, MS81b, MC86, MP89a]. haemangiomas [FOB85]. Haematological [BM86a, HL84, WSEL88, AURE80]. haematology [CHRB87, MHSF84]. haematopoietic [BM83, BG84, FGL82, FTH87, HL85, LH87, Lan89, MPJ83a, MB83, MPJ83b, MP85b, MW86, SSKP88, Tra86]. Haematractidium [BM85a]. Haemic [Mix83]. haemocyte [MS80, PCS87]. haemoflagellate [BM85b, JWS86, JW87]. Haemogregarina [Kir80]. haemolysin [FCM88, TJGA86]. haemolysins [EBS88]. haemolytic [RLP85, TW89b]. haemoparasite [BM85a]. haemopoietic [LFYL81]. haemorrhagic [AURE80, BGB82, DG82, EH83, ENH86, GD80, MW88, Neu84]. Haliotis [Els83, EW83]. Haliphthoros [LPBLS85]. hamanaensis [BE80]. Hamilton [AM80, De 80, KSDM86, MK83]. Hamilton-Buchanan [De 80]. Haplorchis [Som82a, Som82b]. haplosporean [DLF88]. hardened [EKPP84]. hardening [EPPK86]. Harding [Sha81]. hardy [KMF89]. harengus [MHBM86, BM86b]. hatchery [Ber84, Els80, FTH87, KMKK83, Lle80, MPJ83b, Rob83, VK81]. hatchery-reared [Ber84, KMKK83]. Hawaii [KYBN88, BNH^{+}86, BNV^{+}86, LRBB84]. Hawke [SP87]. head [BM86b, KMF89, Sha81]. Heckel [MIT84]. Heinz [Ahn86]. Heinz-Hermann [Ahn86]. held [REM89]. Henneguy [BR82, Bul85, CW84, RRB81, RB87, RBRR87, STF^{+}88]. Henneguya [Sha82, SI80]. Hennerty [Lle80]. Hepatic [RM86, BE86]. hepaticus [RM86]. hepato [BWF84]. hepato-splenic [BWF84]. Hepatocellular [CRS^{+}89, Fit88, Cor86, MBH87, RAL81]. hepatocytes [Cor86, KOA81]. hepatopancreas [DLF88, LR85]. Hepatopancreatic [BNH^{+}86, RPL89]. Herault [MLG^{+}88]. herbicide [Cou84, KOA81]. Hermann [Ahn86]. Hermaphroditism [SK82]. hernia [MK83]. heron [PN86]. Herpes [KNMY83]. Herpesvirus [WS81]. herring [BM86b, MHBM86]. heterogeneity [SDC89]. Heteropneustes [MM88]. Heteropolaria [FHM85]. Hidden [BG84]. high [BC85b]. higher [DPS88]. Hippoglossoides [MS83a]. hippurus [RL87]. Histochemical [RAL81, EH81]. histologic [YKN85]. Histological [BWF84, GFMD88, SR80, BR81a, HG84, Miy87]. histology [BSRF89, HF80]. Histopathologic [SHNS83, ECM84]. Histopathological [MR82, SO82, AURE80, EMM85, MB87, Mey83, Mix83]. Histopathology [BE81, Bru86b, CPH88, Cou84, DL89, EDS83, Els83, HNS^{+}82, MH83, MP85a, MCSZ84, PMW80, RN81a, RLP85, Rou86, SFH89, Sod85, CNR80, Ela82, HF80, JC82, MLG^{+}88, MRS88, MHSF84, MCSZ81, PHK80, Ran86, RLRF84, SRS80, Sha81, SR89]. Historesin(R) [HC88]. Historesin(R)-embedded [HC88]. history [Hor87, Som82a]. histozoic [CNR83]. Hitra [EACR81, EHHJ88, FH85, HAEE87, HJ87, WSEL88]. Hofer [HC87]. hoferi [TT84]. Hoffman [MS83c, OM86]. Holacanthus [GW85]. Holl [JC82]. Hollande [MB86b]. Homarus [Bog86, Gje84]. homogenates [BM83, RPT^{+}88]. homogeneity [PK88]. homologous [SLH88]. hormone [LPPM86]. hormone-implantation [LPPM86]. hornorum [LRM^{+}88]. Host [GC85a, SKS86, SW86, BSRF89, HC87, HB83a, LS83, MS86, RB87]. Host-parasite [SW86, MS86, RB87]. HPV [RPL89]. humoral [HE88]. hybrid [LRM^{+}88]. hydrophila [BIII88a, BIII88b, BH86, DH84, LD84, LdHvM85, MP80, TJGA86, VG88]. Hyperoplus [DP86]. hyperosmotic [AGA80, AKF^{+}84]. hyperplasia [YKN85]. hyperplasias [BMN86]. hyperplastic [KNMY83]. hyphal [LPSBL82]. hypoxia [SR80]. Hysterothylacium [MRS88]. hystrix [RN81a].

Ichthyobodo [Bul85, CW84, Dia87, MC86, RB87, RBRR87]. Ichthyophonus [TT84]. Ichthyophthiriasis [VK81, CVM80]. Ichthyophthirius [CP88, DLG85, FH80, GGL^{+}83, SS89]. Ichtyobodo [BR82, RRB81]. Ictalums [SI80]. ictaluri [ACWM86, KRP87, MP85a, NWGP89, PS83, PK88, SP87, WAC89]. Ictalurus [BMJ^{+}85, BMM85, DPS88, DLG85, GSS81, LMS85, Lew85, MP80, MP85a, NWGP89, SP87, SR80, SFH89, TJGA86, VG88, WAC88, WAC89, WB85, WBB85, ZEK84]. Identification [BR81b, PR85, BMR^{+}89, Lai80]. identified [WMH86]. Idiopathic [NCL87, LRBB84]. IFA [ACWM86]. IHHN [BL87]. II [BNS86, OJBM88]. immersion [AGA80, JFA82a, JFA82b, JA84, LdHvM85, Miy87, NM82, NWGP89]. immersion-exposure [NWGP89]. Immune [BIII88b, GNV87, HE88, JW87, Jør82b, LdHvM85, PDW81, Sak84b, SM80]. immunity [EdGLV88, JFA82a, JFA82b, MAJB83, SLK89, THJS87]. Immunization [AHW83, MCA84, BIII88b, Lil89, ML89]. immuno [GNV87, HCLW89]. immuno-pharmokinetic [GNV87]. Immunodiagnosis [BMR^{+}89]. immunofluorescent [HC88]. Immunogenicity [OEL85]. Immunoglobulin [SDC89]. immunological [BSRF89, LdHvM85]. immunoperoxidase [FA84]. immunosorbent [ABG^{+}86, Lew86, RPT^{+}88, WAC89]. Immunosuppression [JWS86]. impact [BC85b]. imperfect [LRM^{+}88]. implantation [LPPM86]. importance [Nog86]. imported [MR84, MW86]. inactivated [Sak85]. Inactivation [DH83, Sak84a]. inapparent [Jør82a]. including [LPD85]. inclusion [KOA81]. inclusion-like [KOA81]. inclusions [PSKS83]. Incompatibility [MSWL84]. incorporation [SLMH86]. increased [HD82]. increases [WLL87]. incubation [MB83, SS88]. Indian [AM80]. indirect [ACWM86, Lai80, SG81, WAC89]. induced [EdGLV88, HKF80, KNBLR86, KOA81]. Induction [OR86, ES80a, Mat84]. inert [GFDM86, GFMD88]. infected [BLFR86, Bru86a, BM86a, Bru86b, CPH88, HC88, Hay80, HCLW89, JC82, MP80, MRS88, RLP85, SD80, SR89, SD86, SS84, TW88, TW89a, WS81, DARD87]. infecting [DLF88, LPD85, Ste86]. Infection [BNS86, CK89b, Her88, KYBN88, BKF83, BE81, BMR^{+}89, BM85b, BH86, BNH^{+}86, CNR80, CP88, Cop82, CMZS84, CB81, De 80, DLG85, EEC82, EKPP84, FSC85, GGM85, GCM83, GC85a, Hof89, Hog89, HKF80, Jør82b, KK82, Kni81, Lan86, LR86, LPPM86, Mac81, ML89, Mey83, MS80, MS88, MCSZ84, Mur81, Neu84, NB87a, Pap82, Pap83, PVdM87, PAPBJ89, RBML87, RGLS81, RVF86, SCM^{+}82, SR89, SS84, SCG82, TH89, TT84]. infection-response [CP88]. infections [BDDS88, CW82, Cop83, DR82, DL80, Ela82, Els86, Jør82a, MA88, MGB83, MKS83, Sco85, SRS80, SE80b, SE80a, Som82b, SS89, VG88]. Infectious [FTH87, MR84, AF81, AMJS82, ARW83, Ahn83, BM83, BG84, DH83, Dix87, DT81, FGL82, HL85, HCLW89, HBF81, IIK^{+}84, LFYL81, MA88, MPJ83a, MB83, MPJ83b, MP85b, MW86, OSHF83, RPT^{+}88, SD80, SMKH84, SSKP88, SG81, SCG82, Tra86, WWK^{+}87, YMM82]. infectivity [BM83, Dix87, GC85b, LFYL81, REM89, Sak85]. infects [WC88]. infestation [RN81a, Ran86, Rou86, RBRR87, SI80]. infestations [Ben80, Pap80]. infested [Bul85, RRB81, Sha81, TR81]. infiltration [AKF^{+}84, CRS^{+}89]. inflammation [CKGB^{+}84]. inflammatory [EBF^{+}87, Wat82]. Influence [HHF84, DT81, GNV87, PCS87]. influencing [nCJS88]. Ingram [Rou86]. Inhibition [BR80, Dix87, MB86a, SLK89, WWB86]. initiation [Nog86]. injected [PR84]. injection [GNV87, HTV88, Miy87, SCG82]. insecticide [MR82]. interaction [SW86]. interface [RB87]. interlake [LS81]. Interlamellar [SI80]. intermediate [HC87]. intermoult [Bog86]. internal [CNR81]. Intestinal [JC82, Lan86, SD86]. intestinalis [HA82, TH89]. intoxication [Els83]. intra [EKPP84]. intra-ovum [EKPP84]. intracardiac [RR89]. intramuscular [TH89]. intraperitoneal [AHW83, DLG85, SCG82]. intravascularly [PR84]. Introduction [JJ88]. intubation [JA83b]. invasion [Som81]. invasive [GBC^{+}89]. investigation [Bru89, CB82]. Investigations [BM88, GNV87]. iodine [EKPP84, EPPK86]. ionized [Sod85]. Iowa [MGB83]. IPN [AF81, OJBM88]. IPN-like [OJBM88]. IPNV [RPT^{+}88]. iridovirus [BSG83, LHW^{+}86, LHW88]. iridovirus-like [LHW^{+}86]. Irminger [BB89]. irradians [MS82b, Lei89, MS83b]. irradiation [PW85]. Ishii [Cop82]. Islands [Tho82]. Isles [ACCD81, DBFW80]. isolated [AFS86, BE80, BA82, DP86, FFCM86, FCM88, HAEE87, KNMY83, KAF^{+}83, KNY^{+}87, LPSBL82, MRLB89, Sak84b]. isolates [BSG83, DF89, HGB83, MJ82]. isolating [AMJS82, ARW83]. Isolation [AF81, AMS82, AT86, BC85a, BMJ^{+}85, DSK87, HSJE88, HBF81, MLM88, OJBM88, SMKH84, AP86, CPH88, DTC^{+}84, FA84, LHW^{+}86, MR84, MP85b]. Isopoda [Ran86]. Isurus [Ben80]. Italy [GC82]. Iversen [IKA87].

Japan [AKK81, KAF^{+}83]. Japanese [HFOW88]. japonica [Els86, SON81, SLK89]. japonicus [BE81, LRBB84, MS88, MS89, TB87]. jiroveci [LFDK89]. John [Hog89]. Johnstone [PM84]. johnstonei [PM84]. Jordan [FOB85, FBV88]. juvenile [Bog86, BM85b, BR82, Els83, EW83, FTH87, GFMD88, KWYE87, MK89a, Mey80, NM82, RM86, RDB84].

K. [MS86]. Kabata [MS86, SW86, Ste86]. Kaneohe [KYBN88]. kaolin [GFDM86, GFMD88]. Katsuwonus [BBN87]. Katz [JWS86, JW87]. Ken [SM89]. keta [FTH87]. Kidney [BM80, Aus85, AR85, BT86, Bru86b, CNR81, CHBR84, CHA87, CHRB87, CHBR87, DMdK84, ENH86, Ell80, EMM85, EPPK86, FB87, Fer81, HF80, HL84, Lai80, LPHA85, LPD85, MCA84, MS83b, MP89a, MP89b, PDW81, SAAK87a, SBHM81, SMRF84]. kidney-disease [EPPK86]. kidney-infecting [LPD85]. kidneys [ARV86]. killifishes [HSO81, HSO83a, HSO83b]. killing [nCJS88]. kinds [Eck87]. kinetics [KPBS80]. Kinetoplastida [BS81a, TW89a, TW89b]. Kingdom [HBF81, SE80b]. Kirmse [Kir80]. kisutch [EPPK86, LS81, NM82]. Klinke [Ahn86]. koi [CS86]. Kokanee [Tra86]. Kritsky [CRC88, MRS88]. Kudo [CS86]. Kudoa [HS85, MS86, Pap82, SW86, Ste86]. Kuruma [BE81, LRBB84, MS88, MS89].

L. [ALWH87, AE80, AMS82, Ahn85, AM85, AL87, ABLY88, ARV86, BIII88a, BIII88b, Bal83, BL86, BSB89, BC85a, BS84b, BSG83, BMN86, BM85a, Bog86, BMR^{+}89, BA82, BH86, Bru86a, BM86a, Bru86b, BE86, BS87, BM88, Bru89, Buc80, BWF84, BR82, Bul85, BM86b, BR81b, BF85, BVN80, CK89a, CFV89, CB82, CH88, CNR81, CPH88, Cop81, Cop82, CW82, Cop83, CMZS84, Dav85, DS88, DG82, DBFW80, DBT89, Dia87, DTC^{+}84, DARD87, DSK87, DLG83, Eck87, EJL81, Ela82, ES80b, ES80a, Ell80, EDS83, EWM83, Ell84, EMM85, EHHJ88, EH81, FR80, FRR^{+}86, FODS89, FH85, FBV88, FFCM86, FCM88, Gje84, GNV87, GND^{+}87, GC85a, HNS^{+}82, HS85, HK85, HAEE87, Hof89, HKM86, HFSPP88, Hog89, HJ87, HA82, HBF81, HM82, JSF82, JJ88, KN89, KNDLR86]. L. [KNBLR86, KRP87, KWYE87, Kir80, KP85, KOA81, LPHA85, LdHvM85, LHW^{+}86, LH87, Lan87, Lan89, LL85a, LFDK89, LR81, MBH87, MG83, MRLR84, MM80, MB87, MB86a, Mcv87, MW88, Mix83, Mol80, Mol82, MB86b, Mol88, MS81a, MS81b, MCSZ81, MS83c, MCSZ84, MC86, MP89a, MP89b, MS83d, MRLB89, MW86, ML84, ME81, Mur81, NP85, NNS88, NND89, NB87a, OMS83, OJBM88, OR86, OM86, PR84, PHK80, Pap82, PSC82, Pap83, PD84, PDW81, PAPBJ89, Poy87, PM84, RN81a, Ran86, RL87, RBML87, RN81b, RGLS81, RVF86, RAL81, RB81, REM89, RBRR87, RLF89, SOAK89, SM80, SKS86, SD86, STEM80, STF^{+}88, SK82, SE80b, SWDM83, SHNS83, TBS82, TH89, TB88, Tho82, TM88, TH80, TT84, TTR85, TNH88, TRT89, VK81, VA88, WSEL88, WTR88]. L. [Wat82, Wen87, WC88, WWB86, Woo82, WM82]. Labbé [LD82]. labelling [HC88]. Labeo [KSDM86]. Laboratory [BS81b, APF87, AC80, Bru86b, BM86b]. laboratory-reared [BM86b]. labracis [Hog89]. labrax [ABLY88, PAPBJ89]. labrosus [PM82, RM86]. Lacépède [Hay80, Nog86, SZ87, SFH89, TH80]. Lacistorhynchus [MSWL84]. Lack [EBS88, AJ84]. Lactobacillus [Con82]. laevis [WTR88]. Lagenidium [LPSBL82]. lake [HKM86, KH89, SK87, SZ87, Tra86, Urd82]. Lakes [DL89, LS81]. Lamarck [Lei89, MS83b]. lamna [Ben80]. Lampetra [BT86, PR84]. lamprey [BT86, PR84]. lancea [DP86]. lanceolatus [DP86]. large [SKS86]. largemouth [Nog86, TH80]. larva [HA82]. larvae [BM86b, Eck87, LPSBL82, MS88, NB87a, RHI80]. larval [CC84, Ela82, Els80, Lei89, MGB83, MRS88]. Lates [LK84]. latipes [HFOW88]. Latrobe [ND86, SNZH80]. lavage [CCK83]. lavalaei [DK82]. Laveran [Bal83]. layer [HSJE88, EBS88]. leech [Rou86]. Léger [MB86b]. Leiby [MRS88]. Leidy [BR84a]. lemon [CNRC82, GGM85]. length [TBS82]. leniusculus [PCS87]. lenses [Gat87]. Lepomis [JC82, Mey80, MGB83, TR81]. Lereboullet [AP85]. Lernaea [Nog86, Sha81, SKS86, SR89, TH80]. Lernaeenicus [RR89]. Lernaeidae [TH80]. Lernaeoceridae [RN81a]. Lesauvege [DP86]. lesion [SR89]. Lesions [GW85, TB88, VG88, BB89, BM88, BR81a, CNRC82, DBT89, ECM84, Hay80, KNDLR86, KNBLR86, Nog86, TR81, Woo82]. less [TBS82]. Lesueur [DPS88, SFH89, TBRH89]. lethal [Sak84a, TBS83]. leucocytes [Ela82, WAC88, YMM82]. leukaemia [AL87]. Leukaemic [Ber84]. level [HTV88]. levels [BAL87, LS81, SS84, Sod85, SKSA^{+}86]. life [DL81, HC87, LPD85, Som82a, VA88, WMH86]. lifestage [SWM^{+}87]. Light [DK82, MB86b, MS81b, GGL^{+}83, LFDK89, MB81, RDB84]. Ligula [HA82, TH89]. like [BMN86, Els86, GC85b, KOA81, LHW^{+}86, MLM88, MS82b, NNAS88, Pap83, RPL89, YKN85, LHW88, OJBM88]. Limanda [BMN86, BWF84, Dia87, OJBM88, Wat82]. line [FHW89]. lines [FGL82, SLMH86, SLH88]. linked [ABG^{+}86, Lew86, RPT^{+}88, WAC89]. lipopolysaccharide [BIII88b, SP87]. liquefaciens [Lle80]. Liquid [BL85]. lituratus [HHC^{+}81]. liver [Ela82, EH81, HSO81, HD82, KP85, MJG84, MR82, TB88]. livers [PMW80]. living [GSS81, Jør82a]. Lmk [MS82b]. loach [WTR88]. loading [MB83]. lobsters [Bog86, Gje84]. localization [SM80]. locations [ES80b]. Lom [LD82, LS83]. Loma [Hau84, MS83c, MM86]. long [HNS^{+}82, HTM^{+}82, SHNS83]. long-term [HNS^{+}82, HTM^{+}82, SHNS83]. longnose [HFH85]. Looss [Som82a, Som82b]. Lorn [Mol88]. Lota [DSK87]. low [HTM^{+}82, Jør82a, Jør82b]. low-virulence [Jør82b]. lower [DL89, Hog89, Mur81, TB88]. luciopercae [AC80]. lucius [SD86, Tho82, TM88]. lungfish [MIT84]. Lycodes [DK82]. Lymphocystis [PVdM87, Urd82, EPB85a, EPB85b, PSC82]. lymphocytes [PR85]. lymphoid [CK89b]. lymphoma [Tho82]. lymphoproliferative [BMM85]. Lymphosarcoma [HFSPP88, Ber84, WGA^{+}84]. Lysozyme [Gri89].

MacCallum [KYBN88]. Macfarlane [MS82a]. Mackenzie [DSK87]. mackerel [BM85a, Buc80]. Macquaria [Lan86]. macracanthus [DKHJ83]. Macrobrachium [BNS86, NCL87]. macrochirus [JC82, Mey80, MGB83, TR81]. macrogamont [HSO83a]. macrophage [BG85, nCJS88]. macrophages [BNBKS81]. maculatus [SO82]. magellanicus [GCM83, GC85b]. magna [Min81]. maintained [LMS85]. major [AM80, RBML87, SA80]. makos [Ben80]. Malachite [Ald85, Bog86, CHA87]. Malaysia [NNAS88]. male [Mur81]. Malmberg [JJ88]. mammalian [SLH88]. Man [BNS86, NCL87, RPL89]. mangrove [BE80]. Manning [IKA87]. Maracaibo [Urd82]. marble [Sha82]. marginatus [BNH^{+}86, BNV^{+}86]. marine [BAL87, BS81a, CNR80, DK82, KYBN88, LR86, LD82, MB86a, MC86, NB86, NB87a, Pap80, SE80b, SE80a, TM82]. marinum [CNR80]. markers [BDDS88]. marmoratus [Sha82]. marquandii [LRM^{+}88]. Martius [BE81]. mass [GCM83]. Mastigophora [BS81a]. Matthews [PM84]. maximus [BL86, DBT89, Kir80, MM80]. MBV [CK89b]. measure [MS80]. measured [SLK89, THJS87]. mebendazole [BR84a]. mechanical [Ahn85]. mechanism [TW88]. medaka [HFOW88]. mediated [BIII88a, KSDM86, SLK89]. Mediterranean [PAPBJ89]. Meinert [Ran86]. Melanogrammus [BM88, MS81b, MC86, MP89a]. Melanotic [MGB83]. melas [GSS81]. melleni [KYBN88]. memory [LdHvM85]. menhaden [DLN^{+}89, FVO87, NND89, ND86, SNZH80]. Menidia [SO82]. mentella [BB89]. Mercenaria [EEC82]. merguiensis [RPL89]. Merlangius [Ela82]. merlangus [Ela82]. Merluccius [MS86, SW86, Ste86]. mesenchymal [BM86b]. mesentery [Pap82]. Mesothelioma [Her85]. metacercariae [MM88]. metacercarial [RGLS81]. metacestode [MSWL84]. metchnikovi [Bal83]. Method [Bai83, ARW83, AKF^{+}84, HCLW89, MK89b, MS82a, MS88]. methods [HL85, JA83a, MK89a, SAAK87b, SAAK87a]. micro [MS82a]. micro-method [MS82a]. microbiological [ES80b, WGC^{+}87]. Microbiology [MCSZ81]. microcolonies [CC85, CRC88]. Microcultures [LM85]. Microcystis [PRSC85]. Microgadus [Cor86, CRS^{+}89]. microgamete [HSO83b]. microgamont [HSO83b]. Microgemma [RM86]. Micromesistius [Mac81]. Micropogon [EH81]. Micropterus [Nog86, TH80]. microscope [DK82, LFDK89, MS81b, RDB84]. microscopic [MB81, MB86b, RPL89, YKN85]. microscopical [AURE80, MS81a, RAL81, SE80a]. microscopy [BSRF89, CPH88, SS88]. Microspora [CN80, IKA87, MM80]. microsporida [MS81a, MS81b, Her88, MS83c]. microsporidan [Hau84, VA88]. Microsporidian [MS81c, CB81, De 80, DL80, MLG^{+}88, RBML87]. microsporidiosis [RM86]. microsporidium [Her88, Mod81]. mid [MS88, MS89]. mid-gut [MS88, MS89]. midgut [Eck87, LRBB84]. migration [BR80, SLK89]. milkfish [LPPM86]. Miller [CP88]. minimum [MGB83]. minnow [Min81]. minnows [MKS83]. mirandellae [MLG^{+}88]. Mirex [LS80]. Mitchill [BG85, DR82, HFH85, HS82, HB83b, KNMY83, LL85b, ML89, MLM88, MS83c, SCG82, YKN85]. Mizelle [CRC88]. model [ALWH87, CP88, Mcc83]. Moderation [TTR85]. Modification [AKF^{+}84, LPHA85]. molluscs [Els86]. Molnár [ARV86]. monitor [MKS83]. Monoclonal [ABG^{+}86, EHHJ88, ACWM86, ASF87, BMR^{+}89, PK88]. monodon [CK89b, CK89b, LR85, LPSBL82, NAS88, NNAS88]. monodon-type [CK89b]. Monogenea [CC85, CRC88, Sco85, TR81]. monogenean [KYBN88]. monogeneans [CNRC82]. Montta [VK81]. mordax [HB83b, MLM88]. morhua [CMZS84, EJL81, MS81a, MCSZ84, MM86, MP89b, RN81b, SE80b]. moribund [FTH87]. Morone [CC84, MSWL84, SMKH84, SRH85, SCR87, WWK^{+}87]. Morphological [DF89, TNH88]. morphology [CNR80]. Morrison [MM86]. mortalities [DT81, EEC82, FR80, Lle80, TBRH89]. Mortality [Hog89, BBN87, Eck87, GCM83, Hau84, MHBM86, MS83d, MB83, SS88]. mosquitofish [CB81]. mossambicm [WB80]. mossambicus [LM85, KYBN88, LRM^{+}88, Mat84, SS88, SS89]. mouse [HC88]. mouthbrooding [SS89]. Moxostoma [FBV88]. mrigala [AM80]. mucophilosis [MB81]. mucus [WWB86]. Mueller [TR81]. Mugil [BR81b, CK89a, CFV89, LR81, RB81]. muksun [RVF86]. Muller [HC87, AC80, WTR88, WMH86]. mullet [BR81b, CK89a, CFV89, LR81, PM82, RM86, RB81]. multicystic [BF85]. multifiliis [CP88, DLG85, FH80, GGL^{+}83, SS89]. multinucleate [Sec85]. Multiple [BE86, FOB85, LR81]. multiplication [Neu84]. Multivalvulida [HS85]. Murray [PWL^{+}89]. muscle [NCL87, SW86, Ste86]. Muscular [FH85]. muscularis [CNR83]. mussel [GC85a]. mussels [Mix83]. mustela [DS88, Dav85]. Mustelus [LL85b]. mutant [Sak87]. mycobacterial [BM80]. Mycobacterium [AFS86, BNS86, LR86]. Mycoplasma [SLMH86]. mycosis [DLN^{+}89, HFEJ86, LRM^{+}88, ND86]. Mycotic [PWL^{+}89]. Mycteroperca [FOB85]. myocardial [FH85]. myopathy [HS82, LMS85]. Mytilus [GC85a, MB86a, Mix83]. Myxidium [Cop81, Cop83]. myxoboliasis [LFDK89, SA80]. myxobolid [MKS83]. Myxobolus [Cop82, CS86, LFDK89, PM82]. Myxosoma [HC87, HC88]. Myxosporea [CNR83, Dav85, DS88, HS85, LPD85, LA89, PM82]. myxosporean [BSRF89, HSR^{+}88, Lan87, LPD85, Mar89, MS86, Pap83, Ste86]. myxosporeans [SW86]. myxosporidan [WMC^{+}81]. Myxozoa [Dav85, DS88, Lan87, Min81, Sha82]. myxozoan [Min81].

N [Fij81, MM80, Mod81]. naive [SR89]. namaycush [KH89, SK87]. Nardo [SKSA^{+}86]. naso [HHC^{+}81, HHC^{+}81]. Natural [AM88, CB81, MW86, VG88]. naturally [Bru86b, SR89]. NCMB [NB86, NB87a, NB87b]. nebulosus [DPS88, SFH89]. necator [BR82, Bul85, CW84, RRB81, RB87, RBRR87]. necrosi [CB82]. Necrosis [LR85, AF81, AMJS82, ARW83, Ahn83, BM83, BG84, DH83, Dix87, DT81, FGL82, FODS89, FTH87, HL85, HCLW89, HBF81, IIK^{+}84, JSF82, LH87, Lan89, LFYL81, MA88, MR84, Mcv87, MHBM86, MS88, MS89, MPJ83a, MB83, MPJ83b, MP85b, MW86, NCL87, OSHF83, RN81b, RPT^{+}88, SD80, SMKH84, SE80b, SSKP88, SG81, SCG82, Tra86, WWK^{+}87, YMM82]. Negaprion [CNRC82, GGM85]. Nematalosa [PWL^{+}89]. Nematoda [JC82]. nematode [Ela82]. Nemesis [Ben80]. Neobenedenia [KYBN88]. neoplasia [HR80, MBH87]. neoplasms [DL89, MH83, Mix83].

Nephroblastoma [HB83b]. nephrocalcinosis [HHF84]. nerka [AGA80, BG84, MB83, MP85b, Tra86]. Nerocila [Ran86]. nerve [LRBB84]. newborn [MRLB89]. Newfoundland [Con82]. newly [Lei89]. Niacin [PW85]. Nicoll [Som81]. Nie [Lom80]. nifurpirinol [CVM80]. niloticd [TBS82]. niloticus [TBS82, OR86]. nobilis [SZA86, SR89]. noble [STF^{+}88, BSRF89, HSR^{+}88]. noblis [Sha81]. nocardiosis [AFP86]. Nodular [DF85, Mol82, DF86]. nodules [BL86, KWYE87]. Noemacheilus [WTR88]. nom [PM84]. Non [WMC^{+}81, AF81, EPB85a, EPB85b, LD84, RLP85, Sak84a]. non-haemolytic [RLP85]. non-lymphocystis [EPB85a, EPB85b]. Non-myxosporidan [WMC^{+}81]. non-salmonid [AF81]. non-specific [Sak84a]. non-virulent [LD84]. North [DL89, Nog86, APGL86, CW84, EPB85a, EPB85b, GC82, MC86, BB89, BWF84, Dia87, FHM85, HS82, MJ82, SMRF84]. north-eastern [GC82]. north-west [APGL86, CW84, EPB85a, EPB85b, MC86]. northern [ARV86, JJ88, RN81b, RVF86, Rou86, SD86, Tho82, TM88, VK81]. Norway [JJ88, STF^{+}88]. Norwegian [EACR81]. Nosematidae [Her88]. note [Buc80, BR82, Don86]. Notes [LPD85, DL81]. nov [BE80, CNR83, HFH85, Lan87, LFDK89, MS83c, PM84, Sha82, VA88]. novel [nCJS88]. Nuclear [PSKS83]. nucleic [WB85]. number [PCS87]. Nutritional [AM80].

O [HTV88]. O-antigen [HTV88]. O. [FTH87, LRM^{+}88]. observation [RPL89]. Observations [AC80, BVN80, DDMS89, Kir80, MS86, ARW83, AURE80, BGB82, BM85a, DK82, DPS88, Ela82, EMM85, LFDK89, MM86, VA88, YKN85]. observed [NNAS88]. obtained [DF89]. Occurrence [BLFR86, De 80, Els86, GD80, Gje84, Hof89, MF82, PAPBJ89, RVF86, ACCD81, BR82, BR81b, Cop81, DBFW80, HHF84, Jør82a, KMKK83, Mix83]. occurring [SMKH84]. Ocean [BLFR86, CW84, MC86, RN81b]. ocellaris [BW83]. ocellatum [CNR81, Pap80]. Ocular [RGLS81]. offshore [SE80b]. oil [EH81, HNS^{+}82, MHSF84, SO82, SHNS83]. old [Tra86]. Oleg [Fij81]. oligochaete [WMH86]. oncogene [SLH88]. Oncorhynchm [MB83, NM82]. Oncorhynchus [AGA80, BKF83, BM85b, BG84, EPPK86, FTH87, Hau84, LS81, MK89a, MM85, MP85b, Tra86]. Onset [JFA82b]. Ontario [SZ87, DPS88, SFH89]. ontogenetic [TM85a, TM85b]. Oodinium [CNR81]. Oomycete [ND86]. operations [Ber80, MR82]. Ophiocephalus [AM83]. Ophryoglenidae [FH80]. opisthorchis [MS82a]. opportunistic [GBC^{+}89]. Opsanus [BF85, RGLS81, TB88]. Opsonization [Sak84b]. optimum [PW85]. oquassa [SK85]. Oral [Lil89, AHW83, JA83b]. ordalii [RLRF84]. Oregon [FGL82, MM85, Mix83]. Oreochromis [KYBN88, LRM^{+}88, Mat84, OR86, SS88, SS89]. organ [CK89b]. organism [GC85b, Lai80]. organisms [MS82b, Olu86]. organochlorines [LS80]. organs [BWF84, PMW80]. origin [WGA^{+}84]. ornamental [ME81].

Ortholineidae [Lan87]. Oryzia [HFOW88]. Osmerus [AM85, HB83b, MLM88]. Osteichthyes [Ran86]. osteochondroma [NP85]. Ostrea [BMR^{+}89, EEC82, NB87a]. ostreae [BMR^{+}89]. other [CNR81, FHW89, Lan89, NB86, PMW80]. Outbreak [LLGV83, CT82, CT84, FBS83, MJG84, PSC82, SMRF84, SWDM83]. Outbreaks [LHW88, WGC^{+}87]. ova [BE80]. ovarian [MPJ83b]. ovaries [SHNS83]. overlay [BL85]. Overstreet [HSO81]. ovum [EKPP84]. oxalate [BW83]. oxidized [MHSF84]. oxolinic [Jac89]. Oxyeleotns [Sha82]. oxygen [Els83]. oxyrinchus [Ben80]. oxytetracycline [Bru89, GNV87, GND^{+}87, Jac89]. oyster [EBF^{+}87, RGLS81, TB88]. oysters [Els80, EEC82, Mey80].

p [SSKP88]. P. [DKHJ83, LR85]. p53 [SLH88]. Pacifastacus [PCS87]. Pacific [BLFR86, BT86, BM85b, EBF^{+}87, KPBS80, MHBM86, MS86, PSKS83, RLRF84, SW86, Ste86]. paddlefish [BKUW87, MRS88]. Paecilomyces [LRM^{+}88]. Pagenstecher [Don86]. pagurus [VA88]. Pallas [BS84a, PMW80, RVF86, SCM^{+}82]. pallasi [MHBM86]. pallipes [AP85, AFP86]. pancreas [FRR^{+}86, TB88]. pancreatic [AF81, AMJS82, ARW83, Ahn83, DH83, Dix87, DT81, FBV88, HCLW89, HBF81, IIK^{+}84, MA88, MR84, OSHF83, RPT^{+}88, SD80, SMKH84, SG81, SCG82, WWK^{+}87, YMM82]. paniformis [MS86, SW86, Ste86]. Papillomas [GSS81, PSKS83, SZ87, SFH89]. papillomata [BVN80]. Papillomatosis [OMS83, AM85]. parameters [Bru86a, EH81, MM88]. Paramoeba [RLF89]. paramoebic [RLF89]. Paramoebidae [RLF89]. Paranthias [Ran86]. Parasite [RL87, AP85, BSRF89, CNR83, HFH85, HSR^{+}88, MLG^{+}88, MS86, RB87, Sha82, SW86]. Parasite-associated [RL87]. parasites [CC86, KMF89, MS86, MS81c]. parasitic [Don86, PCS87, Wen87]. parasitism [RR89]. parasitized [MM88]. parasitizing [CRC88, HSO83a, HSO83b, MB86b, STF^{+}88, VA88]. parasitology [CK89a]. Parastacidae [Her88]. parenterally [WS81]. parenterally-infected [WS81]. Parophrys [CMRM85, KPBS80]. parr [EWM83, EMM85, Hog89, JJ88, Mur81]. parrotfish [GW83]. particles [BMN86, YKN85]. particular [RB81]. parva [AT86]. parvo [RPL89]. parvo-like [RPL89]. passage [BA82]. passive [ML89]. Pasteurella [AK85]. patch [CB82, Mcv87]. pathogen [BS84a, GBC^{+}89, Hof89, LHW^{+}86, NB87b]. Pathogenesis [EW83, RRB81, SCG82, Dia87, DDMS89, HB83a, KK82]. pathogenic [AK85, BR84b, BR81b, DH84, DTC^{+}84, LD84, NB86, NB87a, PN86]. Pathogenicity [GC85b, APF87, DL81, HAEE87, Kir80, Lan89, Mey80]. pathogens [Gri89, KYW83, SS84]. Pathological [Eck87, KP85, NAS88, CHBR87, DK82, FFCM86, WS81]. Pathology [AM83, CFV89, EBF^{+}87, MGJ81, NWGP89, PHK80, BM86b, CW82, Cop83, DF84, DS88, EHM81, JSF82, Pap80, RL87, SD86, Som82b, TM88, WB80, WTR88]. Pathophysiology [BH86]. Patinopecten [Els86]. patronus [FVO87]. patterns [MPJ83b]. PCB’s [LS80]. pectoralis [PVdM87]. Peculiar [RR89]. pelagicus [Wen87]. pelamis [BBN87]. peled [RVF86]. Pellérdy [ARV86]. penaei [BNV^{+}86]. penaeid [BNH^{+}86, DLF88, Lew86, RPL89]. Penaeus [BL87, BE81, BNH^{+}86, BNV^{+}86, CK89b, DLF88, IKA87, LRBB84, LR85, LR86, LPSBL82, MS88, MS89, NAS88, NNAS88, RPL89, TB87]. Peniculisa [RN81a]. pennellid [RR89]. Pentacapsula [CNR83]. Pentacapsulidea [CNR83]. Perca [BG85, DTC^{+}84, LHW^{+}86, LH87, Lan87, Lan89]. percae [Lan87]. perch [BG85, DTC^{+}84, LHW^{+}86, Lan86, LH87, Lan87, Lan89, LK84]. Percoll [WAC88]. Percoll(R) [HC88]. Percoll(R)-purified [HC88]. Perfusion [HF80]. peripheral [BS84b, WAC88]. peritoneal [Sak84b]. peritoneum [Pap82]. peritrichous [BAL87]. Persistence [EPPK86, HTM^{+}82]. Peters [HFOW88, LM85, Mat84, Sco85, SS88, SS89, WB80]. phagocytes [ABLY88, SRH85, SCR87]. Phagocytosis [GFDM86, Sak84b]. Pharmacokinetics [GND^{+}87]. pharmokinetic [GNV87]. phase [ENH86]. phenax [FOB85]. Phialophora [EWM83]. philippinensis [LPBLS85]. pholis [SE80b]. Phoma [HFEJ86]. Phylum [CN80]. picorna [MLM88]. picorna-like [MLM88]. pigmented [BB89]. pike [AMS82, AT86, SD86, Tho82, TM88]. pilchard [RR89]. pilchardus [RR89]. pili [LD84]. pillulare [LS83]. Pimephales [Min81, MKS83]. Pink [IKA87]. pirarucu [KNY^{+}87]. piscicida [AK85]. Piscicola [Ahn85]. piscinae [Sha81]. piscine [PSZ81, SE80b]. Pisdnoodinium [LS83]. pituitary [Cou84, ML84]. pituitary-gonad [ML84]. PKD [CHRB87, SBHM81]. PKX [FB87]. Placopecten [GCM83, GC85b]. plague [APFH84, BS89]. plaice [BR82, Bul85, Ell80, Ell84, MS83a, PM84, SHNS83, TT84, TTR85]. plankton [MKS83]. plaquing [BL85]. plasma [HD82]. plasmid [SB87]. plasmids [AK85]. platessa [BR82, Bul85, Ell80, Ell84, HNS^{+}82, PM84, SHNS83, TT84, TTR85]. platessoides [MS83a]. Platichthys [FBV88, PMW80]. Plecoglossus [AKK81, AKF^{+}84, HFEJ86, KMKK83, Miy87]. Plehn [HK85]. Pleistophora [CN80, DKHJ83, MLG^{+}88]. pleomorphic [PSZ81]. plerocercoid [HA82]. Pleuronectes [BR82, Bul85, Ell80, Ell84, HNS^{+}82, PM84, SHNS83, TT84, TTR85]. plumbeus [SKSA^{+}86]. Poecilia [HFOW88, Sco85]. Poecilostomatoida [Hog89]. Poey [CNRC82, GGM85]. Point [TH80]. poliovirus [TM82]. polluted [CMRM85]. polybrene [LFYL81]. polymorpha [SR89]. Polyodon [BKUW87, MRS88]. Polyonchobothrium [WB80]. Pomphorhynchus [WTR88]. pond [GSS81, NRC^{+}88]. pond-cultured [NRC^{+}88]. ponds [NAS88, Sod85]. Poorly [AL87]. population [BR82, BG84, CB81, SS84]. populations [JSF82, MPJ83b, ML84, SFH89, Wen87]. pork [CNR81]. Portunus [Wen87]. position [Lev84]. Possible [AFP86, CKGB^{+}84, MC86, SWM^{+}87, SS89]. Post [FR80]. Post-spawning [FR80]. Posthodiplostomum [MGB83]. postmetamorphic [Lei89]. potassium [HRTW84]. potency [AJCM83].

Potential [DF86, GBC^{+}89, JA84, Wil89b]. potentially [NAS88]. poutassou [Mac81]. povidone [EKPP84, EPPK86]. povidone-iodine [EKPP84, EPPK86]. prawn [BE81, BNS86, LRBB84, NCL87, NAS88, NNAS88, RPL89]. precocious [Mur81]. Preliminary [Ber80, WWB86, DTC^{+}84, MCA84, Wil85]. preparation [FCM88, LA89]. Presence [Ahn83, BR80, Don86, EPB85a, EPB85b, EKPP84]. present [SWM^{+}87]. Presporogonic [Mol88]. presumptive [AURE80, EH83]. Prevalence [BM86b, TH80, HM82, MKS83, Mix83, RN81b, RB81, SBHM81, SS84, SE80b, SFH89, WB80, YKN85]. Priacanthus [DKHJ83]. primary [GNV87]. pro [PM84]. pro- [PM84]. probable [LR86]. probe [WB85]. produced [BR84b, NB87a, NB87b]. Production [ASF87, NB87a, HC88, HTV88]. products [EHM81, FCM88, HE88, Sak84a]. productus [MS86, SW86, Ste86]. Professor [San87, Ahn86, Bau80, Fij81]. profile [SB87]. profiles [BDDS88]. progression [Mix83]. Prokaryotic [GCM83]. proliferation [DK82]. proliferations [Ben80]. Proliferative [CHBR84, EMM85, SMRF84, CHA87, CHRB87, CHBR87, DMdK84, DF85, FB87, Fer81, HL84, LRBB84, SBHM81]. prolonged [SR80]. promelas [Min81, MKS83]. properties [KN89]. prophylaxis [Mcv87]. proposed [HC87]. protease [FFCM86, FCM88, Sak85, Sak87]. protease-deficient [Sak87]. Proteases [PSDS89, EBS88]. protected [Lil89]. Protection [CD86, BIII88a, BIII88b, HTM^{+}82, MJ82, PW85, SS89]. protective [EdGLV88, MAJB83, THJS87]. protein [ENH86]. proteinase [NB87a, NB87b]. proteinases [NB87a]. Proteocephalus [BR84a, MGB83]. Protopterus [MIT84]. Protozoa [FHM85, Lev84, Min81, Sha82]. protozoal [NSRS89]. protozoan [CKGB^{+}84, WMH86]. pseudobranch [BM88]. Pseudobranchial [EJL81]. Pseudomonas [EDS83]. Pseudophyllidea [ACCD81, BS84a, SD86, TH89]. pseudophyllidean [HA82]. Pseudopleuronectes [CW84, EPB85a, EPB85b]. Pseudorasbora [AT86]. pseudotuberculosis [MGJ81]. Puerto [GW85]. Puget [CMRM85]. Pugheadedness [SZA86]. pumilio [Som82a, Som82b]. punctatm [MP80]. punctatus [BMM85, WB85, AM83, BMJ^{+}85, DLG85, LMS85, Lew85, MP85a, NWGP89, SP87, SR80, SI80, TJGA86, VG88, WAC88, WAC89, WBB85, ZEK84]. Purification [NB87b, TJGA86]. purified [HC88, LMS85]. putative [EBS88]. Putnam [FO85]. Putz [MS83c].

quadricarinatus [Her88]. Queensland [BR81b, Wen87]. quinolone [TAR89].

R [AK85]. rabbits [HE88]. Radhakrishnan [RN81a]. radiation [Bul85, BC85b, MB87]. Rafinesque [Ben80, BMJ^{+}85, BMM85, DLG85, GSS81, JC82, LMS85, Lew85, Mey80, Min81, MP80, MGB83, MKS83, MP85a, NWGP89, SP87, SR80, SI80, TR81, TJGA86, VG88, WAC88, WAC89, WB85, WBB85, ZEK84]. rainbow [ALWH87, AMJS82, ARW83, AHW83, Ahn83, AURE80, Aus85, AM88, BKF83, BL86, BSB89, Ber84, Bru86a, BM86a, Bru86b, BR80, BC85b, BS81b, CT82, CHRB87, CHBR87, Con82, CD86, DMdK84, DF84, DF85, DF86, Dix87, DT81, EH83, ENH86, FB87, Fer81, FSC85, GD80, GC82, GFDM86, GFMD88, Gri89, HC88, HE88, Her85, HG84, HHF84, HD82, HL84, HTM^{+}82, HRTW84, Jac89, JA83a, JWS86, JW87, Jør82a, LHW88, LS80, LLGV83, Lil89, MJG84, MA88, MR84, MCA84, Mey83, MF82, MHSF84, MS83c, Neu84, NM82, OBB^{+}85, PR85, PRSC85, PW85, Rob83, RPT^{+}88, Sak84a, SOAK89, SDC89, SD80, SRS80, Sod85, SSKP88, TM85a, TM85b, TW88, THJS87, WGA^{+}84, WS81]. Randall [BNH^{+}86, BNV^{+}86]. Rapid [Wil85, ABG^{+}86, MS82a, SG81]. rate [EKPP84, TH80]. rates [REM89]. Rathbun [MS80]. rations [HG84, LMS85]. react [Mar89]. reaction [LdHvM85, TT84]. reactions [DL80, HG82, Hau84, MM80]. reared [BS84a, Ber84, BM86b, DBT89, Eck87, Els80, GD80, KMKK83, Lle80, MB87, MW86, Rob83, SK87]. recipient [HKF80]. recirculation [GGL^{+}83]. record [BNV^{+}86, CS86]. recovery [DF86, HD82]. Red [Pap80, PSC82, BS84b, BR81b, CK89a, CFV89, Els83, EW83, LR85, LRM^{+}88, RB81]. redband [GW83]. redescription [CN80]. redfin [LHW^{+}86, LH87, Lan87, Lan89]. redfish [BB89]. redhorse [FBV88]. redmouth [ABG^{+}86, BP83, CD86, FBS83, Kni81, LLGV83, Lle80]. reduce [EKPP84]. reducing [BAL87]. reduction [TTR85]. reference [Bru86b, RB81]. Regan [PVdM87]. region [MW86]. Regression [SZ87]. Reichenbach [Ahn86]. Reichenbach-Klinke [Ahn86]. Reichenow [Bal83]. related [BB89]. relation [Eck87, SWM^{+}87]. relations [MS86]. relationship [BSB89, EBS88, NB87a, TW89a]. relationships [IIK^{+}84, OSHF83]. Renal [Mol80, BW83, ENH86, LRM^{+}88]. Renibacterium [AJ84, ASF87, AR85, BLFR86, BT86, Bru86a, BM86a, EKPP84, EPPK86, PDW81, SOAK89]. Renibacteriwn [SAAK87b]. renicola [Mol88]. reo [NNAS88]. reo-like [NNAS88]. reovirus [BP88, Mey80, Mey83]. Replication [YMM82]. report [BR81a, CC85, HFSPP88, MBH87, MHBM86, Rob83]. reported [Lei89, SMRF84]. requirement [PW85]. requirements [SS84]. rerio [De 80]. Reservoir [TH80]. residual [CT84]. residues [Bru89]. Resistance [BT86, AKK81, DH84, HG82, PCS87, SSKP88, TAR89]. resistant [SR89]. response [BSRF89, CP88, GNV87, GC85a, HE88, HA82, JW87, Jør82b, Miy87, PDW81, PM84, RBML87, Som81, TH89]. responses [Mey83, SRH85, SCR87]. restricted [BA82]. Retention [BM83]. reticulata [HFOW88, Sco85]. reticulo [PR84, TM85a, TM85b]. reticulo-endothelial [PR84, TM85a, TM85b]. retinal [BW83]. Retinoblastoma [FO85]. retinopathy [NND89]. retrovirus [YKN85]. retrovirus-like [YKN85]. returning [MM85]. review [Ald85, CHBR84, CC86, Ano81a, Ano81b, Ano82a, Ano82b, Ano82c, Ano84a, Ano86a, Ano87a, Ano88a, Ano89a, Ano89b, Ano89c]. Reviews [Ano80a, Ano80b, Ano80c, Ano80d, Ano80e, Ano80f, Ano81c, Ano81d, Ano81e, Ano83b, Ano83c, Ano84b, Ano85a, Ano86b, Ano86c, Ano86d, Ano87b, Ano89d, Ano82d, Ano82e]. rhabdomyoma [KMKK83]. Rhabdomyosarcoma [HFOW88]. rhabdophilia [Mod81]. rhabdovirus [AMS82, AT86, DTC^{+}84, FHW89]. rhabdoviruses [FHW89]. Rhinichthys [HFH85]. Rhipidocotyle [PM84]. Richardson [ALWH87, AMJS82, ARW83, AHW83, Ahn83, AURE80, Aus85, AM88, BKF83, BL86, BSB89, Ber84, BA82, Bru86a, BM86a, Bru86b, BR80, BC85b, BS81b, CT82, CHRB87, CHBR87, Con82, CD86, CRC88, DMdK84, DF84, DF85, DF86, DKHJ83, Dix87, DT81, EH83, ENH86, FB87, Fer81, Fit88, FSC85, GD80, GC82, GFDM86, GFMD88, Gri89, HC88, HE88, Her85, HG84, HHF84, HD82, HL84, HTM^{+}82, HRTW84, HKF80, Jac89, JA83a, JWS86, JW87, Jør82a, Kni81, Lan86, LHW88, LS80, LLGV83, Lil89, MJG84, MA88, MR84, MCA84, Mey83, MF82, MHSF84, MS83c, Neu84, NM82, OBB^{+}85, PR85, PRSC85, PW85, Rob83, RPT^{+}88, RBRR87, RDB84, Sak84a, SOAK89, SDC89, SD80, SRS80, Sha81, SZA86, SR89, Sod85, SSKP88, TM85a, TM85b, TW88, TW89a, THJS87]. Richardson [WGA^{+}84, WS81, WLL87]. rickettsia [GC85b]. rickettsia-like [GC85b]. rickettsial [BNH^{+}86, GC85a]. rickettsiales [Els86]. rickettsiales-like [Els86]. Rickettsias [MS83b]. Rico [GW85]. Risso [Ben80, Mac81, PM82, RM86]. River [DSK87, HSR^{+}88, Hog89, Hor87, MM85, Nog86, PMW80, TB88, MPJ83a, PR84, Sak87, HM82, JJ88, PWL^{+}89]. rivers [MLG^{+}88]. roach [DLG83, HA82, TH89]. róbalo [Urd82]. robustus [TR81]. rock [GW85]. rockling [Dav85]. rodlet [Mod81]. Rogue [MM85]. rohita [KSDM86]. rohu [KSDM86]. role [DDMS89, EHM81, HC87, HB83a, Mcv87, TJGA86]. rosenbergii [BNS86, NCL87]. rostrata [DH84]. routes [AHW83, GND^{+}87, REM89]. ruckeri [JFA82a, JFA82b, MJ82, AJCM83, AJ84, CGBF87, CD86, DF89, DSK87, FSC85, HTV88, HKF80, JA83a, JA83b, Kni81, NM82, RVF86, Rob83, SCR87, SA84, Wil89b]. rufescens [Els83, EW83]. ruffe [KP85]. Runyon [BNS86]. russelli [DS88, Dav85]. Rutilis [TH89]. Rutilus [ARV86, DLG83, HA82].

S [HD82, TBS82]. S-sulphobromophthalein [HD82]. S. [BA82, LPD85, Wpl80]. sac [KH89]. Sacculina [Wen87]. sachai [Kir80]. Sacramento [Hor87]. Saddleback [TBS83, MCSZ81]. Saint [Hog89]. salar [ALWH87, BS84b, Bru86a, BM86a, Bru86b, BE86, BS87, Bru89, BVN80, CH88, DG82, EWM83, EMM85, EHHJ88, FRR^{+}86, FH85, FFCM86, FCM88, HS85, HAEE87, Hof89, Hog89, HJ87, JSF82, JJ88, KWYE87, LPHA85, MCSZ81, MW86, Mur81, PDW81, RVF86, RAL81, REM89, RBRR87, RLF89, TNH88, TRT89, VK81]. salaris [JJ88]. salinity [BM85b]. Salmo [ALWH87, AMJS82, ARW83, AHW83, Ahn83, AURE80, Aus85, AM88, BKF83, BL86, BSB89, BS84b, Ber84, BA82, Bru86a, BM86a, Bru86b, BE86, BS87, Bru89, BR80, BC85b, BVN80, BS81b, CT82, CH88, CHRB87, CHBR87, Con82, CD86, CRC88, DMdK84, DF84, DF85, DF86, DG82, Dix87, DT81, DARD87, EH83, ENH86, EWM83, EMM85, EHHJ88, FB87, FR80, Fer81, FRR^{+}86, FH85, FSC85, FFCM86, FCM88, GD80, GC82, GFDM86, GFMD88, Gri89, HC88, HS85, HE88, Her85, HG84, HHF84, HD82, HAEE87, Hof89, HL84, Hog89, HJ87, HTM^{+}82, HRTW84, HKF80, Jac89, JSF82, JJ88, JA83a, JWS86, JW87, Jør82a, KNDLR86, KNBLR86, KWYE87, Kni81, LPHA85, LHW88, LS80, LLGV83, Lil89, MJG84, MA88, MR84, MB87, MCA84, Mey83, MF82, MHSF84, MCSZ81]. Salmo [MS83c, MW86, Mur81, Neu84, NM82, OBB^{+}85, PDW81, PR85, PRSC85, PW85, Poy87, RVF86, RAL81, Rob83, RPT^{+}88, REM89, RBRR87, RLF89, RDB84, Sak84a, SOAK89, SDC89, SD80, SRS80, Sod85, SSKP88, TM85a, TM85b, TW88, TW89a, THJS87, TNH88, TRT89, VK81, WSEL88, WGA^{+}84, WS81, WLL87, WWB86]. salmoides [Nog86, TH80]. salmon [ALWH87, AGA80, BKF83, BS84b, BM85b, BP88, Bru86a, BM86a, Bru86b, BE86, BS87, Bru89, BG84, BVN80, CH88, DG82, EWM83, EMM85, EHHJ88, EPPK86, FRR^{+}86, FH85, FTH87, FFCM86, FCM88, HS85, Hau84, HAEE87, Hof89, Hog89, HJ87, JSF82, JJ88, KWYE87, LPHA85, LS81, MK89a, MB87, MM85, MCSZ81, MB83, MPJ83b, MP85b, MW86, Mur81, NM82, PDW81, RLRF84, RVF86, RAL81, REM89, RLF89, SWM^{+}87, TNH88, VK81, WSEL88]. salmonae [MS83c]. Salmonella [KNY^{+}87]. salmonicida [HG82, HGB83, MCSZ84, OEL85, REM89, Sak85, AJ84, BDDS88, BR80, CMZS84, DP86, EHM81, EBS88, EHHJ88, EdGLV88, FCM88, nCJS88, HGB83, HE88, HSJE88, Hof89, JA84, MK89a, MK89b, ML89, MAJB83, Mcc83, PSDS89, REM89, Sak84a, Sak87, TAR89, TRT89, WGC^{+}87, WC88]. Salmonid [BNBKS81, Mar89, WMH86, AF81, AJ84, AFS86, AR85, BLFR86, BR81a, CHBR84, EACR81, EKPP84, JFA82b, MH83, Mod81, MS81c, Sak84b, TM82, WC88, Wil89a]. salmonidda [FFCM86]. salmonids [BS84a, BT86, CCK83, HLGC87, JFA82a, JA83b, LHC^{+}86, Lle80, MK89b, NSRS89, RRB81, RB87, RBRR87, WMC^{+}81]. salmoninarum [AJ84, ASF87, AR85, BLFR86, BT86, Bru86a, BM86a, EKPP84, EPPK86, PDW81, SAAK87b, SOAK89]. salmonis [WS81]. salmositica [BM85b, JWS86, JW87, TW88, TW89a, TW89b, WLL87]. salvelini [HKM86]. Salvelinus [HFH85, HKM86, KH89, ML89, MS83c, SK85, SK87, SCG82]. sample [SS84]. sampling [CCK83]. San [MSWL84]. sand [DP86, Mcv87, Wen87]. sand-eels [DP86]. Sanguinicola [HFH85]. Sanguinicolidae [HFH85]. Saprolegnia [CW82, DARD87, Wil85, Wil89a, WWB86]. Saprolegnia-infected [DARD87]. Saprolegniasis [BS87]. Sarcomastigophora [TW89a, TW89b]. Sardina [RR89]. Sarotherodon [HR80, LM85, RHI80, TBS82, TBS83]. saxatilis [SRH85, CC84, MSWL84, RBML87, SMKH84, SCR87, WWK^{+}87]. scallop [GCM83, GC85b, MS82b, MS83b]. scallops [Lei89]. scamp [FOB85]. Scanning [SS88, BSRF89]. scardinii [ARV86]. Scardinius [ARV86]. Schaperclaus [Bau80, LS83]. Schioedte [Ran86]. Schlegel [AT86, AKK81, AKF^{+}84, HFEJ86, HFOW88, KMKK83, Miy87, SON81, SLK89]. Schmidt [MRS88]. Schneider [HHC^{+}81, LD82, SO82]. Schrank [HKM86]. Schubert [FHM85]. Scomber [BM85a, Buc80]. scombri [BM85a]. scombrus [BM85a, Buc80]. Scophthalmus [BL86, DBT89, Kir80, MM80]. Scotland [EDS83, HM82, SWDM83, WM82]. Scottish [BM88]. screening [Bai83, MS82a, Wil85]. sculpin [Fit88]. Scylla [BE80, LPSBL82]. sea [ABLY88, BR81b, CK89a, CFV89, GD80, GCM83, GC85b, JSF82, NP85, PAPBJ89, RB81, REM89, BB89, BWF84, Dia87, DKHJ83, EJL81, Pap80, PSC82]. sea-water [GD80]. Seasonal [RB81, LS81]. seasonality [Mix83]. Seattle [PMW80]. Sebastes [BB89, ECM84]. secreted [PSDS89]. sediments [Sak87]. Sekiten [EDS83, SWDM83]. Sekiten-byo [EDS83, SWDM83]. sekoke [KN89]. selected [EH81, HGB83, NS81, OSHF83]. selenium [HHF84]. seminal [Ahn83]. seminoma [MIT84]. sensitive [AMJS82]. sensitivities [SAAK87a]. Sensitivity [Bog86, CGBF87, FGL82, SAAK87b, SOAK89]. separation [BNBKS81, WAC88]. septicaemia [AURE80, BGB82, DG82, EH83, ENH86, GD80, KNY^{+}87, MW88, Neu84, NWGP89]. sequence [SLH88]. sequential [CHBR87]. sera [MS82a]. sergeant [RBML87]. Seriola [KK82]. serious [DLN^{+}89, Els80, Lei89]. serogroup [OJBM88]. Serological [BP88, HGB83, IIK^{+}84, Mey83, SP87, SA84, FHW89]. Serology [AFS86]. serotypic [MJ82]. Serotyping [KAF^{+}83]. Serranidae [Ran86]. serrata [BE80, LPSBL82]. Serratia [Lle80]. Serum [CMRM85, Bru86a, DARD87, LS81, SP87]. seven [ES80b]. several [JA83a, SAAK87a]. Severe [FRR^{+}86]. sewage [GSS81]. shad [PWL^{+}89, TBRH89]. shaharini [Sha82]. shark [SKSA^{+}86]. sharks [CNRC82, GGM85, GBC^{+}89, MRLB89]. shasta [BSRF89, CCK83, HSR^{+}88]. sheat [OMS83]. sheat-fish [OMS83]. shedding [REM89]. shell [EEC82]. shellfish [BR84b, NB87a]. shellfish-pathogenic [BR84b]. shiner [BP88]. shortfin [Ben80]. shrimp [BAL87, BL87, BNH^{+}86, DLF88, LR86, MS88, MS89, TB87, IKA87]. significance [Els86, EBF^{+}87, MF82]. Silurus [CPH88, NNS88, OMS83]. similarities [Lle80]. sites [LPPM86]. size [BL87, SS84]. sizes [SS84]. Skin [Woo82, BR80, Bul85, BC85b, CNRC82, DBT89, KPBS80, Mcv87, Mur81, Nog86, PSKS83, RB87, RBRR87]. skipjack [BBN87]. slimy [Fit88]. smelt [AM85, HB83b, MLM88]. smolts [BG84]. smooth [LL85b]. snakehead [NRC^{+}88]. snakeskin [PVdM87]. Snieszko [Wpl80]. sobria [TBRH89]. sockeye [AGA80, BG84, MB83, MP85b]. soils [NAS88]. Solangi [HSO81]. solani [BE81, MRLB89]. Solar [MB87, BC85b, MB87, WSEL88]. sole [CB82, CMRM85, KPBS80]. Solea [CB82, Mcv87]. soluble [EH81, SO82]. Some [AJCM83, LD82, Ela82, LPD85, PN86]. Sound [CMRM85]. sources [FGL82]. South [PWL^{+}89, Rou86, Bal83, BR81b, WGC^{+}87, DKHJ83]. south-east [Bal83]. south-eastern [BR81b, WGC^{+}87]. southeastern [DLN^{+}89]. sp [BE80, BR84b, BS81a, CNRC82, CNR83, Con82, Dia87, Hau84, Her88, HAEE87, HFH85, Lan87, LR86, LFDK89, Mac81, MM80, MGB83, Mod81, MS83c, MC86, MM88, PM84, RLP85, RLF89, Sha82, VA88]. sp.n [Min81]. sp.nov [DLG83, LPD85, RM86]. Spain [APGL86].

Sparisoma [GW83]. Sparus [NP85, PHK80, Pap82, PSC82, Pap83]. spathaceum [Gat87]. spathula [BKUW87, MRS88]. Spawning [AM85, FR80, FTH87, SWM^{+}87]. special [LS83]. species [CN80, CVM80, DKHJ83, EWM83, Gat87, GGM85, HFEJ86, LPD85, LA89, MH83, MS81a, MP89b, PS83, RLRF84, Som81]. specific [CGBF87, HCLW89, Sak84a, Sak85]. specimen [Ell84, HR80]. spermatocoeles [BF85]. spermatocytic [MIT84]. Sphaerospora [HK85, LPD85, Mol88]. sphaerosporosis [Mol80]. Sphyrna [MRLB89]. spill [HNS^{+}82, SHNS83]. spilurus [HR80, HR80]. Spinal [Lan87]. Spinitectus [JC82]. spiny [LL85a, Woo82]. Spirurida [JC82]. spleen [Ell80]. splendens [CP88]. splenic [BWF84]. spondylosis [LK84]. Spontaneous [MIT84, MGJ81, MH83]. spores [HC88]. Sporogenesis [Ste86, DS88, LPSBL82]. sporulating [DLG83]. spot [BR81b, CK89a, CFV89, RB81, SCM^{+}82]. spp [APGL86, BM85b, BNS86, ECM84, KK82, LPSBL82, SRS80, SRH85]. Sprague [MM86]. Spring [FA84, Ahn85, BC85a, FO85, MK89a, SNZH80]. Squalus [LL85a, Woo82]. Squamous [FVO87]. SRBC [GNV87]. ssp [HGB83]. stability [TM82]. stages [LPD85, Mar89, MS89, VA88]. staining [HF80]. standardization [CP88]. starry [PMW80]. state [MA88, MK89b, YMM82]. states [Bog86, DLN^{+}89]. static [Sod85]. static-water [Sod85]. status [HLGC87, LHC^{+}86]. Steatitis [SK85]. steelhead [HKF80, Kni81, RDB84]. Steindachner [TBS82, TBS83]. stellatus [PMW80]. Stephanochasmus [Som81]. Stizostedion [HS82, KNMY83, YKN85]. stock [CT84]. stomach [Pap83]. stone [WTR88]. stored [BM83]. strain [DT81, NB86, OEL85, WC88]. strains [AK85, IIK^{+}84, LD84, OSHF83, PSDS89, Sak85, SCR87, SA84]. Strawberry [OBB^{+}85]. streptococcal [KK82]. Streptococcus [RLP85]. Stress [HKF80, KSDM86]. Stress-induced [HKF80]. Stress-mediated [KSDM86]. striatus [NRC^{+}88]. striped [CC84, MSWL84, SMKH84, SRH85, SCR87, WWK^{+}87]. structure [Cor86]. Studies [CNR80, CKGB^{+}84, HC87, JSF82, KK82, MA88, RLF89, DS88, HAEE87, HL84, HA82, ML89, MB86b, PSZ81, RRB81, SM80, SE80a]. study [BBN87, CHBR87, CRC88, DG82, DTC^{+}84, FFCM86, KOA81, LS83, MB87, Miy87, MB81, MS81a, MS81b, Neu84, PM82, PM84, RDB84, SWM^{+}87, Som81, SW86, SS88, SS89, TH89, TW88, TW89b, TNH88]. Stumpbody [TBS82]. stylirostris [BL87, LR85]. subcellular [BMR^{+}89]. subcutaneous [Wat82]. sublethal [BKF83, SR80]. subsp [REM89]. success [BSB89]. successful [HJ87]. sucker [SFH89]. suckers [Hay80, SZ87]. Sueur [DH84, Nog86]. suffering [EH83, FH85, HAEE87, ML84, WSEL88]. sulphadiazine [KYW83]. sulphate [NAS88]. sulphobromophthalein [HD82]. summer [MB87]. Sunapee [SK85]. Sunburn [BR81a]. superficial [TRT89]. supernatant [NB86]. supersaturation [CC84, KH89]. surface [AM88, CC85, CRC88, Don86, EEC82, HSJE88]. survey [MJ82, SBHM81, SHNS83]. Susceptibility [BKF83, Kni81, PS83, DG82, FHW89, SKS86, WC88, WLL87]. susceptible [MS89]. suspended [GFDM86, GFMD88]. SVCV [Ahn85, FA84]. Swain [FOB85]. Swainson [Els83, EW83]. Swedish [JSF82]. swimbladder [CKGB^{+}84, Mol88]. syndrome [BBN87, LRBB84, MB87]. system [GGL^{+}83, GNV87, Hor87, LdHvM85, PR84, SM80, SKS86, TM85a, TM85b]. Systemic [NNS88, EWM83, NSRS89, PHK80]. Syuzo [San87].

T. [TBS82]. T3 [LS81]. T4 [LS81]. tabacinum [AP85]. Tables [SS84]. tail [TBS82, Woo82]. tail-less [TBS82]. tang [HHC^{+}81]. Tanner [MS80]. Tapes [Els86]. tapeworm [BR84a]. tapeworms [Hay80]. Tasmania [RLF89]. tau [BF85, RGLS81, TB88]. taxonomic [Lev84]. taxonomy [MLG^{+}88]. tayenus [DKHJ83]. TCDD [SSKP88]. technique [AMJS82, ACWM86, CCK83, Lai80, Olu86, WAC88]. techniques [BSRF89, FA84]. teleost [Cou84, HF80, Sec85, WB80]. Teleostei [Dav85, DS88]. teleosts [Lan89]. Telogaster [MS82a]. Temminck [AT86, AKK81, AKF^{+}84, HFEJ86, HFOW88, KMKK83, Miy87, SON81, SLK89]. temperature [ABLY88, BM85b, DT81, Fer81, HG82, Jør82a, Jør82b, TTR85]. temperature-dependent [Jør82b]. temperatures [BM83, HTM^{+}82]. tench [AMS82, HK85, Mol82]. tenuis [MSWL84]. term [HNS^{+}82, HTM^{+}82, SHNS83]. test [Mar89, MJ82, SLK89, SG81]. tested [EdGLV88]. testes [BF85]. testing [Ald82]. testosterone [LS80]. tests [BS81b]. tetrachlorodibenzo [SSKP88]. Tetramicra [MM80]. Thailand [KRP87, NCL87]. their [Eck87, Els86, LM85, SWM^{+}87, TBS82]. Thelohan [PM82, LPD85]. Thelohania [Her88, IKA87, STF^{+}88]. Thelohaniidae [IKA87]. Therapeutic [SKSA^{+}86]. Thersitina [Don86]. three [BM83, MH83, RLRF84, SCM^{+}82, SS84]. three-spot [SCM^{+}82]. Thünberg [Els80, EBF^{+}87]. Thymallus [Eck87, HFSPP88, MW88]. thymic [WGA^{+}84]. Thyroid [LS81, LS80]. thyrsites [SW86]. thyrsitis [HS85, MS86]. tiamulin [BP83]. tiburo [MRLB89]. tiger [NAS88, NNAS88, TB87]. tilapia [KYBN88, LRM^{+}88, Mat84, OR86, HR80, TBS82, TBS83]. tilapias [Olu86, Som82a, Som82b]. time [SWM^{+}87]. times [Jac89]. Tinca [AMS82, HK85, Mol82]. tincae [HK85]. tinea [AMS82]. Tissue [Ben80, DL80, BM83, Dix87, GND^{+}87, Hau84, KNMY83, LFYL81, MM80, RPT^{+}88, Som81, Wat82]. tissues [BA82, CNR81, HB83a, Min81, Olu86, SG81]. Titre [MPJ83b]. titres [THJS87]. toadfish [BF85, RGLS81, TB88]. Tolerance [KH89]. tomcod [Cor86, CRS^{+}89, Cor86, CRS^{+}89]. total [FCM88, LS81]. Toxicity [MP80, NB86, DF84, PRSC85, Sak84a]. tract [WTR88]. transferable [AK85]. Transmission [PN86, Sak87, WWK^{+}87, DMdK84, HKF80, Lan89, MPJ83a, MP85b]. trapping [Ell80]. Travin [BB89]. treated [EPPK86]. Treatment [CNRC82, Jac89, KYBN88, Mcv87, SKS86]. Trematoda [AC80]. Tremblay [DK82]. Trewavas [KYBN88]. Triaenophoridae [BS84a]. Triaenophorus [BS84a, DR82, SD86]. Triangula [Lan87]. Tribrissen(R) [BP83]. tributaries [HM82]. Trichogaster [PVdM87, SCM^{+}82]. trichopterus [SCM^{+}82]. tricolor [GW85]. tridentata [BT86]. trifluralin [Cou84]. trimethoprim [KYW83]. trimetoprim [Jac89]. Trinectes [SO82]. Tripathi [Dav85, DS88]. trout [ALWH87, AMJS82, ARW83, AHW83, Ahn83, AURE80, Aus85, AM88, BKF83, BL86, BSB89, Ber84, BA82, Bru86a, BM86a, Bru86b, BR80, BC85b, BS81b, CT82, CT84, CHRB87, CHBR87, Con82, CD86, DMdK84, DF84, DF85, DF86, Dix87, DT81, DARD87, EH83, ENH86, EMM85, FB87, FR80, Fer81, FSC85, GD80, GC82, GFDM86, GFMD88, Gri89, HC88, HE88, Her85, HG84, HHF84, HD82, HFH85, HL84, HTM^{+}82, HRTW84, HKF80, Jac89, JSF82, JA83a, JWS86, JW87, Jør82a, Jør82b, KNDLR86, KNBLR86, Kni81, KH89, LHW88, LS80, LLGV83, Lil89, MJG84, MA88, ML89, MR84, MCA84, Mey83, MF82, MHSF84, MS83c, Neu84, NM82, OBB^{+}85, PR85, PRSC85, PW85, Poy87, Rob83, RPT^{+}88, RDB84, SK85, Sak84a, SOAK89, SDC89, SD80, SRS80, Sod85]. trout [SK87, SSKP88, SCG82, TM85a, TM85b, TW88, THJS87, WGA^{+}84, WS81, WWB86, YMM82]. trutta [JSF82, BSB89, BA82, DARD87, EMM85, FR80, KNDLR86, KNBLR86, Poy87, WWB86]. tsetse [MR82]. tshawytscha [MM85, BKF83, FTH87, Hau84, MK89a]. tubercle [BM80]. tuberculosis [MGJ81, MG83]. Tubifex [HC87, WMH86]. tubules [MP89b]. tumour [BM80]. tumours [EJL81, KPBS80, TM88]. tuna [BBN87]. tunicate [DH84]. turbinata [DH84]. turbot [BL86, DBT89, Kir80, MM80]. Turkey [BS89]. Turnbull [Sco85]. Tweed [HM82]. Two [BSG83, Hay80, Els86, FGL82, MS86, ML84, SO82, TBS82, TM82, Wen87]. type [CN80, CK89b, EPB85a, EPB85b, FGL82, Wat82, YKN85]. type-C [YKN85]. typicalis [CN80]. tyrannus [NND89, ND86, SNZH80].

U.K. [MJG84]. U.S.A. [MGB83, PMW80, TH80]. UDN [JSF82]. ulcer [CH88, WGC^{+}87]. Ulcerative [DBT89, DLN^{+}89, FHW89, ES80b, ES80a, JSF82, ML84, ND86, STEM80]. Ultrastructural [BGB82, BM85a, DPS88, Els80, FB87, HSO81, HA82, KOA81, LS83, PSZ81, RBRR87, TM88, VA88, CN80, DS88, Lom80, PM82, PM84, SW86, TH89, TNH88]. Ultrastructure [Dia87, HSO83a, HSO83b, IKA87, RN81b]. Ultraviolet [GGL^{+}83, Bul85, BC85b, MB87, PW85]. ultraviolet-B [Bul85, PW85]. un-ionized [Sod85]. undecimalis [Urd82]. underyearling [MCSZ81]. undulatus [EH81]. unfertilized [SWM^{+}87]. Unicauda [Min81]. unicellular [Lan86]. unidentified [DK82, Wat82]. Unilateral [ECM84]. unique [DF85]. United [DLN^{+}89, HBF81, SE80b]. unusual [TH89]. upon [Bul85, BC85b, CHA87]. upper [Hor87]. uptake [LS81]. uracil [SLMH86]. uridine [SLMH86]. Urine [KN89]. Uronema [CNR80]. Uronematidae [CNR80]. USA [HSR^{+}88, Hor87, KYBN88, MHBM86, Nog86, TB88]. Use [ACWM86, MKS83, WAC88, WAC89, CVM80, DH83, HC88, HRTW84, LM85, Mur81, WB85]. used [SAAK87b, SAAK87a]. using [JA83a, MS80, PK88].

vaccinated [JFA82a, JFA82b, Miy87]. Vaccination [HTM^{+}82, AGA80, AKF^{+}84, CD86, HJ87, JA84, LdHvM85, NM82, Sak85, THJS87]. Vaccine [EdGLV88, DH83]. Vaccine-induced [EdGLV88]. vaccines [HRTW84, Lil89, OEL85]. Valenciennes [DKHJ83, GW83, HFH85, MHBM86]. valenciennesi [FBV88]. Vaney [MLG^{+}88]. vannamei [DLF88, LR86]. variation [RB81, SA84]. variations [BWF84]. various [AJ84, Bog86, Gat87, SS84]. vectors [Ahn85, CC86, Wil89b]. Vefsna [JJ88]. VEN [RN81b]. Venezuela [Ber80, Urd82]. ventral [LRBB84]. Vertebral [Poy87, TBS82]. vertebrates [DPS88]. Vertical [MP85b]. vetulus [CMRM85, KPBS80]. VHS [AURE80, MW88]. Vibrio [AJ84, AGA80, AKK81, BKF83, BR84b, EEC82, EHHJ88, GGM85, GBC^{+}89, HSJE88, HAEE87, Hof89, HB83a, JFA82a, JFA82b, JA83b, KAF^{+}83, Miy87, NB86, NB87a, NB87b, RLRF84, SRH85]. vibrios [MB86a, NB86, NB87a]. Vibriosis [GC82, Lew85, AHW83, AKF^{+}84, EACR81, EW83, EHHJ88, HJ87, HRTW84, Lil89, Miy87, THJS87, TNH88, WSEL88]. viraemia [Ahn85, BC85a, FA84]. Viral [MW88, AURE80, BGB82, CC86, DG82, EH83, ENH86, EPB85a, EPB85b, GD80, LHC^{+}86, LH87, LL85b, MHBM86, Neu84, PAPBJ89, RN81b, SON81, SNZH80, TB87]. virazole [SD80]. Virginia [TB88]. virginica [EEC82, Mey80]. virginicus [CNR81]. viridans [Gje84]. Virulence [CH88, EBS88, Jør82b, TJGA86]. virulent [HG82, Jør82a, LD84, OEL85]. Virus [SE80b, SE80a, AF81, AMJS82, ARW83, Ahn83, Ahn85, AURE80, BGB82, BM80, BC85a, BMJ^{+}85, BP88, BL85, BM83, BG84, CPH88, DH83, Dix87, DT81, FA84, FGL82, FTH87, HL85, HCLW89, HBF81, IIK^{+}84, Jør82a, Jør82b, KNMY83, LHW^{+}86, Lan89, LFYL81, MA88, MR84, MS88, MS89, MLM88, MPJ83a, MPJ83b, MP85b, NNAS88, Neu84, OSHF83, OJBM88, RN81b, RPT^{+}88, RPL89, SD80, SMKH84, SSKP88, SG81, SCG82, WWK^{+}87, WB85, WBB85, YMM82]. viruses [LHC^{+}86, LM85, PN86, TM82]. Visceral [FODS89, HFEJ86, MGB83]. visualization [AURE80]. vitamin [AM83, MHSF84]. vitreum [KNMY83, YKN85, HS82, KNMY83]. vitro [Ald82, Bai83, BA82, BNBKS81, CGBF87, KYW83, LPSBL82, LPBLS85, Sec85, TW88, TW89b]. vivo [BA82, ENH86, TW88]. Vladykov [DK82].

W [Bau80]. Walbaum [AGA80, BKF83, BKUW87, BG84, CW84, Cor86, CRS^{+}89, CC84, EPB85a, EPB85b, EPPK86, FTH87, Hau84, KH89, LS81, MK89a, MM85, MRS88, MSWL84, MB83, MP85b, NM82, RR89, SMKH84, SK87, SRH85, SCR87, Tra86, WWK^{+}87]. Wales [Dav85, Rou86, SBHM81]. walking [KRP87]. walleye [HS82, KNMY83, YKN85]. warm [Ber80]. Washington [PMW80, CMRM85]. water [Ber80, CVM80, DT81, EACR81, EH81, EKPP84, EPPK86, Fer81, GD80, GC82, KNDLR86, KNBLR86, MPJ83a, Neu84, REM89, Sod85, SO82, TNH88]. water-borne [Neu84]. water-hardened [EKPP84]. water-hardening [EPPK86]. water-soluble [EH81, SO82]. waterborne [MPJ83a]. waters [BM88, RL87, SE80b, TM82]. weight [HD82]. west [APGL86, CW84, EPB85a, EPB85b, MC86, FBS83, TH80]. whirling [Hor87, Mar89, WMH86]. Whitaker [MS86, SW86, Ste86]. white [AMS82, Hay80, SZ87, SFH89]. whitefish [DR82, RVF86]. whiting [Ela82, Mac81, MS86, SW86, Ste86]. whole [AHW83]. wide [DF89]. wild [BR82, BVN80, Cop81, Cop82, Cop83, Dia87, FGL82, HFSPP88, TT84, Wil89b]. wilsoni [RN81a]. winter [CW84, EPB85a, EPB85b]. Withdrawal [Jac89]. within [SWM^{+}87]. Wolf [SM89]. Woodland [KMF89, WB80]. Wrac’h [HNS^{+}82, SHNS83].

X [LRM^{+}88, PSKS83]. X-cells [PSKS83].

Yamaguti [ACCD81, KYBN88]. year [TH80, Tra86]. yellow [BG85]. yellowfin [Rou86]. yellowtails [KK82]. Yersinia [AJCM83, AJ84, CGBF87, CD86, DF89, DSK87, FSC85, HTV88, HKF80, JFA82a, JFA82b, JA83a, JA83b, Kni81, MJ82, NM82, RVF86, Rob83, SCR87, SA84, Wil89b]. yessoensis [Els86]. York [TB88]. young [TH80]. young-of-the-year [TH80]. Yu [SR89].

zebra [De 80]. Zschokkella [Dav85, DS88].

References

Austin:1986:MAB

[ABG^{+}86] B. Austin, I. Bishop, C. Gray, B. Watt, and J. Dawes. Monoclonal antibody-based enzyme-linked immunosorbent assays for the rapid diagnosis of clinical cases of enteric redmouth and furunculosis in fish farms. Journal of Fish Diseases, 9(5):469–474, September 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Angelidis:1988:ETC

[ABLY88] P. Angelidis, F. Baudin-Laurencin, and P. Youinou. Effects of temperature on chemiluminescence of phagocytes from sea bass, Dicentrarchus labrax L. Journal of Fish Diseases, 11(4):281–288, July 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Andrews:1980:ODB

[AC80] C. Andrews and J. C. Chubb. Observations on the development of Bunodera luciopercae (Müller, 1776) (Trematoda: Allocreadiidae) under field and laboratory conditions. Journal of Fish Diseases, 3 (6):481–493, November 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Andrews:1981:OBA

[ACCD81] C. Andrews, J. C. Chubb, T. Coles, and A. Dearsley. The occurrence of Bothriocephalus acheilognathi Yamaguti, 1934 (B. gowkongensis) (Cestoda: Pseudophyllidea) in the British Isles. Journal of Fish Diseases, 4(1):89–93, January 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ainsworth:1986:UMA

[ACWM86] A. J. Ainsworth, G. Capley, P. Waterstreet, and D. Munson. Use of monoclonal antibodies in the indirect fluorescent antibody technique (IFA) for the diagnosis of Edwardsiella ictaluri. Journal of Fish Diseases, 9(5):439–444, September 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ahmed:1980:DFG

[AE80] A. T. A. Ahmed and S. Egusa. Dermal fibrosarcoma in goldfish Carassius auratus (L.). Journal of Fish Diseases, 3(3):249–254, May 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Adair:1981:IIP

[AF81] B. M. Adair and H. W. Ferguson. Isolation of infectious pancreatic necrosis (IPN) virus from non-salmonid fish. Journal of Fish Diseases, 4(1):69–76, January 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Alderman:1986:PNC

[AFP86] D. J. Alderman, S. W. Feist, and Jane L. Polglase. Possible nocardiosis of crayfish, Austropotamobius pallipes. Journal of Fish Diseases, 9(4):345–347, July 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Arakawa:1986:SMC

[AFS86] C. K. Arakawa, J. L. Fryer, and J. E. Sanders. Serology of Mycobacterium chelonei isolated from salmonid fish. Journal of Fish Diseases, 9(3):269–271, May 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Antipa:1980:VAV

[AGA80] R. Antipa, R. Gould, and D. F. Amend. Vibrio anguillarum vaccination of sockeye salmon Oncorhynchus nerka (Walbaum) by direct and hyperosmotic immersion. Journal of Fish Diseases, 3(2): 161–165, March 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ahne:1983:PIP

[Ahn83] W. Ahne. Presence of infectious pancreatic necrosis virus in the seminal fluid of rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 6(4):377, July 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ahne:1985:AFP

[Ahn85] W. Ahne. Argulus foliaceus L. and Piscicola geometra L. as mechanical vectors of spring viraemia of carp virus (SVCV). Journal of Fish Diseases, 8(2):241–242, March 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ahne:1986:BPD

[Ahn86] W. Ahne. Biography: Professor Dr Heinz-Hermann Reichenbach-Klinke. Journal of Fish Diseases, 9(5):365–366, September 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Agius:1983:IRT

[AHW83] C. Agius, M. T. Horne, and P. D. Ward. Immunization of rainbow trout, Salmo gairdneri Richardson, against vibriosis: comparison of an extract antigen with whole cell bacterins by oral and intraperitoneal routes. Journal of Fish Diseases, 6(2):129–134, March 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Amend:1984:ELA

[AJ84] D. F. Amend and K. A. Johnson. Evidence for lack of antigenic competition among various combinations of Vibrio anguillarum, Yersinia ruckeri, Aeromonas salmonicida and Renibacterium salmoninarum bacterins when administered to salmonid fishes. Journal of Fish Diseases, 7(4):293–299, July 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Amend:1983:SFA

[AJCM83] D. F. Amend, K. A. Johnson, T. R. Croy, and D. H. Mccarthy. Some factors affecting the potency of Yersinia ruckeri bacterins. Journal of Fish Diseases, 6(4):337–344, July 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Aoki:1985:DTR

[AK85] T. Aoki and T. Kitao. Detection of transferable R plasmids in strains of the fish–pathogenic bacterium, Pasteurella piscicida. Journal of Fish Diseases, 8(4):345–350, July 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Aoki:1984:MHI

[AKF^{+}84] T. Aoki, T. Kitao, M. Fukudome, S. Takahashi, and S. Egusa. Modification of the hyperosmotic infiltration method of vaccination against vibriosis in cultured ayu, Plecoglossus altivelis Temminck and Schlegel. Journal of Fish Diseases, 7(2):149–156, March 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Aoki:1981:CDR

[AKK81] T. Aoki, T. Kitao, and K. Kawano. Changes in drug resistance of Vibrio anguillarum in cultured ayu, Plecoglossus altivelis Temminck and Schlegel, in Japan. Journal of Fish Diseases, 4(3): 223–230, May 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anderson:1987:PDG

[AL87] W. I. Anderson and Pamela B. Luther. Poorly differentiated granuloplastic leukaemia in a bowfin, Amia calva L. Journal of Fish Diseases, 10(5):411–413, September 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Alderman:1982:VTF

[Ald82] D. J. Alderman. In vitro testing of fisheries chemotherapeutants. Journal of Fish Diseases, 5(2):113–123, March 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Alderman:1985:MGR

[Ald85] D. J. Alderman. Malachite green: a review. Journal of Fish Diseases, 8(3):289–298, May 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Adams:1987:BCM

[ALWH87] Alexandra Adams, W. Leschen, Anne Wilson, and M. T. Horne. A bath challenge model for furunculosis in rainbow trout, Salmo gairdneri Richardson, and Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 10(6):495–504, November 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Agrawal:1980:NDD

[AM80] N. K. Agrawal and C. L. Mahajan. Nutritional deficiency disease in an Indian major carp, Cirrhina mrigala Hamilton, due to avitaminosis C during early growth. Journal of Fish Diseases, 3(3): 231–248, May 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Agrawal:1983:PVB

[AM83] N. K. Agrawal and C. L. Mahajan. Pathology of vitamin B-6 deficiency in Channa (= Ophiocephalus) punctatus Bloch. Journal of Fish Diseases, 6(5):439–450, September 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anders:1985:SPS

[AM85] Kerstin Anders and H. Möller. Spawning papillomatosis of smelt, Osmerus eperlanus L., from the Elbe estuary. Journal of Fish Diseases, 8(2):233–235, March 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Austin:1988:NAC

[AM88] B. Austin and D. Mcintosh. Natural antibacterial compounds on the surface of rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 11(3):275–277, May 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Agius:1982:MST

[AMJS82] C. Agius, H. Mangunwiryo, R. H. Johnson, and D. A. Smail. A more sensitive technique for isolating infectious pancreatic necrosis virus from asymptomatic carrier rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 5(4):285–292, July 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ahne:1982:IPF

[AMS82] W. Ahne, H. Mahnel, and P. Steinhagen. Isolation of pike fry rhabdovirus from tench, Tinca tinea L., and white bream, Blicca bjoerkna (L.). Journal of Fish Diseases, 5(6):535–537, November 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1980:BRa

[Ano80a] Anonymous. Book reviews. Journal of Fish Diseases, 3(1):75–76, January 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1980:BRb

[Ano80b] Anonymous. Book reviews. Journal of Fish Diseases, 3(2):177–179, March 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1980:BRc

[Ano80c] Anonymous. Book reviews. Journal of Fish Diseases, 3(3):261–263, May 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1980:BRd

[Ano80d] Anonymous. Book reviews. Journal of Fish Diseases, 3(4):359–361, July 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1980:BRe

[Ano80e] Anonymous. Book reviews. Journal of Fish Diseases, 3(5):447– 449, September 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1980:BRf

[Ano80f] Anonymous. Book reviews. Journal of Fish Diseases, 3(6):533– 535, November 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1981:BRc

[Ano81a] Anonymous. Book Review. Journal of Fish Diseases, 4(4):359, July 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1981:BRe

[Ano81b] Anonymous. Book Review. Journal of Fish Diseases, 4(6):533– 534, November 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1981:BRa

[Ano81c] Anonymous. Book reviews. Journal of Fish Diseases, 4(1):95–97, January 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1981:BRb

[Ano81d] Anonymous. Book reviews. Journal of Fish Diseases, 4(2):185–186, March 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1981:BRd

[Ano81e] Anonymous. Book reviews. Journal of Fish Diseases, 4(5):441, September 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1982:BRa

[Ano82a] Anonymous. Book Review. Journal of Fish Diseases, 5(2):171, March 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1982:BRb

[Ano82b] Anonymous. Book Review. Journal of Fish Diseases, 5(3):263, May 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1982:BRd

[Ano82c] Anonymous. Book Review. Journal of Fish Diseases, 5(5):445– 446, September 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1982:BRc

[Ano82d] Anonymous. Book reviews. Journal of Fish Diseases, 5(4):349–350, July 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1982:BRe

[Ano82e] Anonymous. Book reviews. Journal of Fish Diseases, 5(6):553– 554, November 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1983:A

[Ano83a] Anonymous. Addendum. Journal of Fish Diseases, 6(3):260, May 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1983:BRa

[Ano83b] Anonymous. Book reviews. Journal of Fish Diseases, 6(4):397–398, July 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1983:BRb

[Ano83c] Anonymous. Book reviews. Journal of Fish Diseases, 6(6):553– 554, November 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1984:BRa

[Ano84a] Anonymous. Book Review. Journal of Fish Diseases, 7(4):327, July 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1984:BRb

[Ano84b] Anonymous. Book reviews. Journal of Fish Diseases, 7(6):525– 527, November 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1985:BR

[Ano85a] Anonymous. Book reviews. Journal of Fish Diseases, 8(3):337–338, May 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1985:E

[Ano85b] Anonymous. ERRATUM. Journal of Fish Diseases, 8(5):484, September 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1986:BRa

[Ano86a] Anonymous. Book Review. Journal of Fish Diseases, 9(2):177, March 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1986:BRb

[Ano86b] Anonymous. Book reviews. Journal of Fish Diseases, 9(3):287–290, May 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1986:BRc

[Ano86c] Anonymous. Book reviews. Journal of Fish Diseases, 9(5):475– 477, September 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1986:BRd

[Ano86d] Anonymous. Book reviews. Journal of Fish Diseases, 9(6):561, November 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1986:E

[Ano86e] Anonymous. Erratum. Journal of Fish Diseases, 9(5):478, September 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1987:BRa

[Ano87a] Anonymous. Book Review. Journal of Fish Diseases, 10(2):143, March 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1987:BRb

[Ano87b] Anonymous. Book reviews. Journal of Fish Diseases, 10(5):423– 424, September 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1988:BR

[Ano88a] Anonymous. Book Review. Journal of Fish Diseases, 11(3):279, May 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1988:E

[Ano88b] Anonymous. Erratum. Journal of Fish Diseases, 11(4):369, July 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1989:BRa

[Ano89a] Anonymous. Book Review. Journal of Fish Diseases, 12(2):203, March 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1989:BRb

[Ano89b] Anonymous. Book Review. Journal of Fish Diseases, 12(3):279, May 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1989:BRd

[Ano89c] Anonymous. Book Review. Journal of Fish Diseases, 12(6):599, November 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Anonymous:1989:BRc

[Ano89d] Anonymous. Book reviews. Journal of Fish Diseases, 12(4):399– 400, July 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Alderman:1985:FTB

[AP85] D. J. Alderman and Jane L. Polglase. Fusarium tabacinum (Beyma) Gams, as a gill parasite in the crayfish, Austropotamobius pallipes Lereboullet. Journal of Fish Diseases, 8(2):249–252, March 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Alderman:1986:AAI

[AP86] D. J. Alderman and Jane L. Polglase. Aphanomyces astaci: isolation and culture. Journal of Fish Diseases, 9(5):367–379, September 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Alderman:1987:AAP

[APF87] D. J. Alderman, Jane L. Polglase, and M. Frayling. Aphanomyces astaci pathogenicity under laboratory and field conditions. Journal of Fish Diseases, 10(5):385–393, September 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Alderman:1984:CPB

[APFH84] D. J. Alderman, Jane L. Polglase, M. Frayling, and J. Hogger. Crayfish plague in Britain. Journal of Fish Diseases, 7(5):401– 405, September 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Alvarez-Pellitero:1986:ESC

[APGL86] Maria-Pilar Alvarez-Pellitero and Maria-Camino Gonzalez-Lanza. Eimeria spp. from cyprinid fish of the Duero basin (north-west Spain). Journal of Fish Diseases, 9(4):325–336, July 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Austin:1985:ERS

[AR85] B. Austin and J. N. Rayment. Epizootiology of Renibacterium salmoninarum, the causal agent of bacterial kidney disease in salmonid fish. Journal of Fish Diseases, 8(6):505–509, November 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Athanassopoulou-Raptopoulou:1986:ESP

[ARV86] F. Athanassopoulou-Raptopoulou and J. Vlemmas. Eimeria scardinii Pellérdy and Molnár, 1968 in the kidneys of Rutilus rutilus (L.), and Scardinius erythrophthalmus (L.) from northern Greece. Journal of Fish Diseases, 9(5):411–416, September 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Agius:1983:FOC

[ARW83] C. Agius, A. Richardson, and W. Walker. Further observations on the co-cultivation method for isolating infectious pancreatic necrosis virus from asymptomatic carrier rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 6(5):477–480, September 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Arakawa:1987:PMA

[ASF87] C. K. Arakawa, J. E. Sanders, and J. L. Fryer. Production of monoclonal antibodies against Renibacterium salmoninarum. Journal of Fish Diseases, 10(3):249–253, May 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ahne:1986:IPF

[AT86] W. Ahne and I. Thomsen. Isolation of pike fry rhabdovirus from Pseudorasbora parva (Temminck & Schlegel). Journal of Fish Diseases, 9(6):555–556, November 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Amlacher:1980:DEM

[AURE80] E. Amlacher, J. Ude, Christiane Rudolph, and G. Ernst. Direct electron microscopical visualization of the presumptive virus of viral haemorrhagic septicaemia (VHS) in the rainbow trout Salmo gairdneri Richardson and additional histopathological and haematological observations. Journal of Fish Diseases, 3(1):55–62, January 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Austin:1985:EAC

[Aus85] B. Austin. Evaluation of antimicrobial compounds for the control of bacterial kidney disease in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 8(2):209–220, March 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bowers:1982:VVP

[BA82] A. Bowers and J. B. Alexander. In vitro and in vivo passage of bacteria across restricted areas and isolated tissues of trout, Salmo trutta L. and S. gairdneri Richardson. Journal of Fish Diseases, 5(2):145–151, March 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bailey:1983:MVS

[Bai83] T. A. Bailey. Method for in vitro screening of aquatic fungicides. Journal of Fish Diseases, 6(2):91–100, March 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ball:1983:EML

[Bal83] S. J. Ball. Eimeria metchnikovi (Laveran, 1897) Reichenow, 1921 from the gudgeon, Gobio gobio L., in south-east England. Journal of Fish Diseases, 6(2):201–203, March 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bell:1987:EFR

[BAL87] T. A. Bell, C. S. Arume, and D. V. Lightner. Efficacy of formalin in reducing the levels of peritrichous ciliates on cultured marine shrimp. Journal of Fish Diseases, 10(1):45–51, January 1987. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bauer:1980:PWS

[Bau80] O. N. Bauer. Professor W. Schaperclaus. Journal of Fish Diseases, 3(2):79–80, March 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bogovski:1989:CRP

[BB89] S. P. Bogovski and Y. I. Bakai. Chromatoblastomas and related pigmented lesions in deepwater redfish, Sebastes mentella (Travin), from North Atlantic areas, especially the Irminger Sea. Journal of Fish Diseases, 12(1):1–13, January 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bourke:1987:SDC

[BBN87] R. E. Bourke, J. Brock, and R. M. Nakamura. A study of delayed capture mortality syndrome in skipjack tuna, Katsuwonus pelamis (1). Journal of Fish Diseases, 10(4):275–287, July 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bekesi:1985:ISV

[BC85a] L. Békési and L. Csontos. Isolation of spring viraemia of carp virus from asymptomatic broodstock carp, Cyprinus carpio L. Journal of Fish Diseases, 8(5):471–472, September 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bullock:1985:ISU

[BC85b] A. M. Bullock and R. R. Coutts. The impact of solar ultraviolet radiation upon the skin of rainbow trout, Salmo gairdneri Richardson, farmed at high altitude in Bolivia. Journal of Fish Diseases, 8 (3):263–272, May 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bast:1988:EPA

[BDDS88] L. Bast, J. G. Daly, S. A. Degrandis, and R. M. W. Stevenson. Evaluation of profiles of Aeromonas salmonicida as epidemiological markers of furunculosis infections in fish. Journal of Fish Diseases, 11(2):133–145, March 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bian:1980:AHS

[BE80] B. Z. Bian and S. Egusa. Atkinsiella hamanaensis sp. nov. isolated from cultivated ova of the mangrove crab, Scylla serrata (Forsskål).

Journal of Fish Diseases, 3(5):373–385, September 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bian:1981:HBG

[BE81] B. Z. Bian and S. Egusa. Histopathology of black gill disease caused by Fusarium solani (Martius) infection in the kuruma prawn, Penaeus japonicus Bate. Journal of Fish Diseases, 4(3):195–201, May 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bruno:1986:MHC

[BE86] D. W. Bruno and A. E. Ellis. Multiple hepatic cysts in farmed Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 9(1): 79–81, January 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Benz:1980:TPA

[Ben80] G. W. Benz. Tissue proliferations associated with Nemesis lamna Risso, 1826 (Copepoda: Eudactylinidae) infestations on the gill filaments of shortfin makos (Isurus oxyrinchus Rafinesque). Journal of Fish Diseases, 3(5):443–446, September 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bermudez:1980:PEC

[Ber80] D. Bermudez. Preliminary experiences in the control of fish diseases in warm water aquaculture operations in Venezuela. Journal of Fish Diseases, 3(4):355–357, July 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bernstein:1984:LLH

[Ber84] J. W. Bernstein. Leukaemic lymphosarcoma in a hatchery-reared rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 7(1):83–86, January 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Burns:1985:BMS

[BF85] J. C. Burns and M. L. Fine. Bilateral multicystic spermatocoeles in the testes of a toadfish, Opsanus tau (L.). Journal of Fish Diseases, 8(1):143–144, January 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Burke:1984:ECI

[BG84] J. Burke and R. Grischkowsky. An epizootic caused by infectious haematopoietic necrosis virus in an enhanced population of sockeye salmon, Oncorhynchus nerka (Walbaum), smolts at Hidden Creek, Alaska. Journal of Fish Diseases, 7(5):421–429, September 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Brown:1985:ADA

[BG85] C. L. Brown and C. J. George. Age-dependent accumulation of macrophage aggregates in the yellow perch, Perca flavescens (Mitchill). Journal of Fish Diseases, 8(1):135–138, January 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Baroni:1982:UOV

[BGB82] A. Baroni, R. Galesso, and G. Bovo. Ultrastructural observations on the virus of viral haemorrhagic septicaemia in culture. Journal of Fish Diseases, 5(5):439–444, September 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Brenden:1986:PEA

[BH86] R. A. Brenden and H. W. Huizinga. Pathophysiology of experimental Aeromonas hydrophila infection in goldfish, Carassius auratus (L.). Journal of Fish Diseases, 9(2):163–167, March 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bauer:1984:B

[BHS84] O. N. Bauer, G. L. Hoffman, and Yu. A. Strelkov. Biography. Journal of Fish Diseases, 7(6):431–432, November 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Baba:1988:CMP

[BIII88a] T. Baba, J. Imamura, K. Izawa, and K. Iked. Cell-mediated protection in carp, Cyprinus carpio L., against Aeromonas hydrophila. Journal of Fish Diseases, 11(2):171–178, March 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Baba:1988:IPC

[BIII88b] T. Baba, J. Imamura, K. Izawa, and K. Ikeda. Immune protection in carp, Cyprinus carpio L., after immunization with Aeromonas hydrophila crude lipopolysaccharide. Journal of Fish Diseases, 11 (3):237–244, May 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Baker:1983:SCS

[BKF83] R. J. Baker, M. D. Knittel, and J. L. Fryer. Susceptibility of Chinook salmon, Oncorhynchus tshawytscha (Walbaum), and rainbow trout, Salmo gairdneri Richardson, to infection with Vibrio anguillarum following sublethal copper exposure. Journal of Fish Diseases, 6(3):267–275, May 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bean-Knudsen:1987:CCP

[BKUW87] D. E. Bean-Knudsen, L. S. Uhazy, and J. E. Wagner. Cranial chondrosarcoma in a paddlefish, Polyodon spathula (Walbaum). Journal of Fish Diseases, 10(5):363–369, September 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Buck:1985:LOP

[BL85] C. D. Buck and P. C. Loh. Liquid overlay plaquing of channel catfish virus. Journal of Fish Diseases, 8(3):325–328, May 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Balouet:1986:GNF

[BL86] G. Balouet and F. Baudin Laurencin. Granulomatous nodules in fish: an experimental assessment in rainbow trout, Salmo gairdneri Richardson, and turbot, Scophthalmus maximus (L.). Journal of Fish Diseases, 9(5):417–429, September 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bell:1987:IDP

[BL87] T. A. Bell and D. V. Lightner. IHHN disease of Penaeus stylirostris: effects of shrimp size on disease expression. Journal of Fish Diseases, 10(3):165–170, May 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Banner:1986:OSF

[BLFR86] C. R. Banner, J. J. Long, J. L. Fryer, and J. S. Rohovec. Occurrence of salmonid fish infected with Renibacterium salmoninarum in the Pacific Ocean. Journal of Fish Diseases, 9(3):273–275, May 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Beckwith:1980:KTV

[BM80] D. G. Beckwith and R. G. Malsberger. Kidney tumour virus — tumour or mycobacterial tubercle? Journal of Fish Diseases, 3 (4):339–348, July 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Burke:1983:RIH

[BM83] J. Burke and D. Mulcahy. Retention of infectious haematopoietic necrosis virus infectivity in fish tissue homogenates and fluids stored at three temperatures. Journal of Fish Diseases, 6(6):543– 547, November 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bodammer:1985:UOH

[BM85a] J. E. Bodammer and Sharon A. Maclean. Ultrastructural observations on the haemoparasite Haematractidium scombri in the Atlantic mackerel, Scomber scombrus L. Journal of Fish Diseases, 8(2):243–247, March 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bower:1985:ETS

[BM85b] Susan M. Bower and L. Margolis. Effects of temperature and salinity on the course of infection with the haemoflagellate Cryptobia salmositica in juvenile Pacific salmon, Oncorhynchus spp. Journal of Fish Diseases, 8(1):25–33, January 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bruno:1986:HAR

[BM86a] D. W. Bruno and A. L. S. Munro. Haematological assessment of rainbow trout, Salmo gairdneri Richardson, and Atlantic salmon, Salmo salar L., infected with Renibacterium salmoninarum. Journal of Fish Diseases, 9(3):195–204, May 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bullock:1986:PPM

[BM86b] A. M. Bullock and G. Minkoff. Prevalence and pathology of mesenchymal growths on the head of laboratory-reared herring larvae, Clupea harengus L. Journal of Fish Diseases, 9(6):551–554, November 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bruno:1988:IPL

[BM88] D. W. Bruno and A. H. Mcvicar. Investigations into pseudobranch lesions of haddock, Melanogrammus aeglefinus L., caught in Scottish waters. Journal of Fish Diseases, 11(2):191–196, March 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bowser:1985:ICC

[BMJ^{+}85] P. R. Bowser, A. D. Munson, H. H. Jarboe, R. Francis-Floyd, and P. R. Waterstrat. Isolation of channel catfish virus from channel catfish, Ictalurus punctatus (Rafinesque), broodstock. Journal of Fish Diseases, 8(6):557–561, November 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bowser:1985:LDC

[BMM85] P. R. Bowser, C. P. Mccoy, and J. R. MacMillan. A lymphoproliferative disorder in a channel catfish, Ictalurus punctatus (Rafinesque). Journal of Fish Diseases, 8(5):465–469, September 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bloch:1986:ALP

[BMN86] B. Bloch, S. Mellergaard, and E. Nielsen. Adenovirus-like particles associated with epithelial hyperplasias in dab, Limanda limanda (L.). Journal of Fish Diseases, 9(3):281–285, May 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Boulo:1989:IBO

[BMR^{+}89] V. Boulo, E. Mialhe, H. Rogier, F. Paolucci, and H. Grizel. Immunodiagnosis of Bonamia ostreae (Ascetospora) infection of Ostrea edulis L. and subcellular identification of epitopes by monoclonal antibodies. Journal of Fish Diseases, 12(3):257–262, May 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Braun-Nesje:1981:SMS

[BNBKS81] Rosemarie Braun-Nesje, Kjell Bertheussen, Gilla Kaplan, and Rolf Seljelid. Salmonid macrophages: separation, in vitro culture and characterization. Journal of Fish Diseases, 4(2):141–151, March 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Brock:1986:HRI

[BNH^{+}86] J. A. Brock, Lauren K. Nakagawa, T. Hayashi, Stephanie Teruya, and Hana Van Campen. Hepatopancreatic rickettsial infection of the penaeid shrimp, Penaeus marginatus (Randall), from Hawaii. Journal of Fish Diseases, 9(1):73–77, January 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Brock:1986:ICF

[BNS86] J. A. Brock, Lauren K. Nakagawa, and R. J. Shimojo. Infection of a cultured freshwater prawn, Macrobrachium rosenbergii de Man (Crustacea: Decapoda), by Mycobacterium spp., Runyon Group II. Journal of Fish Diseases, 9(4):319–324, July 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Brock:1986:RBP

[BNV^{+}86] J. A. Brock, Lauren K. Nakagawa, Hana Van Campen, T. Hayashi, and Stephanie Teruta. A record of Baculovirus penaei from Penaeus marginatus Randall in Hawaii. Journal of Fish Diseases, 9 (4):353–355, July 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Boghen:1986:SJL

[Bog86] A. D. Boghen. Sensitivity of juvenile lobsters, Homarus americanus L., to different concentrations of malachite green during various states of the intermoult cycle. Journal of Fish Diseases, 9(3):243– 248, May 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bosse:1983:TTC

[BP83] Marie P. Bosse and G. Post. Tribrissen(R) and tiamulin for control of enteric redmouth disease. Journal of Fish Diseases, 6(1):27–32, January 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Brady:1988:SCG

[BP88] Y. J. Brady and J. A. Plumb. Serological comparison of golden shiner virus, chum salmon virus, reovirus 13P2 and catfish reovirus. Journal of Fish Diseases, 11(5):441–443, September 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bullock:1980:IEM

[BR80] A. M. Bullock and R. J. Roberts. Inhibition of epidermal migration in the skin of rainbow trout Salmo gairdneri Richardson in the presence of achromogenic Aeromonas salmonicida. Journal of Fish Diseases, 3(6):517–524, November 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bullock:1981:SLS

[BR81a] A. M. Bullock and R. J. Roberts. Sunburn lesions in salmonid fry: a clinical and histological report. Journal of Fish Diseases, 4(3): 271–275, May 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Burke:1981:IPB

[BR81b] J. Burke and L. Rodgers. Identification of pathogenic bacteria associated with the occurrence of ‘red spot’ in sea mullet, Mugil cephalus L., in south-eastern Queensland. Journal of Fish Diseases, 4(2):153–159, March 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bullock:1982:NOI

[BR82] A. M. Bullock and D. A. Robertson. A note on the occurrence of Ichtyobodo necator (Henneguy, 1883) in a wild population of juvenile plaice, Pleuronectes platessa L. Journal of Fish Diseases, 5(6):531–533, November 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Boonyaratpalin:1984:CBT

[BR84a] S. Boonyaratpalin and W. A. Rogers. Control of the bass tapeworm, Proteocephalus ambloplitis (Leidy), with mebendazole. Journal of Fish Diseases, 7(6):449–456, November 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Brown:1984:CEP

[BR84b] Carolyn Brown and G. Roland. Characterization of exotoxin produced by a shellfish-pathogenic Vibrio sp. Journal of Fish Diseases, 7(2):117–126, March 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bruno:1986:CSP

[Bru86a] D. W. Bruno. Changes in serum parameters of rainbow trout, Salmo gairdneri Richardson, and Atlantic salmon, Salmo salar L., infected with Renibacterium salmoninarum. Journal of Fish Diseases, 9(3):205–211, May 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bruno:1986:HBK

[Bru86b] D. W. Bruno. Histopathology of bacterial kidney disease in laboratory infected rainbow trout, Salmo gairdneri Richardson, and Atlantic salmon, Salmo salar L., with reference to naturally infected fish. Journal of Fish Diseases, 9(6):523–537, November 1986. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bruno:1989:IOR

[Bru89] D. W. Bruno. An investigation into oxytetracycline residues in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 12(2): 77–86, March 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Burreson:1981:CSM

[BS81a] E. M. Burreson and J. P. Sypek. Cryptobia sp. (Mastigophora: Kinetoplastida) from the gills of marine fishes in the Chesapeake Bay. Journal of Fish Diseases, 4(6):519–522, November 1981. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bylund:1981:LTD

[BS81b] G. Bylund and O. Sumari. Laboratory tests with Droncit against diplostomiasis in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 4(3):259–264, May 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bauer:1984:CTC

[BS84a] O. N. Bauer and V. P. Solomatova. The cestode Triaenophorus crassus (Pallas) (Pseudophyllidea: Triaenophoridae) as a pathogen of cage-reared salmonids. Journal of Fish Diseases, 7(6):501–504, November 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Benfey:1984:BRB

[BS84b] T. J. Benfey and A. M. Sutterlin. Binucleated red blood cells in the peripheral blood of an Atlantic salmon, Salmo salar L., alevin. Journal of Fish Diseases, 7(5):415–420, September 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bruno:1987:SAS

[BS87] D. W. Bruno and D. J. Stamps. Saprolegniasis of Atlantic salmon, Salmo salar L., fry. Journal of Fish Diseases, 10(6):513–517, November 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Baran:1989:CPT

[BS89] I. Baran and E. Soylu. Crayfish plague in Turkey. Journal of Fish Diseases, 12(2):193–197, March 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Barker:1989:BFR

[BSB89] G. A. Barker, S. N. Smith, and N. R. Bromage. The bacterial flora of rainbow trout, Salmo gairdneri Richardson, and brown trout, Salmo trutta L., eggs and its relationship to developmental success. Journal of Fish Diseases, 12(4):281–293, July 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Berry:1983:TII

[BSG83] E. S. Berry, T. B. Shea, and J. Gabliks. Two iridovirus isolates from Carassius auratus (L.). Journal of Fish Diseases, 6(6):501– 510, November 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bartholomew:1989:CHR

[BSRF89] J. L. Bartholomew, C. E. Smith, J. S. Rohovec, and J. L. Fryer. Characterization of a host response to the myxosporean parasite, Ceratomyxa shasta (Noble), by histology, scanning electron microscopy and immunological techniques. Journal of Fish Diseases, 12(5):509–522, September 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bell:1986:RPL

[BT86] G. R. Bell and G. S. Traxler. Resistance of the Pacific lamprey, Lampetra tridentata (Gairdner), to challenge by Renibacterium salmoninarum, the causative agent of kidney disease in salmonids. Journal of Fish Diseases, 9(3):277–279, May 1986. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bucke:1980:NAA

[Buc80] D. Bucke. A note on acid-alcohol-fast bacteria in mackerel Scomber scombrus L. Journal of Fish Diseases, 3(2):173–175, March 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bullock:1985:EUB

[Bul85] A. M. Bullock. The effect of ultraviolet-B radiation upon the skin of the plaice, Pleuronectes platessa L., infested with the bodonid ectoparasite Ichthyobodo necator (Henneguy, 1883). Journal of Fish Diseases, 8(6):547–550, November 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bylund:1980:OEP

[BVN80] G. Bylund, E. T. Valtonen, and E. Niemelä. Observations on epidermal papillomata in wild and cultured Atlantic salmon Salmo salar L. in Finland. Journal of Fish Diseases, 3(6):525–528, November 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Blazer:1983:CDR

[BW83] V. S. Blazer and R. E. Wolke. Ceroid deposition, retinal degeneration and renal calcium oxalate crystals in cultured clownfish, Amphiprion ocellaris. Journal of Fish Diseases, 6(4):365–376, July 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Bucke:1984:HVH

[BWF84] D. Bucke, B. Watermann, and S. Feist. Histological variations of hepato-splenic organs from the North Sea dab, Limanda limanda (L.). Journal of Fish Diseases, 7(4):255–268, July 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Crandall:1981:MIN

[CB81] T. A. Crandall and P. R. Bowser. A microsporidian infection in a natural population of mosquitofish, Gambusia affinis (Baird and Girard). Journal of Fish Diseases, 4(4):317–324, July 1981. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Campbell:1982:IBA

[CB82] A. C. Campbell and J. A. Buswell. An investigation into the bacterial aetiology of ‘black patch necrosis’ in Dover sole, Solea solea L. Journal of Fish Diseases, 5(6):495–508, November 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cornacchia:1984:EDG

[CC84] J. W. Cornacchia and J. E. Colt. The effects of dissolved gas supersaturation on larval striped bass, Morone saxatilis (Walbaum). Journal of Fish Diseases, 7(1):15–27, January 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cusack:1985:RBM

[CC85] R. Cusack and D. K. Cone. A report of bacterial microcolonies on the surface of Gyrodactylus (Monogenea). Journal of Fish Diseases, 8(1):125–127, January 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cusack:1986:RPV

[CC86] R. Cusack and D. K. Cone. A review of parasites as vectors of viral and bacterial diseases of fish. Journal of Fish Diseases, 9(2): 169–171, March 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Coley:1983:DLT

[CCK83] T. C. Coley, A. J. Chacko, and G. W. Klontz. Development of a lavage technique for sampling Ceratomyxa shasta in adult salmonids. Journal of Fish Diseases, 6(3):317–319, May 1983. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cossarini-Dunier:1986:PAE

[CD86] Muriel Cossarini-Dunier. Protection against enteric redmouth disease in rainbow trout, Salmo gairdneri Richardson, after vaccination with Yersinia ruckeri bacterin. Journal of Fish Diseases, 9(1): 27–33, January 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Callinan:1989:PRS

[CFV89] R. B. Callinan, G. C. Fraser, and J. L. Virgona. Pathology of red spot disease in sea mullet, Mugil cephalus L., from eastern Australia. Journal of Fish Diseases, 12(5):467–479, September 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ceschia:1987:VSY

[CGBF87] G. Ceschia, G. Giorgetti, G. Bertoldini, and S. Fontebasso. The in vitro sensitivity of Yersinia ruckeri to specific antibiotics. Journal of Fish Diseases, 10(1):65–67, January 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Carson:1988:VAA

[CH88] J. Carson and J. Handlinger. Virulence of the aetiological agent of goldfish ulcer disease in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 11(6):471–479, November 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Clifton-Hadley:1987:EMG

[CHA87] R. S. Clifton-Hadley and D. J. Alderman. The effects of malachite green upon proliferative kidney disease. Journal of Fish Diseases, 10(2):101–107, March 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Clifton-Hadley:1984:PKD

[CHBR84] R. S. Clifton-Hadley, D. Bucke, and R. H. Richards. Proliferative kidney disease of salmonid fish: a review. Journal of Fish Diseases, 7(5):363–377, September 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Clifton-Hadley:1987:SSC

[CHBR87] R. S. Clifton-Hadley, D. Bucke, and R. H. Richards. A study of the sequential clinical and pathological changes during proliferative kidney disease in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 10(5):335–352, September 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Clifton-Hadley:1987:FCH

[CHRB87] R. S. Clifton-Hadley, R. H. Richards, and D. Bucke. Further consideration of the haematology of proliferative kidney disease (PKD) in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 10(6):435–444, November 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Callinan:1989:BPR

[CK89a] R. B. Callinan and J. A. Keep. Bacteriology and parasitology of red spot disease in sea mullet, Mugil cephalus L., from eastern Australia. Journal of Fish Diseases, 12(4):349–356, July 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chen:1989:ICC

[CK89b] S. N. Chen and G. H. Kou. Infection of cultured cells from the lymphoid organ of Penaeus monodon Fabricius by monodon-type baculovirus (MBV). Journal of Fish Diseases, 12(1):73–76, January 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Csaba:1984:SPP

[CKGB^{+}84] G. Csaba, Éva Kovács-Gayer, L. Békési, Maria Bucsek, J. Szakolczai, and K. Molnár. Studies into the possible protozoan aetiology of swimbladder inflammation in carp fry. Journal of Fish Diseases, 7(1):39–56, January 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Casillas:1985:SCD

[CMRM85] E. Casillas, M. S. Myers, Linda D. Rhodes, and B. B. Mccain. Serum chemistry of diseased English sole, Parophrys vetulus Girard, from polluted areas of Puget Sound, Washington. Journal of Fish Diseases, 8(5):437–449, September 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cornick:1984:AAS

[CMZS84] J. W. Cornick, C. M. Morrison, B. Zwicker, and G. Shum. Atypical Aeromonas salmonicida infection in Atlantic cod, Gadus morhua L. Journal of Fish Diseases, 7(6):495–499, November 1984. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Canning:1980:GPP

[CN80] Elizabeth U. Canning and J. P. Nicholas. Genus Pleistophora (Phylum Microspora): redescription of the type species, Pleistophora typicalis Gurley, 1893 and ultrastructural characterization of the genus. Journal of Fish Diseases, 3(4):317–338, July 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cheung:1980:SMU

[CNR80] P. J. Cheung, R. F. Nigrelli, and G. D. Ruggieri. Studies on the morphology of Uronema marinum Dujardin (Ciliatea: Uronematidae) with a description of the histopathology of the infection in marine fishes. Journal of Fish Diseases, 3(4):295–303, July 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cheung:1981:OOB

[CNR81] P. J. Cheung, R. F. Nigrelli, and G. D. Ruggieri. Oodinium ocellatum (Brown, 1931) (Dinoflagellata) in the kidney and other internal tissues of pork fish, Anisotremus virginicus (L.). Journal of Fish Diseases, 4(6):523–525, November 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cheung:1983:PMS

[CNR83] P. J. Cheung, R. F. Nigrelli, and G. D. Ruggieri. Pentacapsula muscularis sp. nov. (Myxosporea: Pentacapsulidea); a histozoic parasite of butterflyfish, Chaetodon collare Bloch. Journal of Fish Diseases, 6(4):393–395, July 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cheung:1982:TSL

[CNRC82] P. J. Cheung, R. F. Nigrelli, G. D. Ruggieri, and A. Cilia. Treatment of skin lesions in captive lemon sharks, Negaprion brevirostris (Poey), caused by monogeneans (Dermophthirius sp.). Journal of Fish Diseases, 5(2):167–170, March 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cone:1982:LSD

[Con82] D. K. Cone. A Lactobacillus sp. from diseased female rainbow trout, Salmo gairdneri Richardson, in Newfoundland, Canada. Journal of Fish Diseases, 5(6):479–485, November 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Copland:1981:ODM

[Cop81] J. W. Copland. The occurrence and distribution of Myxidium giardi Cépède, 1906 in wild and cultured European eels, Anguilla anguilla L., in England. Journal of Fish Diseases, 4(3):231–242, May 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Copland:1982:MDI

[Cop82] J. W. Copland. Myxobolus dermatobia (Ishii, 1915) infection of wild eels, Anguilla anguilla L., in England. Journal of Fish Diseases, 5(6):549–552, November 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Copland:1983:PMG

[Cop83] J. W. Copland. The pathology of Myxidium giardi Cépède, 1906 infections in wild and cultured eels, Anguilla anguilla L. Journal of Fish Diseases, 6(5):451–460, September 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cormier:1986:FSH

[Cor86] Susan M. Cormier. Fine structure of hepatocytes and hepatocellular carcinoma of the Atlantic tomcod, Microgadus tomcod (Walbaum). Journal of Fish Diseases, 9(3):179–194, May 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Couch:1984:HEP

[Cou84] J. A. Couch. Histopathology and enlargement of the pituitary of a teleost exposed to the herbicide trifluralin. Journal of Fish Diseases, 7(2):157–163, March 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Clayton:1988:IMS

[CP88] G. M. Clayton and D. J. Price. Ichthyophthirius multifiliis: standardization of the infection-response model in Ameca splendens (Miller & Fitzsimons). Journal of Fish Diseases, 11(5):371–377, September 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Chumnongsitathum:1988:HEM

[CPH88] B. Chumnongsitathum, J. A. Plumb, and V. Hilge. Histopathology, electron microscopy and isolation of channel catfish virus in experimentally infected European catfish, Silurus glanis L. Journal of Fish Diseases, 11(4):351–357, July 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cusack:1988:SBM

[CRC88] R. Cusack, T. Rand, and D. Cone. A study of bacterial microcolonies associated with the body surface of Gyrodactylus colemanensis Mizelle & Kritsky, 1967 (Monogenea), parasitizing Salmo gairdneri Richardson. Journal of Fish Diseases, 11(3):271–274, May 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cormier:1989:HCF

[CRS^{+}89] S. M. Cormier, R. N. Racine, C. E. Smith, W. P. Dey, and T. H. Peck. Hepatocellular carcinoma and fatty infiltration in the Atlantic tomcod, Microgadus tomcod (Walbaum). Journal of Fish Diseases, 12(2):105–116, March 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Crawshaw:1986:MKK

[CS86] M. T. Crawshaw and R. A. Sweeting. Myxobolus koi Kudo, 1919: a new record for Britain. Journal of Fish Diseases, 9(5):465–467, September 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cann:1982:OBR

[CT82] D. C. Cann and L. Y. Taylor. An outbreak of botulism in rainbow trout, Salmo gairdneri Richardson, farmed in Britain. Journal of Fish Diseases, 5(5):393–399, September 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cann:1984:ERC

[CT84] D. C. Cann and Lesley Y. Taylor. An evaluation of residual contamination by Clostridium botulinum in a trout farm following an outbreak of botulism in the fish stock. Journal of Fish Diseases, 7 (5):391–396, September 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Conroy:1980:UNC

[CVM80] D. A. Conroy, C. Vásquez, and J. Morales. The use of nifurpirinol in the control of chilodonelliasis and ichthyophthiriasis in species of Colombian cold water fish. Journal of Fish Diseases, 3(6):531– 532, November 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Copland:1982:PSI

[CW82] J. W. Copland and L. G. Willoughby. The pathology of Saprolegnia infections of Anguilla anguilla L. elvers. Journal of Fish Diseases, 5(5):421–428, September 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Cone:1984:INH

[CW84] D. K. Cone and M. Wiles. Ichthyobodo necator (Henneguy, 1883) from winter flounder, Pseudopleuronectes americanus (Walbaum), in the north-west Atlantic Ocean. Journal of Fish Diseases, 7(1): 87–89, January 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Duran:1987:CSE

[DARD87] A. Durán, L. B. Rodríguez Aparicio, A. Reglero, and J. Pérez Díaz. Changes in serum enzymes of Saprolegnia-infected brown trout, Salmo trutta L. Journal of Fish Diseases, 10(6):505–507, November 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Davies:1985:ZRT

[Dav85] Angela J. Davies. Zschokkella russelli Tripathi (Myxozoa: Myxosporea) from the five-bearded rockling, Ciliata mustela L., (Teleostei: Gadidae) in Wales. Journal of Fish Diseases, 8(3): 299–308, May 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Delves-Broughton:1980:FOC

[DBFW80] J. Delves-Broughton, J. K. Fawell, and D. Woods. The first occurrence of ‘cauliflower disease’ of eels Anguilla anguilla L. in the British Isles. Journal of Fish Diseases, 3(3):255–256, May 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Devesa:1989:USF

[DBT89] S. Devesa, J. L. Barja, and A. E. Toranzo. Ulcerative skin and fin lesions in reared turbot, Scophthalmus maximus (L.). Journal of Fish Diseases, 12(4):323–333, July 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Drinan:1989:ORC

[DDMS89] E. M. Drinan, J. Donlon, S. Molloy, and P. R. Smith. Observations on the role of the caseinase enzyme in the pathogenesis of furunculosis. Journal of Fish Diseases, 12(1):45–50, January 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

DeKinkelin:1980:OMI

[De 80] P. De Kinkelin. Occurrence of a microsporidian infection in zebra danio Brachydanio rerio (Hamilton-Buchanan). Journal of Fish Diseases, 3(1):71–73, January 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Daoust:1984:PCA

[DF84] P.-Y. Daoust and H. W. Ferguson. The pathology of chronic ammonia toxicity in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 7(3):199–205, May 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Daoust:1985:NGD

[DF85] P.-Y. Daoust and H. W. Ferguson. Nodular gill disease: a unique form of proliferative gill disease in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 8(6):511–522, November 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Daoust:1986:PRN

[DF86] P.-Y. Daoust and H. W. Ferguson. Potential for recovery in nodular gill disease of rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 9(4):313–318, July 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Davies:1989:MBD

[DF89] R. L. Davies and G. N. Frerichs. Morphological and biochemical differences among isolates of Yersinia ruckeri obtained from wide geographical areas. Journal of Fish Diseases, 12(4):357–365, July 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

DeKinkelin:1982:ESS

[DG82] P. De Kinkelin and J. Gastric. An experimental study of the susceptibility of Atlantic salmon fry, Salmo salar L., to viral haemorrhagic septicaemia. Journal of Fish Diseases, 5(1):57–65, January 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dixon:1983:IIP

[DH83] P. F. Dixon and B. J. Hill. Inactivation of infectious pancreatic necrosis virus for vaccine use. Journal of Fish Diseases, 6(5):399– 409, September 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Davis:1984:ERA

[DH84] J. F. Davis and S. S. Hayasaka. The enhancement of resistance of the American eel, Anguilla rostrata Le Sueur, to a pathogenic bacterium, Aeromonas hydrophila, by an extract of the tunicate, Ecteinascidia turbinata. Journal of Fish Diseases, 7(4):311–316, July 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Diamant:1987:UPI

[Dia87] A. Diamant. Ultrastructure and pathogenesis of Ichthyobodo sp. from wild common dab, Limanda limanda L., in the North Sea. Journal of Fish Diseases, 10(3):241–247, May 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dixon:1987:IIP

[Dix87] P. P. Dixon. Inhibition of infectious pancreatic necrosis virus infectivity by extracts of rainbow trout, Salmo gairdneri Richardson, tissue. Journal of Fish Diseases, 10(5):371–378, September 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Desser:1982:LEM

[DK82] S. S. Desser and R. A. Khan. Light and electron microscope observations on pathological changes in the gills of the marine fish Lycodes lavalaei Vladykov and Tremblay associated with the proliferation of an unidentified cell. Journal of Fish Diseases, 5(5):351– 364, September 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ding-Ke:1983:NSP

[DKHJ83] Hua Ding-Ke and Dong Han-Ji. A new species of Pleistophora from big-eyes, Priacanthus tayenus Richardson, and P. macracanthus Cuvier and Valenciennes, from the South China Sea. Journal of Fish Diseases, 6(3):293–301, May 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dykova:1980:TRM

[DL80] Iva Dyková and Jiří Lom. Tissue reactions to microsporidian infections in fish. Journal of Fish Diseases, 3(4):265–283, July 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dykova:1981:FCC

[DL81] Iva Dyková and J. Lom. Fish coccidia: critical notes on life cycles, classification and pathogenicity. Journal of Fish Diseases, 4 (6):487–505, November 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Down:1989:HGN

[DL89] N. E. Down and J. F. Leatherland. Histopathology of gonadal neoplasms in cyprinid fish from the lower Great Lakes of North America. Journal of Fish Diseases, 12(5):415–437, September 1989. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dykova:1988:NHI

[DLF88] I. Dyková, J. Lom, and E. Fajer. A new haplosporean infecting the hepatopancreas in the penaeid shrimp, Penaeus vannamei. Journal of Fish Diseases, 11(1):15–22, January 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dykova:1983:EBS

[DLG83] Iva Dyková, Jiří Lom, and Ginka Grupcheva. Eimeria branchiphila sp.nov. sporulating in the gill filaments of roach, Rutilus rutilus L. Journal of Fish Diseases, 6(1):13–18, January 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dickerson:1985:EII

[DLG85] H. W. Dickerson, A. L. Lohr, and J. B. Gratzek. Experimental intraperitoneal infection of channel catfish, Ictalurus punctatus (Rafinesque), with Ichthyophthirius multifiliis (Fouquet). Journal of Fish Diseases, 8(1):139–142, January 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dykstra:1989:UMS

[DLN^{+}89] M. J. Dykstra, J. F. Levine, E. J. Noga, J. H. Hawkins, P. Gerdes, W. J. Hargis, Jr., H. J. Grier, and D. Te Strake. Ulcerative mycosis: a serious menhaden disease of the southeastern coastal fisheries of the United States. Journal of Fish Diseases, 12(2):175–178, March 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

DSilva:1984:ETP

[DMdK84] J. D’Silva, M. F. Mulcahy, and P. de Kinkelin. Experimental transmission of proliferative kidney disease in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 7(3):235–239, May 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Donoghue:1986:NPB

[Don86] S. Donoghue. A note on the presence of bacteria on the surface of the parasitic copepod Thersitina gasterostei Pagenstecher. Journal of Fish Diseases, 9(4):349–351, July 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dalsgaard:1986:AAS

[DP86] I. Dalsgaard and H. Paulsen. Atypical Aeromonas salmonicida isolated from diseased sand-eels, Ammodytes lancea (Cuvier) and Hyperoplus lanceolatus (Lesauvege). Journal of Fish Diseases, 9 (4):361–364, July 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Desser:1988:UOC

[DPS88] S. Desser, W. Paterson, and D. Steinhagen. Ultrastructural observations on the causative agent of epitheliocystis in the brown bullhead, Ictalurus nebulosus Lesueur, from Ontario and a comparison with the chlamydiae of higher vertebrates. Journal of Fish Diseases, 11(6):453–460, November 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dick:1982:EIW

[DR82] T. A. Dick and R. Rosen. Experimental infections of whitefish, Coregonus clupeaformis (Mitchill), with Triaenophorus crassus Forel. Journal of Fish Diseases, 5(1):83–86, January 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Davies:1988:FSZ

[DS88] A. J. Davies and I. K. Sienkowski. Further studies on Zschokkella russelli Tripathi (Myxozoa: Myxosporea) from Ciliata mustela L. (Teleostei; Gadidae), with emphasis on ultrastructural pathology and sporogenesis. Journal of Fish Diseases, 11(4):325–336, July 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dwilow:1987:IYR

[DSK87] A. G. Dwilow, B. W. Souter, and K. Knight. Isolation of Yersinia ruckeri from burbot. Lota lota (L.), from the Mackenzie River, Canada. Journal of Fish Diseases, 10(4):315–317, July 1987. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dorson:1981:IFA

[DT81] M. Dorson and C. Touchy. The influence of fish age and water temperature on mortalities of rainbow trout, Salmo gairdneri Richardson, caused by a European strain of infectious pancreatic necrosis virus. Journal of Fish Diseases, 4(3):213–221, May 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Dorson:1984:RPP

[DTC^{+}84] M. Dorson, C. Torchy, S. Chilmonczyk, P. de Kinkelin, and C. Michel. A rhabdovirus pathogenic for perch, Perca fluviatilis L.: isolation and preliminary study. Journal of Fish Diseases, 7 (3):241–245, May 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Egidius:1981:CWV

[EACR81] E. Egidius, K. Andersen, E. Clausen, and J. Raa. Cold-water vibriosis or ‘Hitra disease’ in Norwegian salmonid farming. Journal of Fish Diseases, 4(4):353–354, July 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Elston:1987:PSF

[EBF^{+}87] R. A. Elston, J. H. Beattie, C. Friedman, R. Hedrick, and M. L. Kent. Pathology and significance of fatal inflammatory bacteraemia in the Pacific oyster, Crassostrea gigas Thünberg. Journal of Fish Diseases, 10(2):121–132, March 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ellis:1988:LRB

[EBS88] A. E. Ellis, A. S. Burrows, and K. J. Stapleton. Lack of relationship between virulence of Aeromonas salmonicida and the putative virulence factors: A-layer, extracellular proteases and extracellular haemolysins. Journal of Fish Diseases, 11(4):309–323, July 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Eckmann:1987:PCM

[Eck87] R. Eckmann. Pathological changes in the midgut epithelium of grayling, Thymallus thymallus L., larvae reared on different kinds of food, and their relation to mortality and growth. Journal of Fish Diseases, 10(2):91–99, March 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Engelman:1984:UES

[ECM84] R. W. Engelman, L. L. Collier, and J. B. Marliave. Unilateral exophthalmos in Sebastes spp.: histopathologic lesions. Journal of Fish Diseases, 7(6):467–476, November 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Evenberg:1988:VIP

[EdGLV88] D. Evenberg, P. de Graaff, B. Lugtenberg, and W. B. Van Muiswinkel. Vaccine-induced protective immunity against Aeromonas salmonicida tested in experimental carp erythrodermatitis. Journal of Fish Diseases, 11(4):337–350, July 1988. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ellis:1983:HSB

[EDS83] A. E. Ellis, G. Dear, and D. J. Stewart. Histopathology of ‘Sekitenbyo’ caused by Pseudomonas anguittiseptica in the European eel, Anguilla anguilla L., in Scotland. Journal of Fish Diseases, 6(1): 77–79, January 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Elston:1982:CIS

[EEC82] R. Elston, Elisa L. Elliot, and R. R. Colwell. Conchiolin infection and surface coating Vibrio: shell fragility, growth depression and mortalities in cultured oysters and clams, Crassostrea virginica, Ostrea edulis and Mercenaria mercenaria. Journal of Fish Diseases, 5(4):265–284, July 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Eurell:1981:EEW

[EH81] Jo Ann Eurell and W. E. Haensly. The effects of exposure to watersoluble fractions of crude oil on selected histochemical parameters of the liver of Atlantic croaker, Micropogon undulatus L. Journal of Fish Diseases, 4(3):187–194, May 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Elger:1983:GDC

[EH83] M. Elger and H. Hentschel. Glomerular disease in cultured rainbow trout, Salmo gairdneri Richardson, suffering from presumptive chronic viral haemorrhagic septicaemia. Journal of Fish Diseases, 6(3):211–229, May 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Espelid:1988:MAA

[EHHJ88] S. Espelid, K. O. Holm, K. Hjelmeland, and T. Jørgensen. Monoclonal antibodies against Vibrio salmonicida: the causative agent of coldwater vibriosis (‘Hitra disease’) in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 11(3):207–214, May 1988. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ellis:1981:RAS

[EHM81] A. E. Ellis, T. S. Hastings, and A. L. S. Munro. The role of Aeromonas salmonicida extracellular products in the pathology of furunculosis. Journal of Fish Diseases, 4(1):41–51, January 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Egidius:1981:PTA

[EJL81] E. C. Egidius, J. V. Johannessen, and E. Lange. Pseudobranchial tumours in Atlantic cod, Gadus morhua L., from the Barents Sea. Journal of Fish Diseases, 4(6):527–532, November 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Evelyn:1984:FEP

[EKPP84] T. P. T. Evelyn, J. E. Ketcheson, and L. Prosperi-Porta. Further evidence for the presence of Renibacterium salmoninarum in salmonid eggs and for the failure of povidone-iodine to reduce the intra-ovum infection rate in water-hardened eggs. Journal of Fish Diseases, 7(3):173–182, May 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Elarifi:1982:HLA

[Ela82] A. E. Elarifi. The histopathology of larval anisakid nematode infections in the liver of whiting, Merlangius merlangus (L.), with some observations on blood leucocytes of the fish. Journal of Fish Diseases, 5(5):411–419, September 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ellis:1980:ATS

[Ell80] A. E. Ellis. Antigen-trapping in the spleen and kidney of the plaice Pleuronectes platessa L. Journal of Fish Diseases, 3(5):413–426, September 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ellis:1984:BFE

[Ell84] A. E. Ellis. Bizzare forms of erythrocytes in a specimen of plaice, Pleuronectes platessa L. Journal of Fish Diseases, 7(5):411–414, September 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Elston:1980:UAS

[Els80] R. Elston. Ultrastructural aspects of a serious disease of hatchery reared larval oysters, Crassostrea gigas Thünberg. Journal of Fish Diseases, 3(1):1–10, January 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Elston:1983:HOI

[Els83] R. Elston. Histopathology of oxygen intoxication in the juvenile red abalone, Haliotis rufescens Swainson. Journal of Fish Diseases, 6(2):101–110, March 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Elston:1986:OBR

[Els86] R. Elston. Occurrence of branchial rickettsiales-like infections in two bivalve molluscs, Tapes japonica and Patinopecten yessoensis, with comments on their significance. Journal of Fish Diseases, 9(1):69–71, January 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ellis:1985:PKD

[EMM85] A. E. Ellis, A. H. Mcvicar, and A. L. S. Munro. Proliferative kidney disease in brown trout, Salmo trutta L., and Atlantic salmon, Salmo salar L., parr: histopathological and epidemiological observations. Journal of Fish Diseases, 8(2):197–208, March 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Elger:1986:KRT

[ENH86] B. Elger, M. Neukirch, and H. Hentschel. The kidney of rainbow trout, Salmo gairdneri Richardson, in the acute phase of viral haemorrhagic septicaemia: in vivo experiments on the renal excretion of fluid, electrolytes and protein. Journal of Fish Diseases, 9 (5):381–392, September 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Emerson:1985:EPVa

[EPB85a] C. J. Emerson, J. F. Payne, and A. K. Bal. Evidence for the presence of a viral non-lymphocystis type disease in winter flounder, Pseudopleuronectes americanus (Walbaum), from the north-west Atlantic. Journal of Fish Diseases, 8(1):75–89, January 1985. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Emerson:1985:EPVb

[EPB85b] C. J. Emerson, J. F. Payne, and A. K. Bal. Evidence for the presence of a viral non-lymphocystis type disease in winter flounder, Pseudopleuronectes americanus (Walbaum), from the north-west Atlantic. Journal of Fish Diseases, 8(1):91–102, January 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Evelyn:1986:PKD

[EPPK86] T. P. T. Evelyn, L. Prosperi-Porta, and J. E. Ketcheson. Persistence of the kidney-disease bacterium, Renibacterium salmoninarum, in Coho salmon, Oncorhynchus kisutch (Walbaum), eggs treated during and after water-hardening with povidone-iodine. Journal of Fish Diseases, 9(5):461–464, September 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Elliott:1980:AUDb

[ES80a] Diane G. Elliott and E. B. Shotts, Jr. Aetiology of an ulcerative disease in goldfish, Carassius auratus (L.): experimental induction of the disease. Journal of Fish Diseases, 3(2):145–151, March 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Elliott:1980:AUDa

[ES80b] Diane G. Elliott and E. B. Shotts, Jr. Aetiology of an ulcerative disease in goldfish Carassius auratus (L.): microbiological examination of diseased fish from seven locations. Journal of Fish Diseases, 3(2):133–143, March 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Elston:1983:PVC

[EW83] R. Elston and G. S. Lock Wood. Pathogenesis of vibriosis in cultured juvenile red abalone, Haliotis rufescens Swainson. Journal of Fish Diseases, 6(2):111–128, March 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ellis:1983:SFD

[EWM83] A. E. Ellis, I. F. Waddell, and D. W. Minter. A systemic fungal disease in Atlantic salmon parr, Salmo salar L., caused by a species of Phialophora. Journal of Fish Diseases, 6(6):511–523, November 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Faisal:1984:SVC

[FA84] M. Faisal and W. Ahne. Spring viraemia of carp virus (SVCV): comparison of immunoperoxidase, fluorescent antibody and cell culture isolation techniques for detection of antigen. Journal of Fish Diseases, 7(1):57–64, January 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Feist:1987:UAP

[FB87] S. W. Feist and D. Bucke. Ultrastructural aspects of PKX, the causative agent of proliferative kidney disease in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 10(4):323– 327, July 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fuhrmann:1983:OER

[FBS83] H. Fuhrmann, K. H. Böhm, and H.-J. Schlotfeldt. An outbreak of enteric redmouth disease in West Germany. Journal of Fish Diseases, 6(3):309–311, May 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fournie:1988:EPA

[FBV88] J. W. Fournie, J. J. Black, and A. D. Vethaak. Exocrine pancreatic adenomas in the greater redhorse, Moxostoma valenciennesi Jordan, and in the European flounder, Platichthys flesus (L.). Journal of Fish Diseases, 11(5):445–448, September 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fyfe:1988:CEI

[FCM88] Lorn A. Fyfe, G. Coleman, and A. L. S. Munro. The combined effect of isolated Aeromonas salmonicida protease and haemolysin on Atlantic salmon, Salmo salar L., compared with that of a total extracellular products preparation. Journal of Fish Diseases, 11 (1):101–104, January 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ferguson:1981:EWT

[Fer81] H. W. Ferguson. The effects of water temperature on the development of proliferative kidney disease in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 4(2):175–177, March 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fyfe:1986:SPE

[FFCM86] Lorn A. Fyfe, A. Finley, G. Coleman, and A. L. S. Munro. A study of the pathological effect of isolated Aeromonas salmonidda extracellular protease on Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 9(5):403–409, September 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fendrick:1982:CSF

[FGL82] J. L. Fendrick, W. J. Groberg, Jr., and J. C. Leong. Comparative sensitivity of five fish cell lines to wild type infectious haematopoietic necrosis virus from two Oregon sources. Journal of Fish Diseases, 5(2):87–95, March 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Farley:1980:ACI

[FH80] D. G. Farley and R. Heckmann. Attempts to control Ichthyophthirius multifiliis Fouquet (Ciliophora: Ophryoglenidae) by chemotherapy and electrotherapy. Journal of Fish Diseases, 3 (3):203–212, May 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

FjOLstad:1985:MMD

[FH85] M. FjØLstad and A. L. Heyeraas. Muscular and myocardial degeneration in cultured Atlantic salmon, Salmo salar L., suffering from ‘Hitra disease’. Journal of Fish Diseases, 8(4):367–372, July 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Foissner:1985:HCF

[FHM85] W. Foissner, G. L. Hoffman, and A. J. Mitchell. Heteropolaria colisarum Foissner & Schubert, 1977 (Protozoa: Epistylididae) of North American freshwater fishes. Journal of Fish Diseases, 8(2): 145–160, March 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Frerichs:1989:UDR

[FHW89] G. N. Frerichs, B. J. Hill, and K. Way. Ulcerative disease rhabdovirus: cell-line susceptibility and serological comparison with other fish rhabdoviruses. Journal of Fish Diseases, 12(1):51–56, January 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fijan:1981:POB

[Fij81] N. Fijan. Professor Oleg N. Bauer. Journal of Fish Diseases, 4 (6):443–444, November 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fitzsimons:1988:HAS

[Fit88] J. D. Fitzsimons. Hepatocellular adenoma in the slimy sculpin, Coitus cognatus Richardson. Journal of Fish Diseases, 11(1):89– 91, January 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fournie:1985:RSC

[FO85] J. W. Fournie and R. M. Overstreet. Retinoblastoma in the spring cavefish, Chologaster agassizi Putnam. Journal of Fish Diseases, 8 (4):377–381, July 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fournie:1985:MCH

[FOB85] J. W. Fournie, R. M. Overstreet, and L. H. Bullock. Multiple capillary haemangiomas in the scamp, Mycteroperca phenax Jordan and Swain. Journal of Fish Diseases, 8(6):551–555, November 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ferguson:1989:VNG

[FODS89] H. W. Ferguson, V. E. Ostland, J. G. Daly, and R. M. Stevenson. Visceral necrosis of goldfish, Carassius auratus (L.): a new bacterial disease. Journal of Fish Diseases, 12(3):249–256, May 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ferguson:1980:PSM

[FR80] H. W. Ferguson and D. A. Rice. Post-spawning mortalities in brown trout Salmo trutta L. Journal of Fish Diseases, 3(2):153– 160, March 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ferguson:1986:SDC

[FRR^{+}86] H. W. Ferguson, R. J. Roberts, R. H. Richards, R. O. Collins, and D. A. Rice. Severe degenerative cardiomyopathy associated with pancreas disease in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 9(1):95–98, January 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Frerichs:1985:AIR

[FSC85] G. N. Frerichs, J. A. Stewart, and R. O. Collins. Atypical infection of rainbow trout, Salmo gairdneri Richardson, with Yersinia ruckeri. Journal of Fish Diseases, 8(4):383–387, July 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Follett:1987:IHN

[FTH87] J. E. Follett, J. B. Thomas, and A. K. Hauck. Infectious haematopoietic necrosis virus in moribund and dead juvenile churn, Oncorhynchus keta (Walbaum), and Chinook, O. tshawytscha (Walbaum), salmon and spawning adult chum salmon at an Alaskan hatchery. Journal of Fish Diseases, 10(4):309–313, July 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Fournie:1987:SCC

[FVO87] J. W. Fournie, W. K. Vogelbein, and R. M. Overstreet. Squamous cell carcinoma in the gulf menhaden, Brevoortia patronus Goode. Journal of Fish Diseases, 10(2):133–136, March 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gaten:1987:AEF

[Gat87] E. Gaten. Aggregation of the eye fluke Diplostomum spathaceum (Digenea: Diplostomatidae) in the lenses of various species of fish. Journal of Fish Diseases, 10(1):69–74, January 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Grimes:1989:PIF

[GBC^{+}89] D. J. Grimes, J. Burgess, J. A. Crunkleton, Jr., P. R. Brayton, and R. R. Colwell. Potential invasive factors associated with Vibrio carchariae, an opportunistic pathogen for sharks. Journal of Fish Diseases, 12(1):69–72, January 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Giorgetti:1982:VRT

[GC82] G. Giorgetti and G. Ceschia. Vibriosis in rainbow trout, Salmo gairdneri Richardson, in fresh water in north-eastern Italy. Journal of Fish Diseases, 5(2):125–130, March 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gulka:1985:HRR

[GC85a] G. Gulka and Pei Wen Chang. Host response to rickettsial infection in blue mussel, Mytilus edulis L. Journal of Fish Diseases, 8(3): 319–323, May 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gulka:1985:PIR

[GC85b] G. Gulka and Pei Wen Chang. Pathogenicity and infectivity of a rickettsia-like organism in the sea scallop, Placopecten magellanicus. Journal of Fish Diseases, 8(3):309–318, May 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gulka:1983:PIA

[GCM83] G. Gulka, Pei Wen Chang, and Karen A. Marti. Prokaryotic infection associated with a mass mortality of the sea scallop, Placopecten magellanicus. Journal of Fish Diseases, 6(4):355–364, July 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gastric:1980:OVH

[GD80] Jeannette Gastric and P. De Kinkelin. Occurrence of viral haemorrhagic septicaemia in rainbow trout Salmo gairdneri Richardson reared in sea-water. Journal of Fish Diseases, 3(1):21–27, January 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Goldes:1986:PIS

[GFDM86] S. A. Goldes, H. W. Ferguson, P.-Y. Daoust, and R. D. Moccia. Phagocytosis of the inert suspended clay kaolin by the gills of rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 9(2):147–151, March 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Goldes:1988:HEI

[GFMD88] S. A. Goldes, H. W. Ferguson, R. D. Moccia, and P.-Y. Daoust. Histological effects of the inert suspended clay kaolin on the gills of juvenile rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 11(1):23–33, January 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gratzek:1983:ULC

[GGL^{+}83] J. B. Gratzek, Jeannine P. Gilbert, Amy L. Lohr, E. B. Shotts, Jr., and J. Brown. Ultraviolet light control of Ichthyophthirius multifiliis Fouquet in a closed fish culture recirculation system. Journal of Fish Diseases, 6(2):145–153, March 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Grimes:1985:EIL

[GGM85] D. J. Grimes, S. H. Gruber, and E. B. May. Experimental infection of lemon sharks, Negaprion brevirostris (Poey), with Vibrio species. Journal of Fish Diseases, 8(2):173–180, March 1985. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Gjerde:1984:OCA

[Gje84] J. Gjerde. Occurrence and characterization of Aerococcus viridans from lobsters, Homarus gammarus L., dying in captivity. Journal of Fish Diseases, 7(5):355–362, September 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Grondel:1987:PTD

[GND^{+}87] J. L. Grondel, J. F. M. Nouws, M. De Jong, A. R. Schutte, and F. Driessens. Pharmacokinetics and tissue distribution of oxytetracycline in carp, Cyprinus carpio L., following different routes of administration. Journal of Fish Diseases, 10(3):153–163, May 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Grondel:1987:IAI

[GNV87] J. L. Grondel, J. F. M. Nouws, and W. B. Van Muiswinkel. The influence of antibiotics on the immune system: immunopharmokinetic investigations on the primary anti-SRBC response in carp, Cyprinus carpio L., after oxytetracycline injection. Journal of Fish Diseases, 10(1):35–43, January 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Grinde:1989:LRT

[Gri89] B. Grinde. Lysozyme from rainbow trout, Salmo gairdneri Richardson, as an antibacterial agent against fish pathogens. Journal of Fish Diseases, 12(2):95–104, March 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Grizzle:1981:PBB

[GSS81] J. M. Grizzle, T. E. Schwedler, and A. L. Scott. Papillomas of black bullheads, Ictalurus melas (Rafinesque), living in a chlorinated sewage pond. Journal of Fish Diseases, 4(4):345–351, July 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Grizzle:1983:DFR

[GW83] J. M. Grizzle and E. H. Williams, Jr. Dermal fibroma in a redband parrotfish, Sparisoma aurofrenatum (Valenciennes). Journal of Fish Diseases, 6(2):205–209, March 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Grizzle:1985:LRB

[GW85] J. M. Grizzle and E. H. Williams, Jr. Lesions of a rock beauty, Holacanthus tricolor (Bloch), from Puerto Rico. Journal of Fish Diseases, 8(1):129–134, January 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hoole:1982:USC

[HA82] D. Hoole and C. Arme. Ultrastructural studies on the cellular response of roach, Rutilus rutilus L., to the plerocercoid larva of the pseudophyllidean cestode, Ligula intestinalis. Journal of Fish Diseases, 5(2):131–144, March 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hjeltnes:1987:ESP

[HAEE87] B. Hjeltnes, K. Andersen, H.-M. Ellingsen, and E. Egidius. Experimental studies on the pathogenicity of a Vibrio sp. isolated from Atlantic salmon, Salmo salar L., suffering from Hitra Disease. Journal of Fish Diseases, 10(1):21–27, January 1987. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hauck:1984:MAT

[Hau84] A. K. Hauck. A mortality and associated tissue reactions of Chinook salmon, Oncorhynchus tshawytscha (Walbaum), caused by the microsporidan Loma sp. Journal of Fish Diseases, 7(3):217– 229, May 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hayunga:1980:TAL

[Hay80] E. G. Hayunga. Two atypical lesions from white suckers Catostomus commersoni Lacépède infected by caryophyllid tapeworms. Journal of Fish Diseases, 3(2):167–172, March 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Horne:1983:AVA

[HB83a] M. T. Horne and A. Baxendale. The adhesion of Vibrio anguillarum to host tissues and its role in pathogenesis. Journal of Fish Diseases, 6(5):461–471, September 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Huizinga:1983:NSO

[HB83b] H. W. Huizinga and Joan Budd. Nephroblastoma in the smelt, Osmerus mordax (Mitchill). Journal of Fish Diseases, 6(4):389– 391, July 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hudson:1981:IIP

[HBF81] E. B. Hudson, D. Bucke, and A. Forrest. Isolation of infectious pancreatic necrosis virus from eels, Anguilla anguilla L., in the United Kingdom. Journal of Fish Diseases, 4(5):429–431, September 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hamilton:1987:SPR

[HC87] A. J. Hamilton and E. U. Canning. Studies on the proposed role of Tubifex tubifex (Muller) as an intermediate host in the life cycle of Myxosoma cerebralis (Hofer, 1903). Journal of Fish Diseases, 10 (3):145–151, May 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hamilton:1988:PMA

[HC88] A. J. Hamilton and E. U. Canning. The production of mouse anti-Myxosoma cerebralis antiserum from Percoll(R)-purified spores and its use in immunofluorescent labelling of Historesin(R)- embedded cartilage derived from infected rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 11(2):185–190, March 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hsu:1989:SDI

[HCLW89] Ya-Li Hsu, Shu-Yuan Chiang, Shou-Tzu Lin, and Jen-Leih Wu. The specific detection of infectious pancreatic necrosis virus in infected cells and fish by the immuno dot blot method. Journal of Fish Diseases, 12(6):561–571, November 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hilton:1982:EIL

[HD82] J. W. Hilton and D. G. Dixon. Effect of increased liver glycogen and liver weight on liver function in rainbow trout, Salmo gairdneri Richardson: recovery from anaesthesia and plasma ^{35}Ssulphobromophthalein clearance. Journal of Fish Diseases, 5(3): 185–195, May 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hastings:1988:HIR

[HE88] T. S. Hastings and A. E. Ellis. The humoral immune response of rainbow trout, Salmo gairdneri Richardson, and rabbits to Aeromonas salmonicida extracellular products. Journal of Fish Diseases, 11(2):147–160, March 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Herman:1985:MRT

[Her85] R. L. Herman. Mesothelioma in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 8(4):373–376, July 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Herbert:1988:ICQ

[Her88] B. W. Herbert. Infection of Cherax quadricarinatus (Decapoda: Parastacidae) by the microsporidium Thelohania sp. (Microsporida: Nosematidae). Journal of Fish Diseases, 11(4):301– 308, July 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hentschel:1980:PFA

[HF80] H. Hentschel and H. Finkenstadt. Perfusion fixation and alcianblue PAS staining in histology and histopathology of teleost kidney. Journal of Fish Diseases, 3(6):473–479, November 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hatai:1986:VMA

[HFEJ86] K. Hatai, Y. Fujimaki, S. Egusa, and Y. Jo. A visceral mycosis in ayu fry, Plecoglossus altivelis Temminck & Schlegel, caused by a species of Phoma. Journal of Fish Diseases, 9(2):111–116, March 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hoffman:1985:SFS

[HFH85] G. L. Hoffman, B. Fried, and J. E. Harvey. Sanguinicola fontinalis sp nov. (Digenea: Sanguinicolidae): a blood parasite of brook trout, Salvelinus fontinalis (Mitchill), and longnose dace, Rhinichthys cataractae (Valenciennes). Journal of Fish Diseases, 8(6):529–538, November 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hawkins:1988:RJM

[HFOW88] W. E. Hawkins, J. W. Fournie, R. M. Overstreet, and W. W. Walker. Rhabdomyosarcoma in the Japanese medaka, Oryzia latipes (Temminck & Schlegel) and guppy, Poecilia reticulata Peters. Journal of Fish Diseases, 11(3):259–266, May 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hoffmann:1988:LWG

[HFSPP88] R. W. Hoffmann, Th. Fischer-Scherl, and C. Pfeil-Putzien. Lymphosarcoma in a wild grayling, Thymallus thymallus L.: a case report. Journal of Fish Diseases, 11(3):267–270, May 1988. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hahnel:1982:ETB

[HG82] G. B. Hahnel and R. W. Gould. Effects of temperature on biochemical reactions and drug resistance of virulent and a virulent Aeromonas salmonicida. Journal of Fish Diseases, 5(4):329–337, July 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hicks:1984:HAD

[HG84] B. D. Hicks and J. R. Geraci. A histological assessment of damage in rainbow trout, Salmo gairdneri Richardson, fed rations containing erythromycin. Journal of Fish Diseases, 7(6):457–465, November 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hahnel:1983:SCS

[HGB83] G. B. Hahnel, R. W. Gould, and E. S. Boatman. Serological comparison of selected isolates of Aeromonas salmonicida ssp. salmonicida. Journal of Fish Diseases, 6(1):1–11, January 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hoover:1981:ECN

[HHC^{+}81] D. M. Hoover, F. J. Hoerr, W. W. Carlton, E. J. Hinsman, and H. W. Ferguson. Enteric cryptosporidiosis in a naso tang, Naso lituratus Bloch and Schneider. Journal of Fish Diseases, 4(5):425– 428, September 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hicks:1984:IDS

[HHF84] B. D. Hicks, J. W. Hilton, and H. W. Ferguson. Influence of dietary selenium on the occurrence of nephrocalcinosis in the rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 7(5):379–389, September 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Holm:1987:SVA

[HJ87] K. O. Holm and T. Jørgensen. A successful vaccination of Atlantic salmon, Salmo salar L., against ‘Hitra disease’ or coldwater vibriosis. Journal of Fish Diseases, 10(2):85–90, March 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hermanns:1985:STP

[HK85] W. Hermanns and W. Körting. Sphaerospora tincae Plehn, 1925 in tench, Tinca tinca L., fry. Journal of Fish Diseases, 8(3):281–288, May 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hunter:1980:SIT

[HKF80] Valerie A. Hunter, M. D. Knittel, and J. L. Fryer. Stress-induced transmission of Yersinia ruckeri infection from carriers to recipient steelhead trout Salmo gairdneri Richardson. Journal of Fish Diseases, 3(6):467–472, November 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hoffmann:1986:EES

[HKM86] R. Hoffmann, C. R. Kennedy, and J. Meder. Effects of Eubothrium salvelini Schrank, 1790 on Arctic charr, Salvelinus alpinus (L.), in an alpine lake. Journal of Fish Diseases, 9(2):153–157, March 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hoffmann:1984:HSP

[HL84] R. Hoffmann and R. Lommel. Haematological studies in proliferative kidney disease of rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 7(4):323–326, July 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hsu:1985:CDM

[HL85] Y. L. Hsu and J. C. Leong. A comparison of detection methods for infectious haematopoietic necrosis virus. Journal of Fish Diseases, 8(1):1–12, January 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Humphrey:1987:DSA

[HLGC87] J. D. Humphrey, C. E. Lancaster, N. Gudkovs, and J. W. Copland. The disease status of Australian salmonids: bacteria and bacterial diseases. Journal of Fish Diseases, 10(5):403–410, September 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hussein:1982:PCD

[HM82] S. A. Hussein and D. H. Mills. The prevalence of ‘cauliflower disease’ of the eel, Anguilla anguilla L., in tributaries of the River Tweed, Scotland. Journal of Fish Diseases, 5(2):161–165, March 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Haensly:1982:HPP

[HNS^{+}82] W. E. Haensly, J. M. Neff, J. R. Sharp, A. C. Morris, M. F. Bedgood, and P. D. Boem. Histopathology of Pleuronectes platessa L. from Aber Wrac’h and Aber Benoit, Brittany, France: long-term effects of the Amoco Cadiz crude oil spill. Journal of Fish Diseases, 5(5):365–391, September 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hoff:1989:OFP

[Hof89] K. A. Hoff. Occurrence of the fish pathogen Vibrio salmonicida in faeces from Atlantic salmon, Salmo salar L., after experimental infection. Journal of Fish Diseases, 12(6):595–597, November 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hogans:1989:MCA

[Hog89] W. E. Hogans. Mortality of cultured Atlantic salmon, Salmo salar L., parr caused by an infection of Ergasilus labracis (Copepoda: Poecilostomatoida) in the lower Saint John River, New Brunswick, Canada. Journal of Fish Diseases, 12(5):529–531, September 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Horsch:1987:CHW

[Hor87] C. M. Horsch. A case history of whirling disease in a drainage system: Battle Creek drainage of the upper Sacramento River Basin, California, USA. Journal of Fish Diseases, 10(6):453–460, November 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Haller:1980:DNS

[HR80] R. D. Haller and R. J. Roberts. Dual neoplasia in a specimen of Sarotherodon spilurus spilurus (Günther) (= Tilapia spilurus). Journal of Fish Diseases, 3(1):63–66, January 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Horne:1984:EUP

[HRTW84] M. T. Horne, R. J. Roberts, Mary Tatner, and P. Ward. The effects of the use of potassium alum adjuvant in vaccines against vibriosis in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 7(2):91–99, March 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Holloway:1982:MND

[HS82] H. L. Holloway, Jr. and C. E. Smith. A myopathy in North Dakota walleye, Stizostedion vitreum (Mitchill). Journal of Fish Diseases, 5(6):527–530, November 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Harrell:1985:KTG

[HS85] L. W. Harrell and T. M. Scott. Kudoa thyrsitis (Gilchrist) (Myxosporea: Multivalvulida) in Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 8(3):329–332, May 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hjelmeland:1988:ICS

[HSJE88] K. Hjelmeland, K. Stensvåg, T. Jørgensen, and S. Espelid. Isolation and characterization of a surface layer antigen from Vibrio salmonicida. Journal of Fish Diseases, 11(3):197–205, May 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hawkins:1981:UEC

[HSO81] W. E. Hawkins, M. A. Solangi, and R. M. Overstreet. Ultrastructural effects of the coccidium, Eimeria funduli Duszynski, Solangi and Overstreet, 1979 on the liver of killifishes. Journal of Fish Diseases, 4(4):281–295, July 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hawkins:1983:UME

[HSO83a] W. E. Hawkins, M. A. Solangi, and R. M. Overstreet. Ultrastructure of the macrogamont of Eimeria funduli, a coccidium parasitizing killifishes. Journal of Fish Diseases, 6(1):33–43, January 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hawkins:1983:UMM

[HSO83b] W. E. Hawkins, M. A. Solangi, and R. M. Overstreet. Ultrastructure of the microgamont and microgamete of Eimeria funduli, a coccidium parasitizing killifishes. Journal of Fish Diseases, 6(1):

45–57, January 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Hoffmaster:1988:GDM

[HSR^{+}88] J. L. Hoffmaster, J. E. Sanders, J. S. Rohovec, J. L. Fryer, and D. G. Stevens. Geographic distribution of the myxosporean parasite, Ceratomyxa shasta Noble, 1950, in the Columbia River Basin, USA. Journal of Fish Diseases, 11(1):97–100, January 1988. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Horne:1982:VRT

[HTM^{+}82] M. T. Horne, M. Tatner, S. Mcderment, C. Agius, and P. Ward. Vaccination of rainbow trout, Salmo gairdneri Richardson, at low temperatures and the long-term persistence of protection. Journal of Fish Diseases, 5(4):343–345, July 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Henken:1988:EFL

[HTV88] A. M. Henken, A. J. Tigchelaar, and W. B. Van Muiswinkel. Effects of feeding level on antibody production in African catfish, Clarias gariepinus Burchell, after injection of Yersinia ruckeri O-antigen. Journal of Fish Diseases, 11(1):85–88, January 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ishiguro:1984:SRA

[IIK^{+}84] S. Ishiguro, H. Izawa, H. Kodama, M. Onuma, and T. Mikami. Serological relationships among five strains of infectious pancreatic necrosis virus. Journal of Fish Diseases, 7(2):127–135, March 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Iversen:1987:UTD

[IKA87] E. S. Iversen, J. F. Kelly, and D. Alzamora. Ultrastructure of Thelohania duorara Iversen & Manning, 1959 (Microspora, Thelohaniidae) in the Pink Shrimp, Penaeus duroarum Burkenroad. Journal of Fish Diseases, 10(4):299–307, July 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Johnson:1983:CES

[JA83a] K. A. Johnson and D. F. Amend. Comparison of efficacy of several delivery methods using Yersinia ruckeri bacterin on rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 6(4):331– 336, July 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Johnson:1983:EVA

[JA83b] K. A. Johnson and D. F. Amend. Efficacy of Vibrio anguillarum and Yersinia ruckeri bacterins applied by oral and anal intubation of salmonids. Journal of Fish Diseases, 6(5):473–476, September 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Johnson:1984:PIV

[JA84] K. A. Johnson and D. F. Amend. Potential for immersion vaccination against Aeromonas salmonicida. Journal of Fish Diseases, 7(2):101–105, March 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jacobsen:1989:WTF

[Jac89] M. D. Jacobsen. Withdrawal times of freshwater rainbow trout, Salmo gairdneri Richardson, after treatment with oxolinic acid, oxytetracycline and trimetoprim. Journal of Fish Diseases, 12(1): 29–36, January 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jilek:1982:IHC

[JC82] R. Jilek and J. L. Crites. Intestinal histopathology of the common bluegill, Lepomis macrochirus Rafinesque, infected with Spinitectus carolini Holl, 1928 (Spirurida: Nematoda). Journal of Fish Diseases, 5(1):75–77, January 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Johnson:1982:DIS

[JFA82a] K. A. Johnson, J. K. Flynn, and D. F. Amend. Duration of immunity in salmonids vaccinated by direct immersion with Yersinia ruckeri and Vibrio anguillarum bacterins. Journal of Fish Diseases, 5(3):207–213, May 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Johnson:1982:OIS

[JFA82b] K. A. Johnson, J. K. Flynn, and D. F. Amend. Onset of immunity in salmonid fry vaccinated by direct immersion in Vibrio anguillarum and Yersinia ruckeri bacterins. Journal of Fish Diseases, 5 (3):197–205, May 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Johnsen:1988:IEG

[JJ88] B. O. Johnsen and A. J. Jensen. Introduction and establishment of Gyrodactylus salaris Malmberg, 1957, on Atlantic salmon, Salmo salar L., fry and parr in the River Vefsna, northern Norway. Journal of Fish Diseases, 11(1):35–45, January 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jorgensen:1982:EVO

[Jør82a] P. E. Vestergård Jørgensen. Egtved virus: occurrence of inapparent infections with virulent virus in free-living rainbow trout, Salmo gairdneri Richardson, at low temperature. Journal of Fish Diseases, 5(3):251–255, May 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jorgensen:1982:EVT

[Jør82b] P. E. Vestergård Jørgensen. Egtved virus: temperature-dependent immune response of trout to infection with low-virulence virus. Journal of Fish Diseases, 5(1):47–55, January 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Johansson:1982:SPU

[JSF82] N. Johansson, K. M. Svensson, and G. Fridberg. Studies on the pathology of ulcerative dermal necrosis (UDN) in Swedish salmon, Salmo salar L., and sea trout, Salmo trutta L., populations. Journal of Fish Diseases, 5(4):293–308, July 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jones:1987:IRR

[JW87] S. R. M. Jones and P. T. K. Woo. The immune response of rainbow trout, Salmo gairdneri Richardson, to the haemoflagellate, Cryptobia salmositica Katz, 1951. Journal of Fish Diseases, 10 (5):395–402, September 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Jones:1986:IRT

[JWS86] S. R. M. Jones, P. T. K. Woo, and R. M. W. Stevenson. Immunosuppression in rainbow trout, Salmo gairdneri Richardson, caused by the haemoflagellate Cryptobia salmositica Katz, 1951. Journal of Fish Diseases, 9(5):431–438, September 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kitao:1983:SVA

[KAF^{+}83] T. Kitao, T. Aoki, M. Fukudome, K. Kawano, Yo. Wada, and Y. Mizuno. Serotyping of Vibrio anguillarum isolated from diseased freshwater fish in Japan. Journal of Fish Diseases, 6(2): 175–181, March 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Krise:1989:TLT

[KH89] W. F. Krise and R. L. Herman. Tolerance of lake trout, Salvelinus namaycush (Walbaum), sac fry to dissolved gas supersaturation. Journal of Fish Diseases, 12(3):269–273, May 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kirmse:1980:OPH

[Kir80] P. Kirmse. Observations on the pathogenicity of Haemogregarina sachai Kirmse, 1978, in farmed turbot Scophthalmus maximus (L.). Journal of Fish Diseases, 3(2):101–114, March 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kimura:1982:SPS

[KK82] H. Kimura and R. Kusuda. Studies on the pathogenesis of streptococcal infection in cultured yellowtails, Seriola spp.: effect of crude exotoxin fractions from cell-free culture on experimental streptococcal infection. Journal of Fish Diseases, 5(6):471–478, November 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kent:1989:CPA

[KMF89] M. L. Kent, M. Moser, and J. W. Fournie. Coccidian parasites (Apicomplexa: Eucoccidorida) in hardy head fish, Atherinomorus capricornensis (Woodland). Journal of Fish Diseases, 12(2):179– 183, March 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kimura:1983:EOR

[KMKK83] T. Kimura, S. Morikawa, T. Kiriyama, and H. Kitaori. An epizootic occurrence of rhabdomyoma and a case of ganglioneuroma in hatchery-reared ayu, Plecoglossus altivelis Temminck & Schlegel. Journal of Fish Diseases, 6(2):195–200, March 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kakuta:1989:UPC

[KN89] I. Kakuta and K. Namba. Urine properties in carp, Cyprinus carpio L., with sekoke disease. Journal of Fish Diseases, 12(3):241–248, May 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Karlsson-Norrgren:1986:AWAb

[KNBLR86] L. Karlsson-Norrgren, I. Björklund, O. Ljungberg, and P. Runn. Acid water and aluminium exposure: experimentally induced gill lesions in brown trout, Salmo trutta L. Journal of Fish Diseases, 9(1):11–25, January 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Karlsson-Norrgren:1986:AWAa

[KNDLR86] L. Karlsson-Norrgren, W. Dickson, O. Ljungberg, and P. Runn. Acid water and aluminium exposure: gill lesions and aluminium accumulation in farmed brown trout, Salmo trutta L. Journal of Fish Diseases, 9(1):1–9, January 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Knittel:1981:SST

[Kni81] M. D. Knittel. Susceptibility of steelhead trout Salmo gairdneri Richardson to redmouth infection Yersinia ruckeri following exposure to copper. Journal of Fish Diseases, 4(1):33–40, January 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kelly:1983:CHV

[KNMY83] R. K. Kelly, O. Nielsen, S. C. Mitchell, and T. Yamamoto. Characterization of Herpes virus vitreum isolated from hyperplastic epidermal tissue of walleye, Stizostedion vitreum vitreum (Mitchill). Journal of Fish Diseases, 6(3):249–260, May 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kodama:1987:SAI

[KNY^{+}87] H. Kodama, Y. Nakanishi, F. Yamamoto, T. Mikami, H. Izawa, T. Imagawa, Y. Hashimoto, and N. Kudo. Salmonella arizonae isolated from a pirarucu, Arapaima gigas Cuvier, with septicaemia. Journal of Fish Diseases, 10(6):509–512, November 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kubota:1981:USI

[KOA81] S. S. Kubota, T. Ochiai, and S. Akeda. Ultrastructural study of inclusion-like bodies in hepatocytes of Cyprinus carpio L. induced by the herbicide Benthiocarb. Journal of Fish Diseases, 4(1):77– 81, January 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kranz:1985:PCL

[KP85] Heidemarie Kranz and N. Peters. Pathological conditions in the liver of ruffe, Gymnocephalus cernua (L.), from the Elbe estuary. Journal of Fish Diseases, 8(1):13–24, January 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kranz:1980:CKS

[KPBS80] Heidemarie Kranz, N. Peters, Gisela Bresching, and H. F. Stich. On cell kinetics in skin tumours of the Pacific English sole Parophrys vetulus Girard. Journal of Fish Diseases, 3(2):125–132, March 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kasornchandra:1987:EIW

[KRP87] J. Kasornchandra, W. A. Rogers, and J. A. Plumb. Edwardsiella ictaluri from walking catfish, Clarias batrachus L., in Thailand. Journal of Fish Diseases, 10(2):137–138, March 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kumar:1986:SMC

[KSDM86] D. Kumar, K. Suresh, R. K. Dey, and B. K. Mishra. Stressmediated columnaris disease in rohu, Labeo rohita (Hamilton). Journal of Fish Diseases, 9(1):87–89, January 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kent:1987:CNA

[KWYE87] M. L. Kent, S. R. Wellings, W. T. Yasutake, and R. A. Elston. Cranial nodules associated with cranial fenestrae in juvenile Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 10(5):419–421, September 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

KanekoII:1988:ITO

[KYBN88] J. J. Kaneko I.I, R. Yamada, J. A. Brock, and R. M. Nakamura. Infection of tilapia, Oreochromis mossambicus (Trewavas), by a marine monogenean, Neobenedenia melleni (MacCallum, 1927) Yamaguti, 1963 in Kaneohe Bay, Hawaii, USA, and its treatment. Journal of Fish Diseases, 11(4):295–300, July 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Kimura:1983:VAA

[KYW83] T. Kimura, M. Yoshimizu, and M. Wada. In vitro antibacterial activity of the combination of sulphadiazine and trimethoprim on bacterial fish pathogens. Journal of Fish Diseases, 6(6):525–532, November 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lom:1989:GPS

[LA89] J. Lom and J. R. Arthur. A guideline for the preparation of species descriptions in Myxosporea. Journal of Fish Diseases, 12(2):151– 156, March 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Laidler:1980:DIB

[Lai80] L. A. Laidler. Detection and identification of the bacterial kidney disease (BKD) organism by the indirect fluorescent antibody technique. Journal of Fish Diseases, 3(1):67–69, January 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Langdon:1986:IIU

[Lan86] J. S. Langdon. Intestinal infection with a unicellular green alga in the golden perch, Macquaria ambigua (Richardson). Journal of Fish Diseases, 9(2):159–162, March 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Langdon:1987:SCE

[Lan87] J. S. Langdon. Spinal curvatures and an encephalotropic myxosporean, Triangula percae sp. nov. (Myxozoa: Ortholineidae), enzootic in redfin perch, Perca fluviatilis L., in Australia. Journal of Fish Diseases, 10(6):425–434, November 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Langdon:1989:ETP

[Lan89] J. S. Langdon. Experimental transmission and pathogenicity of epizootic haematopoietic necrosis virus (EHNV) in redfin perch, Perca fluviatilis L., and 11 other teleosts. Journal of Fish Diseases, 12 (4):295–310, July 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lom:1982:SMF

[LD82] J. Lom and I. Dyková. Some marine fish coccidia of the genera Epiemeria Schneider, Epieimeria Dyková & Lom and Goussia Labbé. Journal of Fish Diseases, 5(4):309–321, July 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lallier:1984:ADP

[LD84] R. Lallier and P. Daigneault. Antigenic differentiation of pili from non-virulent and fish-pathogenic strains of Aeromonas hydrophila. Journal of Fish Diseases, 7(6):509–512, November 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lamers:1985:RIS

[LdHvM85] C. H. J. Lamers, M. J. H. de Haas, and W. B. van Muiswinkel. The reaction of the immune system of fish to vaccination: development of immunological memory in carp, Cyprinus carpio L., following direct immersion in Aeromonas hydrophila bacterin. Journal of Fish Diseases, 8(3):253–262, May 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Leibovitz:1989:CNR

[Lei89] L. Leibovitz. Chlamydiosis: a newly reported serious disease of larval and postmetamorphic bay scallops, Argopecten irradians (Lamarck). Journal of Fish Diseases, 12(2):125–136, March 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Levine:1984:TPC

[Lev84] N. D. Levine. The taxonomic position of the coccidian genus Crystallospora (Protozoa, Apicomplexa). Journal of Fish Diseases, 7 (4):317–318, July 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lewis:1985:VCC

[Lew85] D. H. Lewis. Vibriosis in channel catfish, Ictalurus punctatus (Rafinesque). Journal of Fish Diseases, 8(6):539–545, November 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lewis:1986:ELI

[Lew86] D. H. Lewis. An enzyme-linked immunosorbent assay (ELISA) for detecting penaeid baculovirus. Journal of Fish Diseases, 9(6):519– 522, November 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lom:1989:BMB

[LFDK89] J. Lom, S. W. Feist, I. Dyková, and T. Kepr. Brain myxoboliasis of bullhead, Cottus gobio L., due to Myxobolus jiroveci sp. nov.: light and electron microscope observations. Journal of Fish Diseases, 12(1):15–27, January 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Leong:1981:EPI

[LFYL81] J. C. Leong, J. L. Fendrick, S. Youngman, and A. Lee. Effect of polybrene on the infectivity of infectious haemopoietic necrosis virus in tissue culture. Journal of Fish Diseases, 4(4):335–344, July 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Langdon:1987:EHN

[LH87] J. S. Langdon and J. D. Humphrey. Epizootic haematopoietic necrosis, a new viral disease in redfin perch, Perca fluviatilis L., in Australia. Journal of Fish Diseases, 10(4):289–297, July 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Langdon:1986:DSA

[LHC^{+}86] J. S. Langdon, J. D. Humphrey, J. Copland, R. Carolane, N. Gudkovs, and C. Lancaster. The disease status of Australian salmonids: viruses and viral diseases. Journal of Fish Diseases, 9(2):129–135, March 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Langdon:1986:FVI

[LHW^{+}86] J. S. Langdon, J. D. Humphrey, L. M. Williams, A. D. Hyatt, and H. A. Westbury. First virus isolation from Australian fish: an iridovirus-like pathogen from redfin perch, Perca fluviatilis L. Journal of Fish Diseases, 9(3):263–268, May 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Langdon:1988:OEL

[LHW88] J. S. Langdon, J. D. Humphrey, and L. M. Williams. Outbreaks of an EHNV-like iridovirus in cultured rainbow trout, Salmo gairdneri Richardson, in Australia. Journal of Fish Diseases, 11(1):93–96, January 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lillehaug:1989:OIR

[Lil89] A. Lillehaug. Oral immunization of rainbow trout, Salmo gairdneri Richardson, against vibriosis with vaccines protected against digestive degradation. Journal of Fish Diseases, 12(6):579–584, November 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lester:1984:ASG

[LK84] R. J. G. Lester and W. R. Kelly. Ankylosing spondylosis in the giant perch, Lates calcarifer (Bloch). Journal of Fish Diseases, 7 (3):193–197, May 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Leibovitz:1985:CAD

[LL85a] L. Leibovitz and S. S. Lebouitz. A coccolithophorid algal dermatitis of the spiny dogfish, Squalus acanthias L. Journal of Fish Diseases, 8(4):351–358, July 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Leibovitz:1985:VDS

[LL85b] L. Leibovitz and S. S. Lebouitz. A viral dermatitis of the smooth dogfish, Mustelus canis (Mitchill). Journal of Fish Diseases, 8(3): 273–279, May 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Llewellyn:1980:BSR

[Lle80] L. C. Llewellyn. A bacterium with similarities to the redmouth bacterium and Serratia liquefaciens (Grimes and Hennerty) causing mortalities in hatchery reared salmonids in Australia. Journal of Fish Diseases, 3(1):29–39, January 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lesel:1983:OER

[LLGV83] R. Lesel, M. Lesel, F. Gavini, and A. Vuillaume. Outbreak of enteric redmouth disease in rainbow trout, Salmo gairdneri Richardson, in France. Journal of Fish Diseases, 6(4):385–387, July 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lewis:1985:MSM

[LM85] D. H. Lewis and J. E. Marks. Microcultures of Sarotherodon mossambicus (Peters) cells: their use in detecting fish viruses.

Journal of Fish Diseases, 8(5):477–478, September 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lewis:1985:DMC

[LMS85] D. H. Lewis, J. E. Marks, and R. R. Stickney. Degenerative myopathy in channel catfish, Ictalurus punctatus (Rafinesque), maintained on rations containing purified fatty acids. Journal of Fish Diseases, 8(6):563–565, November 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lom:1980:CBN

[Lom80] J. Lom. Cryptobia branchialis Nie from fish gills: ultrastructural evidence of ectocommensal function. Journal of Fish Diseases, 3 (5):427–436, September 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lio-Po:1985:VEF

[LPBLS85] G. D. Lio-Po, M. C. L. Baticados, C. R. Lavilla, and M. E. G. Sanvictores. In vitro effects of fungicides on Haliphthoros philippinensis. Journal of Fish Diseases, 8(4):359–365, July 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lom:1985:NKI

[LPD85] Jiří Lom, Milena Pavlásková, and Iva Dyková. Notes on kidneyinfecting species of the genus Sphaerospora Thélohan (Myxosporea), including a new species S. gobionis sp.nov., and on myxosporean life cycle stages in the blood of some freshwater fish. Journal of Fish Diseases, 8(2):221–232, March 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lall:1985:CBK

[LPHA85] S. P. Lall, W. D. Paterson, J. A. Hines, and N. J. Adams. Control of bacterial kidney disease in Atlantic salmon, Salmo salar L., by dietary modification. Journal of Fish Diseases, 8(1):113–124, January 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lio-Po:1986:BAI

[LPPM86] Gild A. Lio-Po, Celia Pitogo, and Clarissa Marte. Bacteria associated with infection at hormone-implantation sites among milkfish, Chanos chanos (Forsskål), adults. Journal of Fish Diseases, 9(4): 337–343, July 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lio-Po:1982:VEF

[LPSBL82] G. D. Lio-Po, M. E. G. Sanvictores, M. C. L. Baticados, and C. R. Lavilla. In vitro effect of fungicides on hyphal growth and sporogenesis of Lagenidium spp. isolated from Penaeus monodon larvae and Scylla serrata eggs. Journal of Fish Diseases, 5(2):97–112, March 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lopez:1981:MCF

[LR81] A. Lopez and A. Raibaut. Multiple cutaneous fibromas in a mullet, Mugil cephalus cephalus L. Journal of Fish Diseases, 4(2):169–174, March 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lightner:1985:NHP

[LR85] D. V. Lightner and R. M. Redman. Necrosis of the hepatopancreas in Penaeus monodon and P. stylirostris (Arthropoda, Decapoda) with red disease. Journal of Fish Diseases, 8(2):181–188, March 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lightner:1986:PMS

[LR86] D. V. Lightner and R. M. Redman. A probable Mycobacterium sp. infection of the marine shrimp Penaeus vannamei (Crustacea: Decapoda). Journal of Fish Diseases, 9(4):357–359, July 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lightner:1984:IPD

[LRBB84] D. V. Lightner, R. M. Redman, T. A. Bell, and J. A. Brock. An idiopathic proliferative disease syndrome of the midgut and ventral nerve in the kuruma prawn, Penaeus japonicus Bate, cultured in Hawaii. Journal of Fish Diseases, 7(3):183–191, May 1984. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lightner:1988:RMA

[LRM^{+}88] D. Lightner, R. M. Redman, L. Mohney, J. Sinski, and D. Priest. A renal mycosis of an adult hybrid red tilapia, Oreochromis mossambicus X O. hornorum, caused by the imperfect fungus, Paecilomyces marquandii. Journal of Fish Diseases, 11(5):437–440, September 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Leatherland:1980:EDM

[LS80] J. F. Leatherland and R. A. Sonstegard. Effect of dietary Mirex and PCB’s in combination with food deprivation and testosterone administration on thyroid activity and bioaccumulation of organochlorines in rainbow trout Salmo gairdneri Richardson. Journal of Fish Diseases, 3(2):115–124, March 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Leatherland:1981:TDG

[LS81] J. F. Leatherland and R. A. Sonstegard. Thyroid dysfunction in Great Lakes Coho salmon, Oncorhynchus kisutch (Walbaum): seasonal and interlake differences in serum T3 uptake and serum total and free T4 and T3 levels. Journal of Fish Diseases, 4(5):413– 423, September 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Lom:1983:USP

[LS83] J. Lom and G. Schubert. Ultrastructural study of Pisdnoodinium pillulare (Schäperclaus, 1954) Lom, 1981 with special emphasis on its attachment to the fish host. Journal of Fish Diseases, 6(5):411– 428, September 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mangunwiryo:1988:SCS

[MA88] H. Mangunwiryo and C. Agius. Studies on the carrier state of infectious pancreatic necrosis virus infections in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 11(2):125–132, March 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mackenzie:1981:EES

[Mac81] K. Mackenzie. The effect of Eimeria sp. infection on the condition of blue whiting, Micromesistius poutassou (Risso). Journal of Fish Diseases, 4(6):473–486, November 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mccarthy:1983:ASD

[MAJB83] D. H. Mccarthy, D. F. Amend, K. A. Johnson, and Joan V. Bloom. Aeromonas salmonicida: determination of an antigen associated with protective immunity and evaluation of an experimental bacterin. Journal of Fish Diseases, 6(2):155–174, March 1983. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Markiw:1989:SWD

[Mar89] M. E. Markiw. Salmonid whirling disease: myxosporean and actinosporean stages cross-react in direct fluorescent antibody test. Journal of Fish Diseases, 12(2):137–141, March 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mater:1984:GIT

[Mat84] K. J. Mater. Gallstone induction in tilapia, Oreochromis mossambicus (Peters). Journal of Fish Diseases, 7(6):521–524, November 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Molnar:1981:LEM

[MB81] K. Molnár and G. Boros. A light and electron microscopic study of the agent of carp mucophilosis. Journal of Fish Diseases, 4(4): 325–334, July 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mulcahy:1983:ELD

[MB83] D. Mulcahy and K. Bauersfeld. Effect of loading density of sockeye salmon, Oncorhynchm nerka (Walbaum), eggs in incubation boxes on mortality caused by infectious haematopoietic necrosis. Journal of Fish Diseases, 6(2):189–193, March 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mchenery:1986:IFM

[MB86a] J. G. Mchenery and T. H. Birkbeck. Inhibition of filtration in Mytilus edulis L. by marine vibrios. Journal of Fish Diseases, 9 (3):257–261, May 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Molnar:1986:LEM

[MB86b] K. Molnár and F. Baska. Light and electron microscopic studies on Epieimeria anguillae (Léger & Hollande, 1922), a coccidium parasitizing the European eel, Anguilla anguilla L. Journal of Fish Diseases, 9(2):99–110, March 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mcardle:1987:SUR

[MB87] J. Mcardle and A. M. Bullock. Solar ultraviolet radiation as a causal factor of ‘summer syndrome’ in cage-reared Atlantic salmon, Salmo solar L.: a clinical and histopathological study. Journal of Fish Diseases, 10(4):255–264, July 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Maccubbin:1987:CRH

[MBH87] A. E. Maccubbin, J. Black, and J. C. Harshbarger. A case report of hepatocellular neoplasia in bowfin, Amia calva L. Journal of Fish Diseases, 10(4):329–331, July 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Morrison:1986:PMF

[MC86] Carol M. Morrison and D. K. Cone. A possible marine form of Ichthyobodo sp. on haddock, Melanogrammus aeglefinus (L.), in the north-west Atlantic Ocean. Journal of Fish Diseases, 9(2): 141–142, March 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mccarthy:1984:IRT

[MCA84] D. H. Mccarthy, T. R. Croy, and D. F. Amend. Immunization of rainbow trout, Salmo gairdneri Richardson, against bacterial kidney disease: preliminary efficacy evaluation. Journal of Fish Diseases, 7(1):65–71, January 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mccarthy:1983:EMF

[Mcc83] D. H. Mccarthy. An experimental model for fish furunculosis caused by Aeromonas salmonicida. Journal of Fish Diseases, 6 (3):231–237, May 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Morrison:1981:MHS

[MCSZ81] Carol Morrison, J. Cornick, G. Shum, and B. Zwicker. Microbiology and histopathology of ‘saddleback’ disease of underyearling Atlantic salmon, Salmo salar L. Journal of Fish Diseases, 4(3): 243–258, May 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Morrison:1984:HAA

[MCSZ84] C. M. Morrison, J. W. Cornick, G. Shum, and E. Zwicker. Histopathology of atypical Aeromonas salmonicida infection in Atlantic cod, Gadus morhua L. Journal of Fish Diseases, 7(6):477– 494, November 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mcvicar:1987:BPN

[Mcv87] A. H. Mcvicar. Black patch necrosis of the skin of Solea solea (L.): the role of sand in prophylaxis and treatment. Journal of Fish Diseases, 10(1):59–63, January 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Murchelano:1981:EOC

[ME81] R. A. Murchelano and R. L. Edwards. An erythrophoroma in ornamental carp, Cyprinus carpio L. Journal of Fish Diseases, 4 (3):265–269, May 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Meyers:1980:EPR

[Mey80] T. R. Meyers. Experimental pathogenicity of reovirus 13p_{2} for juvenile American oysters Crassostrea virginica (Gmelin) and bluegill fingerlings Lepomis macrochirus (Rafinesque). Journal of Fish Diseases, 3(3):187–201, May 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Meyers:1983:SHR

[Mey83] T. R. Meyers. Serological and histopathological responses of rainbow trout, Salmo gairdneri Richardson, to experimental infection with the 13_{P 2} reovirus. Journal of Fish Diseases, 6(3):277–292, May 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Michel:1982:OSA

[MF82] C. Michel and Bernadette Faivre. Occurrence and significance of agglutinating antibodies in experimental furunculosis of rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 5 (5):429–432, September 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Majeed:1983:CTC

[MG83] S. K. Majeed and C. Gopinath. Cutaneous tuberculosis in the carp, Cyprinus carpio L. Journal of Fish Diseases, 6(3):313–316, May 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mitchell:1983:MVF

[MGB83] L. G. Mitchell, J. Ginal, and W. C. Bailey. Melanotic visceral fibrosis associated with larval infections of Posthodiplostomum minimum and Proteocephalus sp. in bluegill, Lepomis macrochirus Rafinesque, in central Iowa, U.S.A. Journal of Fish Diseases, 6(2): 135–144, March 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Majeed:1981:PST

[MGJ81] S. K. Majeed, C. Gopinath, and D. W. Jolly. Pathology of spontaneous tuberculosis and pseudotuberculosis in fish. Journal of Fish Diseases, 4(6):507–512, November 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Meyers:1983:HFS

[MH83] T. R. Meyers and J. D. Hendricks. Histopathology of four spontaneous neoplasms in three species of salmonid fishes. Journal of Fish Diseases, 6(6):481–499, November 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Meyers:1986:FRV

[MHBM86] T. R. Meyers, A. K. Hauck, W. D. Blankenbeckler, and T. Minicucci. First report of viral erythrocytic necrosis in Alaska, USA, associated with epizootic mortality in Pacific herring, Clupea harengus pallasi (Valenciennes). Journal of Fish Diseases, 9 (6):479–491, November 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Moccia:1984:EOF

[MHSF84] R. D. Moccia, S. S. O. Hung, S. J. Slinger, and H. W. Ferguson. Effect of oxidized fish oil, vitamin E and ethoxyquin on the histopathology and haematology of rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 7(4):269–282, July 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Minchew:1981:UMS

[Min81] C. D. Minchew. Unicauda magna sp.n. (Protozoa: Myxozoa): a new myxozoan from the fin tissues of the fathead minnow, Pimephales promelas Rafinesque. Journal of Fish Diseases, 4 (6):513–518, November 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Masahito:1984:SSS

[MIT84] Prince Masahito, T. Ishikawa, and S. Takayama. Spontaneous spermatocytic seminoma in African lungfish, Protopterus aethiopicus Heckel. Journal of Fish Diseases, 7(2):169–172, March 1984. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mix:1983:HNB

[Mix83] M. C. Mix. Haemic neoplasms of bay mussels, Mytilus edulis L., from Oregon: occurrence, prevalence, seasonality and histopathological progression. Journal of Fish Diseases, 6(3):239–248, May 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Miyazaki:1987:HSR

[Miy87] T. Miyazaki. A histological study of the response to challenge with vibriosis in ayu, Plecoglossus altivelis Temminck and Schlegel, vaccinated by immersion and injection with Vibrio anguillarum bacterin. Journal of Fish Diseases, 10(6):445–452, November 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mccarthy:1982:SSC

[MJ82] D. H. Mccarthy and K. A. Johnson. A serotypic survey and crossprotection test of North American field isolates of Yersinia ruckeri. Journal of Fish Diseases, 5(4):323–328, July 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Majeed:1984:OLC

[MJG84] S. K. Majeed, D. W. Jolly, and C. Gopinath. An outbreak of liver cell carcinoma in rainbow trout, Salmo gairdneri Richardson, in the U.K. Journal of Fish Diseases, 7(2):165–168, March 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mishra:1983:CHC

[MK83] B. K. Mishra and D. Kumar. A case of hernia in Catla catla (Hamilton). Journal of Fish Diseases, 6(6):549, November 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Markwardt:1989:EFM

[MK89a] N. M. Markwardt and G. W. Klontz. Evaluation of four methods to establish asymptomatic carriers of Aeromonas salmonicida in juvenile spring Chinook salmon, Oncorhynchus tshawytscha (Walbaum). Journal of Fish Diseases, 12(4):311–315, July 1989. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Markwardt:1989:MEA

[MK89b] N. M. Markwardt and G. W. Klontz. A method to eliminate the asymptomatic carrier state of Aeromonas salmonicida in salmonids. Journal of Fish Diseases, 12(4):317–322, July 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mitchell:1983:UPC

[MKS83] L. G. Mitchell, S. E. Krall, and C. L. Seymour. Use of the plankton centrifuge to diagnose and monitor prevalence of myxobolid infections in fathead minnows, Pimephales promelas Rafinesque. Journal of Fish Diseases, 6(4):379–384, July 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Munkittrick:1984:APG

[ML84] K. R. Munkittrick and J. F. Leatherland. Abnormal pituitarygonad function in two feral populations of goldfish, Carassius auvatus (L.), suffering epizootics of an ulcerative disease. Journal of Fish Diseases, 7(6):433–447, November 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Marquis:1989:ESP

[ML89] H. Marquis and R. Lallier. Efficacy studies of passive immunization against Aeromonas salmonicida infection in brook trout, Salvelinus fontinalis (Mitchill). Journal of Fish Diseases, 12(3):233–240, May 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Maurand:1988:PMV

[MLG^{+}88] J. Maurand, C. Loubes, C. Gasc, J. Pelletier, and J. Barral. Pleistophora mirandellae Vaney & Conte, 1901, a microsporidian parasite in cyprinid fish of rivers in Herault: taxonomy and histopathology. Journal of Fish Diseases, 11(3):251–258, May 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Moore:1988:IPL

[MLM88] A. R. Moore, M. F. Li, and M. Mcmenemy. Isolation of a picornalike virus from smelt, Osmerus mordax (Mitchill). Journal of Fish Diseases, 11(2):179–184, March 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Matthews:1980:CTR

[MM80] R. A. Matthews and B. F. Matthews. Cell and tissue reactions of turbot Scophthalmus maximus (L.) to Tetramicra brevifilum gen. n., sp. n. (Microspora). Journal of Fish Diseases, 3(6):495–515, November 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Meyers:1985:CGR

[MM85] T. R. Meyers and B. P. Mcpherson. Chronic glomerulonephritis in returning fall Chinook salmon, Oncorhynchus tshawytscha (Walbaum), from the Rogue River, Oregon. Journal of Fish Diseases, 8(5):407–416, September 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Morrison:1986:FOL

[MM86] Carol M. Morrison and Vivian Marryatt. Further observations on Loma morhua Morrison & Sprague, 1981. Journal of Fish Diseases, 9(1):63–67, January 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Murad:1988:BPC

[MM88] A. Murad and S. Mustafa. Blood parameters of catfish, Heteropneustes fossilis (Bloch), parasitized by metacercariae of Diplostomulum sp. Journal of Fish Diseases, 11(4):365–368, July 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Modin:1981:MRS

[Mod81] J. C. Modin. Microsporidium rhabdophilia n. sp. from rodlet cells of salmonid fishes. Journal of Fish Diseases, 4(3):203–211, May 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Molnar:1980:RSC

[Mol80] K. Molnár. Renal sphaerosporosis in the common carp Cypnnus carpio L. Journal of Fish Diseases, 3(1):11–19, January 1980. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Molnar:1982:NCG

[Mol82] K. Molnär. Nodular coccidiosis in the gut of the tench, Tinca tinca L. Journal of Fish Diseases, 5(6):461–470, November 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Molnar:1988:PDS

[Mol88] K. Molnár. Presporogonic development of Sphaerospora renicola Dyková & Lorn, 1982, in the swimbladder of the common carp, Cyprinus carpio L. Journal of Fish Diseases, 11(6):489–497, November 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mitchell:1980:TEF

[MP80] A. J. Mitchell and J. A. Plumb. Toxicity and efficacy of Furanace on channel catfish Ictalurus punctatm (Rafinesque) infected experimentally with Aeromonas hydrophila. Journal of Fish Diseases, 3 (2):93–99, March 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Miyazaki:1985:HEI

[MP85a] T. Miyazaki and J. A. Plumb. Histopathology of Edwardsiella ictaluri in channel catfish, Ictalurus punctatus (Rafinesque). Journal of Fish Diseases, 8(4):389–392, July 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mulcahy:1985:VTI

[MP85b] D. Mulcahy and R. J. Pascho. Vertical transmission of infectious haematopoietic necrosis virus in sockeye salmon, Oncorhynchus nerka (Walbaum): isolation of virus from dead eggs and fry. Journal of Fish Diseases, 8(4):393–396, July 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Morrison:1989:CKH

[MP89a] C. Morrison and S. Poynton. A coccidian in the kidney of the haddock, Melanogrammus aeglefinus (L.). Journal of Fish Diseases, 12(6):591–593, November 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Morrison:1989:NSG

[MP89b] C. M. Morrison and S. L. Poynton. A new species of Goussia (Apicomplexa, Coccidia) in the kidney tubules of the cod, Gadus morhua L. Journal of Fish Diseases, 12(6):533–560, November 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mulcahy:1983:DIH

[MPJ83a] D. Mulcahy, R. J. Pascho, and C. K. Jenes. Detection of infectious haematopoietic necrosis virus in river water and demonstration of waterborne transmission. Journal of Fish Diseases, 6(4):321–330, July 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mulcahy:1983:TDP

[MPJ83b] D. Mulcahy, R. J. Pascho, and C. K. Jenes. Titre distribution patterns of infectious haematopoietic necrosis virus in ovarian fluids of hatchery and feral salmon populations. Journal of Fish Diseases, 6(2):183–188, March 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Matthiessen:1982:HCL

[MR82] P. Matthiessen and R. J. Roberts. Histopathological changes in the liver and brain of fish exposed to endosulfan insecticide during tsetse fly control operations in Botswana. Journal of Fish Diseases, 5(2):153–159, March 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mcallister:1984:IPN

[MR84] P. E. Mcallister and Ximena Reyes. Infectious pancreatic necrosis virus: isolation from rainbow trout, Salmo gairdneri Richardson, imported into Chile. Journal of Fish Diseases, 7(4):319–322, July 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Molnar:1988:NCF

[MR88] K. Molnar and K. Rhode. New coccidians from freshwater fishes of Australia. Journal of Fish Diseases, 11(2):161–169, March 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Muhvich:1989:FSI

[MRLB89] A. G. Muhvich, R. Reimschuessel, M. M. Lipsky, and R. O. Bennett. Fusarium solani isolated from newborn bonnethead sharks, Sphyrna tiburo (L.). Journal of Fish Diseases, 12(1):57–62, January 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Manier:1984:CFC

[MRLR84] J. F. Manier, A. Raibaut, A. Lopez, and J.-A. Rioux. A calcified fibroma in the common carp, Cyprinus carpio L. Journal of Fish Diseases, 7(4):283–292, July 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Miyazaki:1988:GHP

[MRS88] T. Miyazaki, W. A. Rogers, and K. J. Semmens. Gastrointestinal histopathology of paddlefish, Polyodon spathula (Walbaum), infected with larval Hysterothylacium dollfusi Schmidt, Leiby & Kritsky, 1974. Journal of Fish Diseases, 11(3):245–250, May 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mix:1980:TCC

[MS80] M. C. Mix and A. K. Sparks. Tanner crab Chionocetes bairdi Rathbun haemocyte classification and an evaluation of using differential counts to measure infection with a fungal disease. Journal of Fish Diseases, 3(4):285–293, July 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Morrison:1981:EMS

[MS81a] Carol M. Morrison and V. Sprague. Electron microscopical study of a new genus and new species of microsporida in the gills of Atlantic cod Gadus morhua L. Journal of Fish Diseases, 4(1): 15–32, January 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Morrison:1981:LEM

[MS81b] Carol M. Morrison and V. Sprague. Light and electron microscope study of microsporida in the gill of haddock, Melanogrammus aeglefinus (L.). Journal of Fish Diseases, 4(2):179–184, March 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Morrison:1981:MPG

[MS81c] Carol M. Morrison and V. Sprague. Microsporidian parasites in the gills of salmonid fishes. Journal of Fish Diseases, 4(5):371– 386, September 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mcarthur:1982:RMM

[MS82a] Carole P. Mcarthur and S. Sengupta. A rapid micro-method for screening eel sera for antibodies against the digenean Telogaster opisthorchis Macfarlane, 1945. Journal of Fish Diseases, 5(1):67– 70, January 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Morrison:1982:CLO

[MS82b] Carol Morrison and G. Shum. Chlamydia-like organisms in the digestive diverticula of the bay scallop, Argopecten irradians (Lmk). Journal of Fish Diseases, 5(3):173–184, May 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Morrison:1983:EAP

[MS83a] Carol Morrison and G. Shum. Epitheliocystis in American plaice, Hippoglossoides platessoides (Fabricius). Journal of Fish Diseases, 6(3):303–308, May 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Morrison:1983:RKB

[MS83b] Carol Morrison and G. Shum. Rickettsias in the kidney of the bay scallop, Argopecten irradians (Lamarck). Journal of Fish Diseases, 6(6):537–541, November 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Morrison:1983:LSP

[MS83c] Carol M. Morrison and V. Sprague. Loma salmonae (Putz, Hoffman and Dunbar, 1965) in the rainbow trout, Salmo gairdneri Richardson, and L. fontinalis sp. nov. (Microsporida) in the brook trout, Salvelinus fontinalis (Mitchill). Journal of Fish Diseases, 6 (4):345–353, July 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mugridge:1983:MEA

[MS83d] R. E. R. Mugridge and H. G. Stallybrass. A mortality of eels, Anguilla anguilla L., attributed to Gnathiidae. Journal of Fish Diseases, 6(1):81–82, January 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Morado:1986:OHP

[MS86] J. F. Morado and A. Sparks. Observations on the host-parasite relations of the Pacific whiting, Merluccius productus (Ayres), and two myxosporean parasites, Kudoa thyrsitis (Gilchrist, 1924) and K. paniformis Kabata & Whitaker, 1981. Journal of Fish Diseases, 9(5):445–455, September 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Momoyama:1988:MEI

[MS88] K. Momoyama and T. Sano. A method of experimental infection of kuruma shrimp larvae, Penaeus japonicus Bate, with baculoviral mid-gut gland necrosis (BMN) virus. Journal of Fish Diseases, 11(2):105–111, March 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Momoyama:1989:DSK

[MS89] K. Momoyama and T. Sano. Developmental stages of kuruma shrimp, Penaeus japonicus Bate, susceptible to baculoviral midgut gland necrosis (BMN) virus. Journal of Fish Diseases, 12 (6):585–589, November 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Moser:1984:IBS

[MSWL84] M. Moser, Judy Sakanari, S. Wellings, and K. Lindstrom. Incompatibility between San Francisco striped bass, Morone saxatilis (Walbaum), and the metacestode, Lacistorhynchus tenuis (Beneden, 1858). Journal of Fish Diseases, 7(5):397–400, September 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Murphy:1981:UCL

[Mur81] T. M. Murphy. The use of chemosterilants to lower the frequency of skin fungal infection amongst precocious male 1+ Atlantic salmon parr, Salmo salar L. Journal of Fish Diseases, 4(5):387–395, September 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Mulcahy:1986:NEI

[MW86] D. Mulcahy and J. Wood. A natural epizootic of infectious haematopoietic necrosis in imported Atlantic salmon, Salmo salar L., reared in the enzootic region. Journal of Fish Diseases, 9(2): 173–175, March 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Meier:1988:VHS

[MW88] W. Meier and T. Wahli. Viral haemorrhagic septicaemia (VHS) in grayling, Thymallus thymallus L. Journal of Fish Diseases, 11 (6):481–487, November 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nash:1988:PCT

[NAS88] G. Nash, I. G. Anderson, and M. Shariff. Pathological changes in the tiger prawn, Penaeus monodon Fabricius, associated with culture in brackishwater ponds developed from potentially acid sulphate soils. Journal of Fish Diseases, 11(2):113–123, March 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nottage:1986:TMB

[NB86] A. S. Nottage and T. H. Birkbeck. Toxicity to marine bivalves of culture supernatant fluids of the bivalve-pathogenic Vibrio strain NCMB 1338 and other marine vibrios. Journal of Fish Diseases, 9 (3):249–256, May 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nottage:1987:PPD

[NB87a] A. S. Nottage and T. H. Birkbeck. Production of proteinase during experimental infection of Ostrea edulis L. larvae with Vibrio alginolyticus NCMB 1339 and the antigenic relationship between proteinases produced by marine vibrios pathogenic for fish and shellfish. Journal of Fish Diseases, 10(4):265–273, July 1987. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nottage:1987:PPP

[NB87b] A. S. Nottage and T. H. Birkbeck. Purification of a proteinase produced by the bivalve pathogen Vibrio alginolyticus NCMB 1339. Journal of Fish Diseases, 10(3):211–220, May 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Graham:1988:NAD

[nCJS88] S. Graham (nee Chung), A. H. Jeffries, and C. J. Secombes. A novel assay to detect macrophage bactericidal activity in fish: factors influencing the killing of Aeromonas salmonicida. Journal of Fish Diseases, 11(5):389–396, September 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nash:1987:IMN

[NCL87] G. Nash, S. Chinabut, and C. Limsuwan. Idiopathic muscle necrosis in the freshwater prawn, Macrobrachium rosenbergii de Man, cultured in Thailand. Journal of Fish Diseases, 10(2):109–120, March 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Noga:1986:OFA

[ND86] E. J. Noga and M. J. Dykstra. Oomycete fungi associated with ulcerative mycosis in menhaden, Brevoortia tyrannus (Latrobe). Journal of Fish Diseases, 9(1):47–53, January 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Neukirch:1984:ESE

[Neu84] M. Neukirch. An experimental study of the entry and multiplication of viral haemorrhagic septicaemia virus in rainbow trout, Salmo gairdneri Richardson, after water-borne infection. Journal of Fish Diseases, 7(3):231–234, May 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Newman:1982:DIV

[NM82] S. G. Newman and J. J. Majnarich. Direct immersion vaccination of juvenile rainbow trout, Salmo gairdneri Richardson, and juvenile Coho salmon, Oncorhynchm kisutch (Walbaum), with a Yersinia ruckeri bacterin. Journal of Fish Diseases, 5(4):339–341, July 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nash:1988:RLV

[NNAS88] M. Nash, G. Nash, I. G. Anderson, and M. Shariff. A reo-like virus observed in the tiger prawn, Penaeus monodon Fabricius, from Malaysia. Journal of Fish Diseases, 11(6):531–535, November 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nasisse:1989:DRC

[NND89] M. P. Nasisse, E. J. Noga, and M. G. Davidson. Degenerative retinopathy in captive Atlantic menhaden, Brevoortia tyrannus (L.). Journal of Fish Diseases, 12(1):37–44, January 1989. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nash:1988:SAC

[NNS88] G. Nash, M. Nash, and H.-J. Schlotfeldt. Systemic amoebiasis in cultured European catfish, Silurus glanis L. Journal of Fish Diseases, 11(1):57–71, January 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Noga:1986:ILC

[Nog86] E. J. Noga. The importance of Lernaea cruciata (Le Sueur) in the initiation of skin lesions in largemouth bass, Micropterus salmoides (Lacépède), in the Chowan River, North Carolina, USA. Journal of Fish Diseases, 9(4):295–302, July 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nash:1985:BOG

[NP85] G. Nash and C. Porter. Branchial osteochondroma in a gilthead sea bream, Sparus aurata L., cultured in the Gulf of Aqaba. Journal of Fish Diseases, 8(3):333–336, May 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nash:1988:EPC

[NRC^{+}88] G. Nash, R. J. Roberts, Supranee Chinabut, Sofa Areerat, and C. Limsuwan. Emaciation of pond-cultured snakehead, Channa striatus (Fowler). Journal of Fish Diseases, 11(3):215–224, May 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nusbaum:1981:ASA

[NS81] K. E. Nusbaum and E. B. Shotts, Jr. Action of selected antibiotics on four common bacteria associated with diseases of fish. Journal of Fish Diseases, 4(5):397–404, September 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Nash:1989:SPD

[NSRS89] G. Nash, P. Southgate, R. H. Richards, and E. Sochon. A systemic protozoal disease of cultured salmonids. Journal of Fish Diseases, 12(2):157–173, March 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Newton:1989:PEE

[NWGP89] J. C. Newton, L. G. Wolfe, J. M. Grizzle, and J. A. Plumb. Pathology of experimental enteric septicaemia in channel catfish, Ictalurus punctatus (Rafinesque), following immersion-exposure to Edwardsiella ictaluri. Journal of Fish Diseases, 12(4):335–347, July 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Olson:1985:SDR

[OBB^{+}85] D. P. Olson, M. H. Beleau, R. A. Busch, S. Roberts, and R. I. Krieger. Strawberry disease in rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 8(1):103–111, January 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Olivier:1985:IVV

[OEL85] G. Olivier, T. P. T. Evelyn, and R. Lallier. Immunogenicity of vaccines from a virulent and an a virulent strain of Aeromonas salmonicida. Journal of Fish Diseases, 8(1):43–55, January 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Olesen:1988:IIL

[OJBM88] N. J. Olesen, P. E. Vestergard Jorgensen, B. Bloch, and S. Mellergaard. Isolation of an IPN-like virus belonging to the serogroup II of the aquatic birnaviruses from dab, Limanda limanda L. Journal of Fish Diseases, 11(5):449–451, September 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Olufemi:1986:AFA

[Olu86] B. E. Olufemi. Application of the fluorescent antibody technique (FAT) to the demonstration of Aspergillus organisms in formalinfixed tissues of tilapias. Journal of Fish Diseases, 9(1):91–93, January 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Overstreet:1986:BGH

[OM86] R. M. Overstreet and F. P. Meyer. Biography: Glenn L. Hoffman. Journal of Fish Diseases, 9(4):291–293, July 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Obradoic:1983:PSF

[OMS83] J. Obradoić, B. Marân, and R. Sabočanec. Papillomatosis in the sheat-fish, Silurus glanis L. Journal of Fish Diseases, 6(1):83–84, January 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Olufemi:1986:ICA

[OR86] B. E. Olufemi and R. J. Roberts. Induction of clinical aspergillomycosis by feeding contaminated diet to tilapia, Oreochromis niloticus (L.). Journal of Fish Diseases, 9(2):123–128, March 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Okamoto:1983:ARS

[OSHF83] N. Okamoto, T. Sano, R. P. Hedrick, and J. L. Fryer. Antigenic relationships of selected strains of infectious pancreatic necrosis virus and European eel virus. Journal of Fish Diseases, 6(1):19– 25, January 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Paperna:1980:AOB

[Pap80] I. Paperna. Amyloodinium ocellatum (Brown, 1931) (Dinoflagellida) infestations in cultured marine fish at Eilat, Red Sea: epizootiology and pathology. Journal of Fish Diseases, 3(5):363–372, September 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Paperna:1982:KIG

[Pap82] I. Paperna. Kudoa infection in the glomeruli, mesentery and peritoneum of cultured Sparus aurata L. Journal of Fish Diseases, 5 (6):539–543, November 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Paperna:1983:CLM

[Pap83] I. Paperna. A Chloromyxum-like myxosporean infection in the stomach of cultured Sparus aurata (L.). Journal of Fish Diseases, 6(1):85–89, January 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pinto:1989:OVE

[PAPBJ89] R. M. Pinto, P. Alvarez-Pellitero, A. Bosch, and J. Jofre. Occurrence of a viral erythrocytic infection in the Mediterranean sea bass, Dicentrarchus labrax (L.). Journal of Fish Diseases, 12(2): 185–191, March 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Persson:1987:IHN

[PCS87] M. Persson, L. Cerenius, and K. Söderhäll. The influence of haemocyte number on the resistance of the freshwater crayfish, Pacifastacus leniusculus Dana, to the parasitic fungus Aphanomyces astaci. Journal of Fish Diseases, 10(6):471–477, November 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Paperna:1984:DCE

[PD84] I. Paperna and A. P. Alves Dematos. The developmental cycle of epitheliocystis in carp, Cyprinus carpio L. Journal of Fish Diseases, 7(2):137–147, March 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Paterson:1981:IRA

[PDW81] W. D. Paterson, D. Desautels, and J. M. Weber. The immune response of Atlantic salmon, Salmo salar L., to the causative agent of bacterial kidney disease, Renibacterium salmoninarum. Journal of Fish Diseases, 4(2):99–111, March 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Paperna:1980:PHS

[PHK80] I. Paperna, J. G. Harrison, and G. W. Kissil. Pathology and histopathology of a systemic granuloma in Sparus aurata (L.) cultured in the Gulf of Aqaba. Journal of Fish Diseases, 3(3):213–221, May 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Plumb:1988:AAH

[PK88] J. A. Plumb and P. Klesius. An assessment of the antigenic homogeneity of Edwardsiella ictaluri using monoclonal antibody. Journal of Fish Diseases, 11(6):499–509, November 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pulsford:1982:USM

[PM82] Ann Pulsford and R. A. Matthews. An ultrastructural study of Myxobolus exlguus Thelohan, 1895 (Myxosporea) from grey mullet, Crenimugil labrosus (Risso). Journal of Fish Diseases, 5(6): 509–526, November 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pulsford:1984:USC

[PM84] Ann Pulsford and R. A. Matthews. An ultrastructural study of the cellular response of the plaice, Pleuronectes platessa L., to Rhipidocotyle johnstonei nom. nov. (pro- Gasterostomum sp. Johnstone, 1905) Matthews, 1968 (Digenea: Bucephalidae). Journal of Fish Diseases, 7(1):3–14, January 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Pierce:1980:HAL

[PMW80] K. V. Pierce, B. B. Mccain, and S. R. Wellings. Histopathology of abnormal livers and other organs of starry flounder Platichthys stellatus (Pallas) from the estuary of the Duwamish River, Seattle, Washington, U.S.A. Journal of Fish Diseases, 3(2):81–91, March 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Peters:1986:TSF

[PN86] F. Peters and M. Neukirch. Transmission of some fish pathogenic viruses by the heron, Ardea cinerea. Journal of Fish Diseases, 9 (6):539–544, November 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Poynton:1987:VCA

[Poy87] Sarah L. Poynton. Vertebral column abnormalities in brown trout, Salmo trutta L. Journal of Fish Diseases, 10(1):53–57, January 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Page:1984:RES

[PR84] M. Page and A. F. Rowley. The reticulo-endothelial system of the adult river lamprey, Lampetra fluviatilis (L.): the fate of intravascularly injected colloidal carbon. Journal of Fish Diseases, 7(5):339–353, September 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Peleteiro:1985:ILE

[PR85] Maria C. Peleteiro and R. H. Richards. Identification of lymphocytes in the epidermis of the rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 8(2):161–172, March 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Phillips:1985:TCM

[PRSC85] M. J. Phillips, R. J. Roberts, J. A. Stewart, and G. A. Codd. The toxicity of the cyanobacterium Microcystis aeruginosa to rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 8(4): 339–344, July 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Plumb:1983:SFS

[PS83] J. A. Plumb and D. J. Sanchez. Susceptibility of five species of fish to Edwardsiella ictaluri. Journal of Fish Diseases, 6(3):261–266, May 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Paperna:1982:OLS

[PSC82] I. Paperna, H. Ilana Sabnai, and A. Colorni. An outbreak of lymphocystis in Sparus aurata L. in the Gulf of Aqaba, Red Sea. Journal of Fish Diseases, 5(5):433–437, September 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Price:1989:PSS

[PSDS89] N. C. Price, L. Stevens, D. Duncan, and M. Snodgrass. Proteases secreted by strains of Aeromonas salmonicida. Journal of Fish Diseases, 12(3):223–232, May 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Peters:1983:NIX

[PSKS83] N. Peters, Wiebke Schmidt, Heidemarie Kranz, and H. F. Stich. Nuclear inclusions in the X-cells of skin papillomas of Pacific flatfish. Journal of Fish Diseases, 6(6):533–536, November 1983. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Paperna:1981:USP

[PSZ81] I. Paperna, I. Sabnai, and A. Zachary. Ultrastructural studies in piscine epitheliocystis: evidence for a pleomorphic developmental cycle. Journal of Fish Diseases, 4(6):459–472, November 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Paperna:1987:LIS

[PVdM87] I. Paperna, T. M. Ventura, and A. P. de Matos. Lymphocystis infection in snakeskin gourami, Trichogaster pectoralis (Regan), (Anabantidae). Journal of Fish Diseases, 10(1):11–19, January 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Poston:1985:NRO

[PW85] H. A. Poston and Marilyn J. Wolfe. Niacin requirement for optimum growth, feed conversion and protection of rainbow trout, Salmo gairdneri Richardson, from ultraviolet-B irradiation. Journal of Fish Diseases, 8(5):451–460, September 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Puckridge:1989:MDF

[PWL^{+}89] J. T. Puckridge, K. F. Walker, J. S. Langdon, C. Daley, and G. W. Beakes. Mycotic dermatitis in a freshwater gizzard shad, the bony bream, Nematalosa erebi (Günther), in the River Murray, South Australia. Journal of Fish Diseases, 12(3):205–221, May 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roald:1981:HFE

[RAL81] S. O. Roald, D. Armstrong, and T. Landsverk. Histochemical, fluorescent and electron microscopical appearance of hepatocellular ceroidosis in the Atlantic salmon Salmo salar L. Journal of Fish Diseases, 4(1):1–14, January 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rand:1986:HIP

[Ran86] T. G. Rand. The histopathology of infestation of Paranthias furdfer (L.) (Osteichthyes: Serranidae) by Nerocila acuminata (Schioedte and Meinert) (Crustacea: Isopoda: Cymothoidae). Journal of Fish Diseases, 9(2):143–146, March 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rodgers:1981:SVP

[RB81] L. J. Rodgers and J. B. Burke. Seasonal variation in the prevalence of ‘red spot’ disease in estuarine fish with particular reference to the sea mullet, Mugil cephalus L. Journal of Fish Diseases, 4(4): 297–307, July 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roubal:1987:DBH

[RB87] F. R. Roubal and A. M. Bullock. Differences between the hostparasite interface of Ichthyobodo necator (Henneguy, 1883) on the skin and gills of salmonids. Journal of Fish Diseases, 10(3):237– 240, May 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Reimschuessel:1987:EGC

[RBML87] R. Reimschuessel, R. O. Bennett, E. B. May, and M. M. Lipsky. Eosinophilic granular cell response to a microsporidian infection in a sergeant major fish, Abudefduf saxatilis (L.). Journal of Fish Diseases, 10(4):319–322, July 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roubal:1987:UAI

[RBRR87] F. R. Roubal, A. M. Bullock, D. A. Robertson, and R. J. Roberts. Ultrastructural aspects of infestation by Ichthyobodo necator (Henneguy, 1883) on the skin and gills of the salmonids Salmo salar L. and Salmo gairdneri Richardson. Journal of Fish Diseases, 10(3): 181–192, May 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rourke:1984:LEM

[RDB84] A. W. Rourke, R. W. Davis, and T. M. Bradley. A light and electron microscope study of epitheliocystis in juvenile steelhead trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 7(4): 301–309, July 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rose:1989:IDR

[REM89] A. S. Rose, A. E. Ellis, and A. L. S. Munro. The infectivity by different routes of exposure and shedding rates of Aeromonas salmonicida subsp. salmonicida in Atlantic salmon, Salmo salar L., held in sea water. Journal of Fish Diseases, 12(6):573–578, November 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Riis:1981:OMI

[RGLS81] R. C. Riis, M. E. Georgi, L. Leibovitz, and J. S. Smith. Ocular metacercarial infection of the oyster toadfish, Opsanus tau (L.). Journal of Fish Diseases, 4(5):433–435, September 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rothbard:1980:ASL

[RHI80] S. Rothbard, G. Hulata, and J. Itzkovitch. Abnormalities in Sarotherodon larvae. Journal of Fish Diseases, 3(5):441–442, September 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Raptopoulou:1987:PAP

[RL87] F. A. Raptopoulou and R. H. Lambertsen. Parasite-associated pathology of the dolphin fish, Coryphaena hippurus L., from Florida waters. Journal of Fish Diseases, 10(5):379–384, September 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roubal:1989:SCG

[RLF89] F. R. Roubal, R. J. G. Lester, and C. K. Foster. Studies on cultured and gill-attached Paramoeba sp. (Gymnamoebae: Paramoebidae) and the cytopathology of paramoebic gill disease in Atlantic salmon, Salmo salar L., from Tasmania. Journal of Fish Diseases, 12(5):481–492, September 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rasheed:1985:HBF

[RLP85] Victoria Rasheed, C. Limsuwan, and J. Plumb. Histopathology of bullminnows, Fundulus grandis Baird & Girard, infected with a non-haemolytic group B Streptococcus sp. Journal of Fish Diseases, 8(1):65–74, January 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ransom:1984:CHC

[RLRF84] D. P. Ransom, C. N. Lannan, J. S. Rohovec, and J. L. Fryer. Comparison of histopathology caused by Vibrio anguillarum and Vibrio ordalii in three species of Pacific salmon. Journal of Fish Diseases, 7(2):107–115, March 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ralphs:1986:HMJ

[RM86] J. R. Ralphs and R. A. Matthews. Hepatic microsporidiosis of juvenile grey mullet, Chelon labrosus (Risso), due to Microgemma hepaticus gen.nov. sp.nov. Journal of Fish Diseases, 9(3):225–242, May 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Radhakrishnan:1981:HID

[RN81a] S. Radhakrishnan and N. B. Nair. Histopathology of the infestation of Diodon hystrix L. by Peniculisa wilsoni Radhakrishnan (Copepoda: Lernaeoceridae). Journal of Fish Diseases, 4(1):83–87, January 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Reno:1981:UPV

[RN81b] P. W. Reno and B. L. Nicholson. Ultrastructure and prevalence of viral erythrocytic necrosis (VEN) virus in Atlantic cod, Gadus morhua L., from the northern Atlantic Ocean. Journal of Fish Diseases, 4(5):361–370, September 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:1983:REH

[Rob83] M. S. Roberts. A report of an epizootic in hatchery reared rainbow trout, Salmo gairdneri Richardson, at an English trout farm, caused by Yersinia ruckeri. Journal of Fish Diseases, 6(6):551– 552, November 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roubal:1986:HLA

[Rou86] F. R. Roubal. Histopathology of leech, Austrobdella bilobata Ingram, infestation on the yellowfin bream, Acanthopagrus australis (Günther), in northern New South Wales. Journal of Fish Diseases, 9(3):213–223, May 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roubal:1989:EMO

[RPL89] F. R. Roubal, J. L. Paynter, and R. J. G. Lester. Electron microscopic observation of hepatopancreatic parvo-like virus (HPV) in the penaeid prawn, Penaeus merguiensis de Man, from Australia. Journal of Fish Diseases, 12(2):199–201, March 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rodak:1988:ELI

[RPT^{+}88] L. Rodák, Z. Pospíšil, J. Tománek, T. Veselý, T. Obr, and L. Valíček. Enzyme-linked immunosorbent assay (ELISA) detection of infectious pancreatic necrosis virus (IPNV) in culture fluids and tissue homogenates of the rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 11(3):225–235, May 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rousset:1989:PCI

[RR89] V. Rousset and A. Raibaut. Peculiar cases of intracardiac parasitism in the pilchard, Sardina pilchardus (Walbaum), by a pennellid copepod belonging to the genus Lernaeenicus. Journal of Fish Diseases, 12(3):263–268, May 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Robertson:1981:PAS

[RRB81] D. A. Robertson, R. J. Roberts, and A. M. Bullock. Pathogenesis and autoradiographic studies of the epidermis of salmonids infested with Ichtyobodo necator (Henneguy, 1883). Journal of Fish Diseases, 4(2):113–125, March 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Rintamaki:1986:OYR

[RVF86] Päivi Rintamäki, E. Tellervo Valtonen, and G. N. Frerichs. Occurrence of Yersinia ruckeri infection in farmed whitefish, Coregonus peled Gmelin and Coregonus muksun Pallas, and Atlantic salmon, Salmo salar L., in northern Finland. Journal of Fish Diseases, 9(2): 137–140, March 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:1980:E

[RW80] Ronald J. Roberts and Rod Wootten. Editorial. Journal of Fish Diseases, 3(2):77, March 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Roberts:1984:E

[RW84] Ronald J. Roberts and Rod Woottien. Editorial. Journal of Fish Diseases, 7(1):1, January 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sanaullah:1980:GMM

[SA80] M. Sanaullah and A. T. A. Ahmed. Gill myxoboliasis of major carps in Bangladesh. Journal of Fish Diseases, 3(4):349–354, July 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stevenson:1984:SVA

[SA84] R. M. W. Stevenson and D. W. Airdrie. Serological variation among Yersinia ruckeri strains. Journal of Fish Diseases, 7(4): 247–254, July 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sakai:1987:CSS

[SAAK87a] M. Sakai, N. Amaaki, S. Atsuta, and M. Kobayashi. Comparative sensitivities of several dot blot methods used to detect bacterial kidney diseases. Journal of Fish Diseases, 10(3):229–231, May 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sakai:1987:CSD

[SAAK87b] M. Sakai, N. Amaaki, S. Atsuta, and M. Kobayashi. Comparative sensitivity of dot blot methods used to detect Renibacteriwn salmoninarum. Journal of Fish Diseases, 10(5):415–418, September 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sakai:1984:NSA

[Sak84a] D. K. Sakai. The non-specific activation of rainbow trout, Salmo gairdneri Richardson, complement by Aeromonas salmonicida extracellular products and the correlation of complement activity with the inactivation of lethal toxicity products. Journal of Fish Diseases, 7(5):329–338, September 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sakai:1984:OFA

[Sak84b] D. K. Sakai. Opsonization by fish antibody and complement in the immune phagocytosis by peritoneal exudate cells isolated from salmonid fishes. Journal of Fish Diseases, 7(1):29–38, January 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sakai:1985:ESA

[Sak85] D. K. Sakai. Efficacy of specific antibody against agglutinating Aeromonas salmonicida strains on infectivity and vaccination with inactivated protease. Journal of Fish Diseases, 8(5):397–405, September 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sakai:1987:TPG

[Sak87] D. K. Sakai. Transmission of protease genes into a proteasedeficient mutant of Aeromonas salmonicida in river sediments. Journal of Fish Diseases, 10(3):171–179, May 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sano:1982:B

[San82] Tokuo Sano. Biography. Journal of Fish Diseases, 5(6):447–448, November 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sano:1987:EPD

[San87] Tokuo Sano. Emeritus Professor Dr Syuzo Egusa. Journal of Fish Diseases, 10(5):333–334, September 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Speyerer:1987:PPE

[SB87] P. D. Speyerer and J. A. Boyle. The plasmid profile of Edwardsiellaictaluri. Journal of Fish Diseases, 10(6):461–469, November 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Seagrave:1981:SPD

[SBHM81] C. P. Seagrave, D. Bucke, E. B. Hudson, and D. Mcgregor. A survey of the prevalence and distribution of proliferative kidney disease (PKD) in England and Wales. Journal of Fish Diseases, 4 (5):437–439, September 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Swanson:1982:PIP

[SCG82] R. N. Swanson, J. C. Carlisle, and J. H. Gillespie. Pathogenesis of infectious pancreatic necrosis virus infection in brook trout, Salvelinus fontinalis (Mitchill), following intraperitoneal injection. Journal of Fish Diseases, 5(6):449–460, November 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Santacana:1982:AFB

[SCM^{+}82] J. A. Santacana, D. A. Conroy, M. E. Mujica, C. Marín, and N. De López. Acid–fast bacterial infection and its control in three-spot gouramies, Trichogaster trichopterus Pallas. Journal of Fish Diseases, 5(6):545–547, November 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Scott:1985:DCI

[Sco85] Marilyn E. Scott. Dynamics of challenge infections of Gyrodactylus bullatarudis Turnbull (Monogenea) on guppies, Poecilia reticulata (Peters). Journal of Fish Diseases, 8(6):495–503, November 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stave:1987:CRS

[SCR87] J. W. Stave, T. M. Cook, and B. S. Roberson. Chemiluminescent responses of striped bass, Morone saxatilis (Walbaum), phagocytes to strains of Yersinia ruckeri. Journal of Fish Diseases, 10(1): 1–10, January 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Savan:1980:EVR

[SD80] M. Savan and P. Dobos. Effect of virazole on rainbow trout Salmo gairdneri Richardson fry infected with infectious pancreatic necrosis virus. Journal of Fish Diseases, 3(5):437–440, September 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shostak:1986:IPN

[SD86] A. W. Shostak and T. A. Dick. Intestinal pathology in northern pike, Esox lucius L., infected with Triaenophorus crassus Forel, 1868 (Cestoda: Pseudophyllidea). Journal of Fish Diseases, 9(1): 35–45, January 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sanchez:1989:IHR

[SDC89] C. Sanchez, J. Dominguez, and J. Coll. Immunoglobulin heterogeneity in the rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 12(5):459–465, September 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Smail:1980:VIMb

[SE80a] D. A. Smail and S. I. Egglestone. Virus infections of marine fish erythrocytes: electron microscopical studies on the blenny virus. Journal of Fish Diseases, 3(1):47–54, January 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Smail:1980:VIMa

[SE80b] D. A. Smail and S. I. Egglestone. Virus infections of marine fish erythrocytes: prevalence of piscine erythrocytic necrosis in cod Gadus morhua L. and blenny Blennius pholis L. in coastal and offshore waters of the United Kingdom. Journal of Fish Diseases, 3(1): 41–46, January 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Secombes:1985:VFT

[Sec85] C. J. Secombes. The in vitro formation of teleost multinucleate giant cells. Journal of Fish Diseases, 8(5):461–464, September 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Smith:1989:HPE

[SFH89] I. R. Smith, H. W. Ferguson, and M. A. Hayes. Histopathology and prevalence of epidermal papillomas epidemic in brown bullhead, Ictalurus nebulosus (Lesueur), and white sucker, Catostomus commersoni (Lacépède), populations from Ontario, Canada. Journal of Fish Diseases, 12(4):373–388, July 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Swanson:1981:IFA

[SG81] R. N. Swanson and J. H. Gillespie. An indirect fluorescent antibody test for the rapid detection of infectious pancreatic necrosis virus in tissues. Journal of Fish Diseases, 4(4):309–315, July 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shariff:1981:HEB

[Sha81] M. Shariff. The histopathology of the eye of big head carp, Aristichthys noblis (Richardson), infested with Lernaea piscinae Harding, 1950. Journal of Fish Diseases, 4(2):161–168, March 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shariff:1982:HSS

[Sha82] M. Shariff. Henneguya shaharini sp. nov. (Protozoa: Myxozoa), a parasite of marble goby, Oxyeleotns marmoratus (Bleeker). Journal of Fish Diseases, 5(1):37–45, January 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stott:1983:HSO

[SHNS83] G. G. Stott, W. E. Haensly, J. M. Neff, and J. R. Sharp. Histopathologic survey of ovaries of plaice, Pleuronectes platessa L., from Aber Wrac’h and Aber Benoit, Brittany, France: longterm effects of the Amoco Cadiz crude oil spill. Journal of Fish Diseases, 6(5):429–437, September 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Smith:1980:IHI

[SI80] C. E. Smith and T. Inslee. Interlamellar Henneguya infestation in adult channel catfish Ictalums punctatus (Rafinesque). Journal of Fish Diseases, 3(3):257–260, May 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Slooff:1982:HBA

[SK82] W. Slooff and E. Klootwijk-Van Dijk. Hermaphroditism in the bream, Abramis brama (L.). Journal of Fish Diseases, 5(1):79–81, January 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

S:1985:SST

[SK85] R. L. Herman U. S. and F. W. Kircheis. Steatitis in Sunapee trout, Salvelinus alpinus oquassa Girard. Journal of Fish Diseases, 8(2): 237–239, March 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Soderberg:1987:FCL

[SK87] R. W. Soderberg and W. F. Krise. Fin condition of lake trout, Salvelinus namaycush Walbaum, reared at different densities. Journal of Fish Diseases, 10(3):233–235, May 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shariff:1986:HSL

[SKS86] M. Shariff, Z. Kabata, and Christina Sommerville. Host susceptibility to Lernaea cyprinacea L. and its treatment in a large aquarium system. Journal of Fish Diseases, 9(5):393–401, September 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stoskopf:1986:TAA

[SKSA^{+}86] M. K. Stoskopf, Suzanne Kennedy-Stoskopf, Jill Arnold, J. Andrews, and Marie T. Perlstein. Therapeutic aminoglycoside antibiotic levels in brown shark, Carcharhinus plumbeus (Nardo). Journal of Fish Diseases, 9(4):303–311, July 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Smith:1988:SHM

[SLH88] C. A. D. Smith, M. J. Louis, and F. M. Hetrick. A sequence homologous to the mammalian p53 oncogene in fish cell lines. Journal of Fish Diseases, 11(6):525–530, November 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Song:1989:CMI

[SLK89] Y. L. Song, T. Lin, and G. H. Kou. Cell-mediated immunity of the eel, Anguilla japonica (Temminck & Schlegel), as measured by the migration inhibition test. Journal of Fish Diseases, 12(2):117–123, March 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Schultz:1986:MCF

[SLMH86] C. L. Schultz, B. C. Lidgerding, P. E. Mcallister, and F. M. Hetrick. Mycoplasma contamination in fish cell lines: an evaluation of detection by differential incorporation of ^{3}H-uridine and ^{14}Curacil. Journal of Fish Diseases, 9(2):117–122, March 1986. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Secombes:1980:CSI

[SM80] C. J. Secombes and Margaret J. Manning. Comparative studies on the immune system of fishes and amphibians: antigen localization in the carp Cyprinus carpio L. Journal of Fish Diseases, 3(5):399– 412, September 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Smith:1989:BKW

[SM89] Charlie E. Smith and Joe McCraren. Biography: Ken Wolf. Journal of Fish Diseases, 12(5):401–403, September 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Schutz:1984:IIP

[SMKH84] Marian Schutz, E. B. May, J. N. Kraeuter, and F. M. Hetrick. Isolation of infectious pancreatic necrosis virus from an epizootic occurring in cultured striped bass, Morone saxatilis (Walbaum). Journal of Fish Diseases, 7(6):505–507, November 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Smith:1984:PKD

[SMRF84] C. E. Smith, J. K. Morrison, H. W. Ramsey, and H. W. Ferguson. Proliferative kidney disease: first reported outbreak in North America. Journal of Fish Diseases, 7(3):207–216, May 1984. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stephens:1980:VAA

[SNZH80] E. B. Stephens, M. W. Newman, A. L. Zachary, and F. M. Hetrick. A viral aetiology for the annual spring epizootics of Atlantic menhaden Brevoortia tyrannus (Latrobe) in Chesapeake Bay. Journal of Fish Diseases, 3(5):387–398, September 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Solangi:1982:HCT

[SO82] M. A. Solangi and R. M. Overstreet. Histopathological changes in two estuarine fishes, Menidia beryllina (Cope) and Trinectes maculatus (Bloch and Schneider), exposed to crude oil and its watersoluble fractions. Journal of Fish Diseases, 5(1):13–35, January 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sakai:1989:CSC

[SOAK89] M. Sakai, K. Ogasawara, S. Atsuta, and M. Kobayashi. Comparative sensitivity of carp, Cyprinus carpio L. and rainbow trout, Salmo gairdneri Richardson, to Renibacterium salmoninarum. Journal of Fish Diseases, 12(4):367–372, July 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Soderberg:1985:HRT

[Sod85] R. W. Soderberg. Histopathology of rainbow trout, Salmo gairdneri Richardson, exposed to diurnally fluctuating un-ionized ammonia levels in static-water ponds. Journal of Fish Diseases, 8(1): 57–64, January 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sommerville:1981:CST

[Som81] Christina Sommerville. A comparative study of the tissue response to invasion and encystment by Stephanochasmus baccatus (Nicoll, 1907) (Digenea: Acanthocolpidae) in four species of flatfish. Journal of Fish Diseases, 4(1):53–68, January 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sommerville:1982:LHH

[Som82a] Christina Sommerville. The life history of Haplorchis pumilio (Looss, 1896) from cultured tilapias. Journal of Fish Diseases, 5 (3):233–241, May 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sommerville:1982:PHP

[Som82b] Christina Sommerville. The pathology of Haplorchis pumilio (Looss, 1896) infections in cultured tilapias. Journal of Fish Diseases, 5(3):243–250, May 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sano:1981:NVE

[SON81] T. Sano, N. Okamoto, and T. Nishimura. A new viral epizootic of Anguilla japonica Temminck and Schlegel. Journal of Fish Diseases, 4(2):127–139, March 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Saeed:1987:SDE

[SP87] M. O. Saeed and J. A. Plumb. Serological detection of Edwardsiella ictaluri Hawke lipopolysaccharide antibody in serum of channel catfish, Ictalurus punctatus Rafinesque. Journal of Fish Diseases, 10(3):205–209, May 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Scott:1980:HEP

[SR80] A. L. Scott and W. A. Rogers. Histological effects of prolonged sublethal hypoxia on channel catfish Ictalurus punctatus (Rafinesque). Journal of Fish Diseases, 3(4):305–316, July 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shariff:1989:EHL

[SR89] M. Shariff and R. J. Roberts. The experimental histopathology of Lernaea polymorpha Yu, 1938 infection in naive Aristichthys nobilis (Richardson) and a comparison with the lesion in naturally infected clinically resistant fish. Journal of Fish Diseases, 12(5):405–414, September 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stave:1985:CRS

[SRH85] J. W. Stave, B. S. Roberson, and F. M. Hetrick. Chemiluminescent responses of striped bass, Morone saxatilis (Walbaum), phagocytes to Vibrio spp. Journal of Fish Diseases, 8(5):479–483, September 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shariff:1980:HAC

[SRS80] M. Shariff, R. H. Richards, and Christina Sommerville. The histopathology of acute and chronic infections of rainbow trout Salmo gairdneri Richardson with eye flukes, Diplostomum spp. Journal of Fish Diseases, 3(6):455–465, November 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Simon:1984:TSS

[SS84] R. C. Simon and W. B. Schtll. Tables of sample size requirements for detection of fish infected by pathogens: three confidence levels for different infection prevalence and various population sizes. Journal of Fish Diseases, 7(6):515–520, November 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Subasinghe:1988:SEM

[SS88] R. P. Subasinghe and C. Sommerville. Scanning electron microscopy study of the causes of mortality in Oreochromis mossambicus (Peters) eggs under artificial incubation. Journal of Fish Diseases, 11(5):409–416, September 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Subasinghe:1989:ESP

[SS89] R. P. Subasinghe and C. Sommerville. An experimental study into the possible protection of fry from Ichthyophthirius multifiliis (Fouquet) infections during mouthbrooding of Oreochromis mossambicus (Peters). Journal of Fish Diseases, 12(2):143–149, March 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Spitsbergen:1988:ETP

[SSKP88] J. M. Spitsbergen, K. A. Schat, J. M. Kleeman, and R. E. Peterson. Effects of 2,3,7,8-tetrachlorodibenzo-p-dioxin (TCDD) or Aroclor 1254 on the resistance of rainbow trout, Salmo gairdneri Richardson, to infectious haematopoietic necrosis virus. Journal of Fish Diseases, 11(1):73–83, January 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stehr:1986:SMK

[Ste86] Carla Stehr. Sporogenesis of the myxosporean Kudoa paniformis Kabata & Whitaker, 1981 infecting the muscle of the Pacific whiting, Merluccius productus (Ayres). Journal of Fish Diseases, 9 (6):493–504, November 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shotts:1980:AUD

[STEM80] E. B. Shotts, Jr., P. D. Talkington, Diane G. Elliott, and D. H. Mccarthy. Aetiology of an ulcerative disease in goldfish, Carassius auratus (L.): characterization of the causative agent. Journal of Fish Diseases, 3(3):181–186, May 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Skurdal:1988:TCH

[STF^{+}88] J. Skurdal, T. Taugbøl, E. Fjeld, D. O. Hessen, and T. Håstein. Thelohania contejeani Henneguy parasitizing the noble crayfish, Astacus astacus L., in Norway. Journal of Fish Diseases, 11(5):433– 435, September 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stehr:1986:HPI

[SW86] Carla Stehr and D. J. Whitaker. Host-parasite interaction of the myxosporeans Kudoa paniformis Kabata & Whitaker, 1981 and Kudoa thyrsites (Gilchrist, 1924) in the muscle of Pacific whiting, Merluccius productus (Ayres): an ultrastructural study. Journal of Fish Diseases, 9(6):505–517, November 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Stewart:1983:OSB

[SWDM83] D. J. Stewart, K. Woldemariam, G. Dear, and Francesca M. Mochaba. An outbreak of ‘Sekiten-byo’ among cultured European eels, Anguilla anguilla L., in Scotland. Journal of Fish Diseases, 6(1):75–76, January 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Sauter:1987:SBP

[SWM^{+}87] R. W. Sauter, C. Williams, E. A. Meyer, B. Celnik, J. L. Banks, and D. A. Leith. A study of bacteria present within unfertilized salmon eggs at the time of spawning and their possible relation to early lifestage disease. Journal of Fish Diseases, 10(3):193–203, May 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Smith:1987:RDE

[SZ87] I. R. Smith and B. A. Zajdlik. Regression and development of epidermal papillomas affecting white suckers, Catostomus commersoni (Lacépède), from Lake Ontario, Canada. Journal of Fish Diseases, 10(6):487–494, November 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Shariff:1986:PBC

[SZA86] M. Shariff, A. T. Zainuddin, and H. Abdullah. Pugheadedness in bighead carp, Aristichthys nobilis (Richardson). Journal of Fish Diseases, 9(5):457–460, September 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tsoumas:1989:ASD

[TAR89] A. Tsoumas, D. J. Alderman, and C. J. Rodgers. Aeromonas salmonicida: development of resistance to 4-quinolone antimicrobials. Journal of Fish Diseases, 12(5):493–507, September 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tsing:1987:NVD

[TB87] A. Tsing and J. R. Bonami. A new viral disease of the tiger shrimp, Penaeus japonicus Bate. Journal of Fish Diseases, 10(2):139–141, March 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thiyagarajah:1988:LPL

[TB88] A. Thiyagarajah and M. E. Bender. Lesions in the pancreas and liver of an oyster toadfish, Opsanus tau (L.), collected from the lower York River, Virginia, USA. Journal of Fish Diseases, 11(4): 359–364, July 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Toranzo:1989:AAS

[TBRH89] A. E. Toranzo, A. M. Baya, J. L. Romalde, and F. M. Hetrick. Association of Aeromonas sobria with mortalities of adult gizzard shad, Dorosoma cepedianum Lesueur. Journal of Fish Diseases, 12 (5):439–448, September 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tave:1982:SSA

[TBS82] D. Tave, J. E. Bartels, and R. O. Smitherman. Stumpbody Sarotherodon aureus (Steindachner) (= Tilapia aurea) and tailless S. niloticus (L.) (= T. niloticd): two vertebral anomalies and their effects on body length. Journal of Fish Diseases, 5(6):487– 494, November 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tave:1983:SDL

[TBS83] D. Tave, J. E. Bartels, and R. O. Smitherman. Saddleback: a dominant, lethal gene in Sarotherodon aureus (Steindachner) (= Tilapia aurea). Journal of Fish Diseases, 6(1):59–73, January 1983. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Timmons:1980:PRL

[TH80] T. J. Timmons and W. G. Hemstreet. Prevalence rate of Lernaea cyprinacea L. (Copepoda: Lernaeidae) on young-of-the-year largemouth bass, Micropterus salmoides (Lacépède), in West Point Reservoir, Alabama–Georgia, U.S.A. Journal of Fish Diseases, 3 (6):529–530, November 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Taylor:1989:LIC

[TH89] M. Taylor and D. Hoole. Ligula intestinalis (L.) (Cestoda: Pseudophyllidea): an ultrastructural study of the cellular response of roach fry, Rutilis rutilis L., to an unusual intramuscular infection. Journal of Fish Diseases, 12(5):523–528, September 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thuvander:1987:DPI

[THJS87] Ann Thuvander, T. Hongslo, Eva Jansson, and Bo Sundquist. Duration of protective immunity and antibody titres measured by ELISA after vaccination of rainbow trout, Salmo gairdneri Richardson, against vibriosis. Journal of Fish Diseases, 10(6):479– 486, November 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thompson:1982:ELN

[Tho82] J. S. Thompson. An epizootic of lymphoma in northern pike, Esox lucius L., from the Åland Islands of Finland. Journal of Fish Diseases, 5(1):1–11, January 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thune:1986:AHB

[TJGA86] R. L. Thune, Maureen C. Johnson, T. E. Graham, and R. L. Amborski. Aeromonas hydrophila B-haemolysin: purification and examination of its role in virulence in 0-group channel catfish, Ictalurus punctatus (Rafinesque). Journal of Fish Diseases, 9(1): 55–61, January 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Toranzo:1982:CST

[TM82] Alicia E. Toranzo and F. M. Metricic. Comparative stability of two salmonid viruses and poliovirus in fresh, estuarine and marine waters. Journal of Fish Diseases, 5(3):223–231, May 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tatner:1985:ODRa

[TM85a] Mary F. Tatner and Margaret J. Manning. The ontogenetic development of the reticulo-endothelial system in the rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 8(1):35–41, January 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Tatner:1985:ODRb

[TM85b] Mary F. Tatner and Margaret J. Manning. The ontogenetic development of the reticulo-endothelial system in the rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 8(2):189– 195, March 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thompson:1988:UPC

[TM88] J. S. Thompson and M. Miettinen. Ultrastructural pathology of cutaneous tumours of northern pike, Esox lucius L. Journal of Fish Diseases, 11(1):47–55, January 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Totland:1988:USM

[TNH88] G. K. Totland, A. Nylund, and K. O. Holm. An ultrastructural study of morphological changes in Atlantic salmon, Salmo salar L., during the development of cold water vibriosis. Journal of Fish Diseases, 11(1):1–13, January 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thune:1981:GLB

[TR81] R. L. Thune and W. A. Rogers. Gill lesions in bluegill, Lepomis macrochirus Rafinesque, infested with Cleidodiscus robustus Mueller, 1934 (Monogenea: Dactylogyridae). Journal of Fish Diseases, 4(3):277–280, May 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Traxler:1986:EIH

[Tra86] G. S. Traxler. An epizootic of infectious haematopoietic necrosis in 2-year-old Kokanee, Oncorhynchus nerka (Walbaum) at Lake Cowichan, British Columbia. Journal of Fish Diseases, 9(6):545– 549, November 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Turnbull:1989:ESB

[TRT89] J. F. Turnbull, R. H. Richards, and M. F. Tatner. Evidence that superficial branchial colonies on the gills of Salmo salar L. are not Aeromonas salmonicida. Journal of Fish Diseases, 12(5):449–458, September 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Timur:1984:GCR

[TT84] G. Timur and M. Timur. Giant-cell reaction associated with Ichthyophonus hoferi infection in wild plaice, Pleuronectes platessa L. Journal of Fish Diseases, 7(6):513–514, November 1984. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Timur:1985:MEC

[TTR85] M. Timur, G. Timor, and R. J. Roberts. Moderation of the experimental carrageenin granuloma in the plaice, Pleuronectes platessa L., by reduction in temperature. Journal of Fish Diseases, 8 (6):523–527, November 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thomas:1988:CSV

[TW88] P. T. Thomas and P. T. K. Woo. Cryptobia salmositica: an in vitro and in vivo study on the mechanism of anaemia in infected rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 11 (5):425–431, September 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thomas:1989:CAS

[TW89a] P. T. Thomas and P. T. K. Woo. Complement activity in Salmo gairdneri Richardson infected with Cryptobia salmositica (Sarcomastigophora: Kinetoplastida) and its relationship to the anaemia in cryptobiosis. Journal of Fish Diseases, 12(4):395–397, July 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Thomas:1989:VSH

[TW89b] P. T. Thomas and P. T. K. Woo. An in vitro study on the haemolytic components from Cryptobia salmositica (Sarcomastigophora: Kinetoplastida). Journal of Fish Diseases, 12(4): 389–393, July 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Urdaneta:1982:LRC

[Urd82] H. E. Urdaneta. Lymphocystis in róbalo, Centropomus undecimalis (Bloch), in Lake Maracaibo, Venezuela. Journal of Fish Diseases, 5 (4):347–348, July 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Vivares:1988:UOL

[VA88] C. P. Vivarès and C. Azevedo. Ultrastructural observations of the life cycle stages of Ameson atlanticum sp. nov., a microsporidan parasitizing Cancer pagurus L. Journal of Fish Diseases, 11 (5):379–387, September 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Ventura:1988:LAN

[VG88] M. T. Ventura and J. M. Grizzle. Lesions associated with natural and experimental infections of Aeromonas hydrophila in channel catfish, Ictalurus punctatus (Rafinesque). Journal of Fish Diseases, 11(5):397–407, September 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Valtonen:1981:IAS

[VK81] E. Tellervo Valtonen and Anna-Liisa Keränen. Ichthyophthiriasis of Atlantic salmon, Salmo salar L., at the Montta Hatchery in northern Finland in 1978–1979. Journal of Fish Diseases, 4(5):405– 411, September 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Waterstrat:1988:UDP

[WAC88] P. R. Waterstrat, A. J. Ainsworth, and G. Capley. Use of a discontinuous Percoll gradient technique for the separation of channel catfish, Ictalurus punctatus (Rafinesque), peripheral blood leucocytes. Journal of Fish Diseases, 11(4):289–294, July 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Waterstrat:1989:UIE

[WAC89] P. Waterstrat, J. Ainsworth, and G. Capley. Use of an indirect enzyme-linked immunosorbent assay (ELISA) in the detection of channel catfish, Ictalurus punctatus (Rafinesque), antibodies to Edwardsiella ictaluri. Journal of Fish Diseases, 12(2):87–94, March 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Watermann:1982:UCT

[Wat82] B. Watermann. An unidentified cell type associated with an inflammatory condition of the subcutaneous connective tissue in dab, Limanda limanda L. Journal of Fish Diseases, 5(3):257–261, May 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wabuke-Bunoti:1980:PPC

[WB80] M. A. N. Wabuke-Bunoti. The prevalence and pathology of the cestode Polyonchobothrium clarias (Woodland, 1925) in the teleost, Clarias mossambicm (Peters). Journal of Fish Diseases, 3(3):223– 230, May 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wise:1985:DCCa

[WB85] Janet A. Wise and J. A. Boyle. Detection of channel catfish virus in channel catfish, Ictalurus punctatus (Rafinesque): use of a nucleic acid probe. Journal of Fish Diseases, 8(5):417–424, September 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wise:1985:DCCb

[WBB85] Janet A. Wise, P. R. Bowser, and J. A. Boyle. Detection of channel catfish virus in asymptomatic adult channel catfish, Ictalurus punctatus (Rafinesque). Journal of Fish Diseases, 8(6):485–493, November 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Whittington:1988:SSF

[WC88] R. J. Whittington and B. Cullis. The susceptibility of salmonid fish to an atypical strain of Aeromonas salmonicida that infects goldfish, Carassius auratus (L.), in Australia. Journal of Fish Diseases, 11(6):461–470, November 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Weng:1987:PBS

[Wen87] H. T. Weng. The parasitic barnacle, Sacculina granifera Boschma, affecting the commercial sand crab, Portunus pelagicus (L.), in populations from two different environments in Queensland. Journal of Fish Diseases, 10(3):221–227, May 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Warr:1984:LTO

[WGA^{+}84] G. W. Warr, B. R. Griffin, D. P. Anderson, P. E. Mcallister, B. Lidgerding, and C. E. Smith. A lymphosarcoma of thymic origin in the rainbow trout, Salmo gairdneri Richardson. Journal of Fish Diseases, 7(1):73–82, January 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Whittington:1987:CME

[WGC^{+}87] R. J. Whittington, N. Gudkovs, M. J. Carrigan, L. D. Ashburner, and S. J. Thurstan. Clinical, microbiological and epidemiological findings in recent outbreaks of goldfish ulcer disease due to atypical Aeromonas salmonicida in south-eastern Australia. Journal of Fish Diseases, 10(5):353–362, September 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Willoughby:1985:RPS

[Wil85] L. G. Willoughby. Rapid preliminary screening of Saprolegnia on fish. Journal of Fish Diseases, 8(5):473–476, September 1985. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Willoughby:1989:CDS

[Wil89a] L. G. Willoughby. Continued defence of salmonid fish against Saprolegnia fungus, after its establishment. Journal of Fish Diseases, 12(1):63–67, January 1989. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Willumsen:1989:BWF

[Wil89b] B. Willumsen. Birds and wild fish as potential vectors of Yersinia ruckeri. Journal of Fish Diseases, 12(3):275–277, May 1989. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Woo:1987:CSC

[WLL87] P. T. K. Woo, J. F. Leatherland, and M. S. Lee. Cryptobia salmositica: cortisol increases the susceptibility of Salmo gairdneri Richardson to experimental cryptobiosis. Journal of Fish Diseases, 10(2):75–83, March 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wootten:1982:DCE

[WM82] R. Wootten and A. H. Mcvicar. Dermocystidium from cultured eels, Anguilla anguilla L., in Scotland. Journal of Fish Diseases, 5 (3):215–222, May 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wolf:1981:NMB

[WMC^{+}81] Ken Wolf, Maria E. Markiw, J. Machado Cruz, M. H. Galhano, J. Eiras, and R. L. Herman. Non-myxosporidan blacktail of salmonids. Journal of Fish Diseases, 4(4):355–357, July 1981. CO-DEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wolf:1986:SWD

[WMH86] K. Wolf, M. E. Markiw, and J. K. Hiltunen. Salmonid whirling disease: Tubifex tubifex (Müller) identified as the essential oligochaete in the protozoan life cycle. Journal of Fish Diseases, 9(1):83–85, January 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Woodhead:1982:SLT

[Woo82] Avril D. Woodhead. Skin lesions in the tail of the spiny dogfish, Squalus acanthias L. Journal of Fish Diseases, 5(1):71–74, January 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wplf:1980:FS

[Wpl80] K. Wplf. S. F. Snieszko. Journal of Fish Diseases, 3(6):451–453, November 1980. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wolf:1981:HSP

[WS81] K. Wolf and C. E. Smith. Herpesvirus salmonis: pathological changes in parenterally-infected rainbow trout, Salmo gairdneri Richardson, fry. Journal of Fish Diseases, 4(6):445–457, November 1981. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

WaagbO:1988:HBA

[WSEL88] R. WaagbØ, K. Sandnes, S. Espelid, and Ø. Lie. Haematological and biochemical analyses of Atlantic salmon, Salmo solar L., suffering from coldwater vibriosis (‘Hitra disease’). Journal of Fish Diseases, 11(5):417–423, September 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wanstall:1988:PCP

[WTR88] S. T. Wanstall, J. S. Thomas, and P. W. J. Robotham. The pathology caused by Pomphorhynchus laevis Müller in the alimentary tract of the stone loach, Noemacheilus barbatulus (L.). Journal of Fish Diseases, 11(6):511–523, November 1988. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wood:1986:PEI

[WWB86] Sarah E. Wood, L. G. Willoughby, and G. W. Beakes. Preliminary evidence for inhibition of Saprolegnia fungus in the mucus of brown trout, Salmo trutta L., following experimental challenge. Journal of Fish Diseases, 9(6):557–560, November 1986. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Wechsler:1987:TIP

[WWK^{+}87] S. J. Wechsler, L. C. Woods, J. N. Kraeuter, F. M. Hetrick, and P. E. Mcallister. Transmission of infectious pancreatic necrosis virus in striped bass, Morone saxatilis (Walbaum). Journal of Fish Diseases, 10(1):29–34, January 1987. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yamamoto:1985:EHW

[YKN85] T. Yamamoto, R. K. Kelly, and O. Nielsen. Epidermal hyperplasia of walleye, Stizostedion vitreum vitreum (Mitchill), associated with retrovirus-like type-C particles: prevalence, histologic and electron microscopic observations. Journal of Fish Diseases, 8(5):425–436, September 1985. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Yu:1982:RIP

[YMM82] K. K-Y. Yu, R. D. Macdonald, and Anne R. Moore. Replication of infectious pancreatic necrosis virus in trout leucocytes and detection of the carrier state. Journal of Fish Diseases, 5(5):401–410, September 1982. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).

Zimmer:1984:EDC

[ZEK84] M. A. Zimmer, M. S. Ewing, and K. M. Kocan. Epitheliocystis disease in the channel catfish, Ictalurus punctatus (Rafinesque). Journal of Fish Diseases, 7(5):407–410, September 1984. CODEN JFIDDI. ISSN 0140-7775 (print), 1365-2761 (electronic).